### v1

# An extrapolated and provably convergent algorithm for nonlinear matrix decomposition with the ReLU function

Abstract
Nonlinear matrix decomposition (NMD) with the ReLU function, denoted ReLU-NMD, is the following problem: given a sparse, nonnegative matrix $X$ and a factorization rank $r$, identify a rank-$r$ matrix $\Theta$ such that $X\approx\max(0,\Theta)$. This decomposition finds application in data compression, matrix completion with entries missing not at random, and manifold learning. The standard ReLU-NMD model minimizes the least squares error, that is, $\|X-\max(0,\Theta)\|_{F}^{2}$. The corresponding optimization problem is nondifferentiable and highly nonconvex. This motivated Saul to propose an alternative model, Latent-ReLU-NMD, where a latent variable $Z$ is introduced and satisfies $\max(0,Z)=X$ while minimizing $\|Z-\Theta\|_{F}^{2}$ (“A nonlinear matrix decomposition for mining the zeros of sparse data”, SIAM J. Math. Data Sci., 2022). Our first contribution is to show that the two formulations may yield different low-rank solutions $\Theta$; in particular, we show that Latent-ReLU-NMD can be ill-posed when ReLU-NMD is not, meaning that there are instances in which the infimum of Latent-ReLU-NMD is not attained while that of ReLU-NMD is. We also consider another alternative model, called 3B-ReLU-NMD, which parameterizes $\Theta=WH$, where $W$ has $r$ columns and $H$ has $r$ rows, allowing one to get rid of the rank constraint in Latent-ReLU-NMD. Our second contribution is to prove the convergence of a block coordinate descent (BCD) applied to 3B-ReLU-NMD and referred to as BCD-NMD. Our third contribution is a novel extrapolated variant of BCD-NMD, dubbed eBCD-NMD, which we prove is also convergent under mild assumptions. We illustrate the significant acceleration effect of eBCD-NMD compared to BCD-NMD, and also show that eBCD-NMD performs well against the state of the art on synthetic and real-world data sets.
keywords:
Low-rank matrix approximations, nonlinear matrix decomposition, rectified liner unit, block coordinate descent, extrapolation
{AMS}
15A23, 65F55, 68Q25, 90C26, 65K05.
1 Introduction
Nonlinear matrix decomposition (NMD) can be defined as follows: given a data matrix $X\in\mathbb{R}^{n\times m}$, a factorization rank $r$, NMD looks for a rank-$r$ matrix $\Theta\in\mathbb{R}^{m\times n}$ such that $X\approx f(\Theta)$, where $f(\Theta)$ applies the scalar function $f$ element-wise on $\Theta$, that is, $[f(\Theta)]_{i,j}=f(\Theta_{i,j})$ for all $i,j$. The choice of $f$ depends on the type of data at hand. For example, one might choose $f(x)=\operatorname{sign}(x)$ for binary data in $\{-1,1\}$ [11], or if $X$ has a multiplicative structure, one can consider the Hadamard product decomposition with $f(x)=x^{2}$, which is also efficient for the compression of dense images  [9, 18] and for representing probabilistic circuits [27]. In this work, we assume $X$ to be nonnegative and sparse, opting for $f$ to be the ReLU function, $f(\cdot)=\max(0,\cdot)$. This NMD decomposition, with $X\approx\max(0,\Theta)$, is referred to as the ReLU decomposition. Intuitively, the ReLU decomposition aims at finding a matrix $\Theta$ that substitutes the zero entries of the target matrix $X$ with negative values in order to decrease its rank. Minimizing the least squares error leads to an optimization problem referred to as ReLU-NMD:
(1) $\min_{\Theta}\lVert X-\max(0,\Theta)\rVert_{F}^{2}\quad\mbox{such that }\quad\operatorname{rank}(\Theta)\leq r.$
Note that some degree of sparsity on the matrix $X$ is an essential condition for ReLU-NMD to be meaningful in practice. In fact, if $X$ has mostly positive entries, then $\Theta$ will also have mostly positive entries and the ReLU nonlinear activation will be active only for few entries, leading to a model that is similar to standard low-rank matrix approximations such as the truncated singular value decomposition (TSVD). However, it is interesting to note that ReLU-NMD will always lead to a lower error than the TSVD. In fact, let $X_{r}$ be an optimal rank-$r$ approximation of $X\geq 0$ (e.g., obtained with the TSVD), then $\|X-X_{r}\|_{F}^{2}\geq\|X-\max(X_{r},0)\|_{F}^{2}$ since negative entries in $X_{r}$, if any, are replaced by zeros which better approximate the corresponding nonnegative entries in $X$.
ReLU-NMD was introduced by Saul in [34, 33]. It is non-differentiable and non-convex, and hence difficult to tackle. Indeed, to the best of our knowledge, there exists only one algorithm that addresses (1) directly, by parametrizing $\Theta=WH$ and using a coordinate descent approach, optimizing one entry of $W$ and $H$ at a time; see [4]. All other existing approaches rely on Saul’s alternative formulation, dubbed Latent-ReLU-NMD, and defined as
(2) $\min_{Z,\Theta}\lVert Z-\Theta\rVert_{F}^{2}\quad\mbox{such that}\quad\begin{cases}\mbox{rank}(\Theta)\leq r,\\ \max(0,Z)=X.\end{cases}$
The objective function in (2) is differentiable, while the subproblems in variables $Z$ and $\Theta$ can be solved up to global optimality; see Section 4. However, for a general matrix $X$, the link between the solutions of (1) and (2) has not been explored in the literature. Our first contribution, in Section 3, is to understand better the connection between the two.
A third formulation was proposed in [35] using the parametrization $\Theta=WH$, where $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$, which is common practice for low-rank matrix approximation, as the rank of $\Theta$ does not exceed $r$ [43, 38]. This formulation, referred to as 3B-ReLU-NMD model, is
(3) $\min_{Z,W,H}\lVert Z-WH\rVert_{F}^{2}\quad\mbox{such that}\quad\max(0,Z)=X,$
where 3B stands for 3 blocks of variables. The objective function in (3) is again jointly nonconvex, but it is convex with respect to each matrix variable, $W$, $H$ and $Z$, and can be solved computationally faster, as the rank constraint is not explicitly involved. Problem (3) can be tackled using the block coordinate descent (BCD) method, optimizing each matrix variable sequentially [35, 41].
Outline and contribution
Section 2 motivates the study of ReLU-NMD by presenting several applications. In Section 3, we clarify the relationship between ReLU-NMD and Latent ReLU-NMD. In particular, we provide a matrix $X$ for which the infimum of the rank-1 Latent-ReLU-NMD is not attained, showing the ill-poseness of Latent-ReLU-NMD (Lemma 3.1). On the other hand, for the same matrix $X$, ReLU-NMD has a non-empty set of optimal rank-1 solutions (Lemma 3.3). This shows that, in general, the two models do not share the same set of optimal low-rank solutions, that is, the sets
[…]
do not necessarily coincide. However, we show that any feasible point $(\widehat{Z},\widehat{\Theta})$ of the Latent-ReLU-NMD model satisfies the relation
$\lVert X-\max(0,\widehat{\Theta})\rVert_{F}\leq 2\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}.$
This guarantees that any approximated solution of Latent-ReLU-NMD provides an approximated solution for ReLU-NMD whose residual is at most twice that of Latent-ReLU-NMD.
In Section 4, we first briefly review the state-of-the-art algorithms for solving Latent-ReLU-NMD and 3B-ReLU-NMD (Section 4.1). We then prove convergence of the BCD scheme for solving 3B-ReLU-NMD (BCD-NMD) (Section 4.2), demonstrating that it falls within the framework of BCD with strictly quasi-convex subproblems [16]. In addition, we introduce a novel extrapolated version of BCD-NMD (eBCD-NMD) (Section 4.3), adapting the well-known LMaFit algorithm for matrix completion [43]. We prove subsequence convergence of eBCD-NMD (Theorem 4.9). To the best of our knowledge, BCD-NMD and eBCD-NMD are the first algorithms for solving 3B-ReLU-NMD with convergence guarantees.
In Section 5, we provide extensive numerical experiments comparing our new approach, eBCD-NMD, with the state of the art. We test the algorithms on various applications: matrix completion with ReLU sampling, Euclidean distance matrix completion (EDMC), compression of sparse data, and low-dimensional embedding. In particular, our experiments show that eBCD-NMD consistently accelerates BCD-NMD and performs favorably with the state of the art, while having stronger theoretical guarantees.
2 Applications of ReLU-NMD
In this section, we first show the link between the ReLU decomposition and neural networks. Then we present some applications of the ReLU decomposition, illustrating how nonlinear models can be used in a variety of different contexts.
2.1 Neural network interpretation
The increasing interest in nonlinear models can be attributed to the rising popularity of neural networks, which use nonlinear functions as activations in their hidden layers. Furthermore, low-rank nonlinear models, involving matrix decompositions, have been studied in the context of neural networks [11, 23, 28, 36]. Let us discuss the link between ReLU-NMD and ReLU networks. Figure 1 illustrates a one hidden-layer neural network.
$g$$f$$W$$Y$$a_{i}$$b_{i}$
Figure 1: One hidden-layer neural network with 5 input and output nodes and 3 hidden nodes.
Let $a_{i}\in\mathbb{R}^{p}$ be $p$-dimensional input and $b_{i}\in\mathbb{R}^{m}$ the corresponding output vectors, for $i=1,\dots,n$. Define the input and output matrices $A=[a_{1},\dots,a_{n}]\in\mathbb{R}^{p\times n}$ and $B=[b_{1},\dots,b_{n}]\in\mathbb{R}^{m\times n}$, respectively. We also denote $Y\in\mathbb{R}^{r\times p}$ the matrix of the weights between the input layer and hidden layer, and $W\in\mathbb{R}^{m\times r}$ the matrix of the weights between the hidden layer and the output layer. The functions $g$ and $f$ are the activation functions for the first and second layer, respectively. We are interested in finding an approximation of a given data $X$ using this neural network with one hidden layer. If we choose the error measure to be the least squares error, we need to minimize
(4) $\lVert X-B\rVert_{F}^{2}\;=\;\lVert X-f(Wg(YA))\rVert_{F}^{2}.$
This shows that the hidden representation of each data point is a nonlinear combination of the inputs of the network, that is, $h_{i}=g(Ya_{i})\in\mathbb{R}^{r}$, or $H=g(YA)\in\mathbb{R}^{r\times n}$ in matrix form. NMD is a relaxation of (4) in which the hidden representation of the input matrix, $H$, can be any real matrix of appropriate dimensions. Furthermore, the number of hidden neurons is equal to the rank of the decomposition, and it needs to be chosen a priori as when designing the structure of the network. If the number of hidden neurons is smaller than the dimension of the inputs, the hidden layer is a low-dimensional embedding of the data.
2.2 Compression of sparse data
An intuitive application of ReLU decomposition models is the compression of sparse data. Let $X\in\mathbb{R}^{m\times n}$ be a sparse and nonnegative matrix with $\operatorname{nnz}(X)$ non-zero entries, and let $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$ be such that $X\approx\max(0,WH)$. As long as $r(m+n)<\operatorname{nnz}(X)$, storing the factors $W$ and $H$ requires less memory than storing $X$. ReLU models can outperform the standard linear compression techniques such as the TSVD on sparse data, such as sparse images or dictionaries [34, 33, 35, 41, 4]. An interesting example is the identity matrix of any dimension that can be exactly reconstructed using a rank-3 ReLU decomposition [11, 34].
2.3 Matrix completion with ReLU sampling
The class of matrix completion problems with entries missing not at random (MNR) is characterized by the fact that the probability of an entry being missing depends on the matrix itself. For example, when completing a survey, people are likely to share nonsensitive information such as the name or the surname, while they might be more hesitant to disclose their phone number. Therefore, we might expect the missing entries to be concentrated in some area of the matrix more than others. Even though MNR matrix completion is still a relatively unexplored area, dropping the assumption that missing entries are independent of the matrix is a challenging problem that attracts more and more people as the increasing number of related works testifies [6, 29, 13, 7, 26]. Among them, Liu et al. [26] investigate the matrix completion problem when only positive values are observed, matrix completion with ReLU sampling. This instance of matrix completion falls within the MNR problems since the missing entries depend on the sign of the target matrix. Assume that $X$ is a nonnegative and sparse matrix and that there exists a rank-$r$ matrix $\Theta$ such that $X=\max(0,\Theta)$. In this case, finding a ReLU decomposition is equivalent to recovering the negative values, observing only the positive ones; thus, it is equivalent to matrix completion with ReLU sampling. This implies that we can use the algorithms for computing a ReLU decomposition to address this matrix completion problem, as illustrated in [26].
Inspired by the connection with ReLU sampling matrix completion, we introduce an equivalent formulation of 3B-ReLU-NMD that we will use throughout the paper. Let $\Omega=\{(i,j)\ |\ X_{ij}>0\}$ be the set of positive entries of $X$ and let $P_{\Omega}(X)$ be equal to $X$ in $\Omega$ and zero elsewhere. We rewrite 3B-ReLU-NMD in (3) as
(5) $\operatorname{argmin}_{Z,W,H}\frac{1}{2}\lVert Z-WH\rVert_{F}^{2}\quad\mbox{such that}\quad P_{\Omega}(Z)=P_{\Omega}(X),\quad P_{\Omega^{C}}(Z)\leq 0,$
where $\Omega^{C}$ is the complement of $\Omega$ and a reformulation of the constraint $\max(0,Z)=X$ is introduced. Using the matrix factorization $\Theta=WH$ is a standard approach in matrix completion [5, 19, 20, 21, 38], see [30, 25] for detailed surveys on the topic. To our knowledge, the problem (5), without the constraint $P_{\Omega^{C}}(Z)\leq 0$, was first addressed in [43] in the context of matrix completion. The novelties in our model are the additional information that the missing entries have negative signs, which translates into the constraints $P_{\Omega^{C}}(Z)\leq 0$, and the explicit dependence of the known index set $\Omega$ from the target matrix $X$. Since the formulations (3) and (5) are equivalent, we will refer to both as 3B-ReLU-NMD.
2.4 Euclidean distance matrix completion (EDMC)
We provide now one practical example of the MNR matrix completion problem that can be solved using the ReLU decomposition. It is the recovery of a low-rank matrix for which only the smallest or the largest entries are observed. Let $\Theta$ be a low-rank matrix, and assume that we observe all the entries in $\Theta$ smaller than a known threshold $d$. We want to recover the missing entries, which are the ones larger than $d$. We can model this problem using a rank-1 modified ReLU decomposition: $X=\max(0,dee^{T}-\Theta)$, where $e$ denotes the vector of all ones of appropriate dimension. The matrix $X$ is nonzero exactly in the entries where $\Theta$ is smaller than $d$. Of course, an analogous model can be derived when only the largest entries are observed. The 3B-ReLU-NMD formulation in (5) can be easily adapted to include the rank-1 modification, that is,
(6) $\operatorname{argmin}_{Z,W,H}\frac{1}{2}\lVert dee^{T}-WH-Z\rVert_{F}^{2}\quad\mbox{such that}\quad P_{\Omega}(Z)=P_{\Omega}(X),\quad P_{\Omega^{C}}(Z)\leq 0.$
The same idea applies also to the Latent-ReLU-NMD and all the algorithms can be adjusted to take into account the additional rank-one term. This formulation is particularly interesting when the matrix $\Theta$ contains the squared distances between a collection of points $p_{1},\dots,p_{n}\in\mathbb{R}^{d}$. This problem is often referred to as Euclidean distance matrix completion (EDMC) [1, 12, 22, 39], and is particularly meaningful in sensor network localization problems where sensors can only communicate with nearby sensors. Let $P=[p_{1},\dots,p_{n}]\in\mathbb{R}^{d\times n}$, then the corresponding Euclidean distance matrix is $\Theta_{i,j}=\|p_{i}-p_{j}\|_{2}^{2}$ for all $i,j$, so that $\Theta=e\text{diag}(P^{T}P)^{T}+\text{diag}(P^{T}P)e^{T}+2P^{T}P$ and it has at most rank $r=d+2$ and thus the rank of the factorization in (6) is known in advance. We will show in Section 5 that this slightly modifed ReLU-NMD model allows us to solve EDMC in this context.
2.5 Manifold learning
Manifold learning aims at learning a norm-preserving and neighborhood-preserving mapping of high-dimensional inputs into a lower-dimensional space [24]. Given a collection of points $\{z_{1},\dots,z_{m}\}$ in $\mathbb{R}^{N}$, Saul [33] looks for a faithful representation $\{y_{1},\dots,y_{m}\}$ in $\mathbb{R}^{r}$, with $r<N$. Let $\tau\in(0,1)$, Saul defines an embedding to be $\tau$-faithful if
(7) $\max(0,\langle y_{i},y_{j}\rangle-\tau\lVert y_{i}\rVert\lVert y_{j}\rVert)=\max(0,\langle z_{i},z_{j}\rangle-\tau\lVert z_{i}\rVert\lVert z_{j}\rVert).$
Equation (7) implies that the embedding must preserve norms (take $i=j$ to obtain $(1-\tau)\lVert y_{i}\rVert^{2}$ $=$ $(1-\tau)\lVert z_{i}\rVert^{2}$) as well as the small angles between two points $(z_{i},z_{j})$: when $\cos(z_{i},z_{j})>\tau$, we need $\cos(z_{i},z_{j})=\cos(y_{i},y_{j})$. Large angles have to remain large but do not need to be preserved, that is, $\cos(z_{i},z_{j})\leq\tau$ only requires $\cos(y_{i},y_{j})\leq\tau$. The condition (7) defines a similarity between points based on norm and angles, which differs substantially from the more common concept of pairwise distances. This makes it suitable, for example, in text analysis where the $z_{i}$’s are sparse vectors of word counts.
This embedding strategy is dubbed threshold similarity matching (TSM). Saul proposes three main steps to compute such an embedding. First, for each couple of points $(z_{i},z_{j})$, compute the right-hand side in (7) and construct the sparse, nonnegative similarity matrix
$X_{ij}=\max(0,\langle z_{i},z_{j}\rangle-\tau\lVert z_{i}\rVert\lVert z_{j}\rVert)\quad\text{ for all }i,j.$
Second, compute a ReLU-decomposition, $X\approx\max(0,\Theta)$, and the matrix $\max(0,\Theta)$ represents the similarities between points in the lower-dimensional space. Finally, the embedded points are recovered from the low-rank approximation $\Theta$; see [33] for more details.
3 ReLU-NMD vs Latent-ReLU-NMD
In this section, we explore the link between ReLU-NMD in (1) and Latent-ReLU-NMD in (2). In particular, we aim to determine if the optimal low-rank approximations provided by the two models coincide. First, when $X$ contains negative values, Latent-ReLU-NMD is infeasible, since there exists no latent variable $Z$ such that $X=\max(0,Z)$. In contrast, ReLU-NMD still admits feasible solutions. Of course, the ReLU decomposition is more meaningful when $X$ is nonnegative which is the case we focus on in the following.
In the exact case, when there exists a low-rank matrix $\Theta$ such that $X=\max(0,\Theta)$, ReLU-NMD and Latent-ReLU-NMD are equivalent: for any low-rank solution $\Theta^{*}$ of (1), there exists a latent matrix $Z^{*}$, such that $(Z^{*},\Theta^{*})$ is a solution of (2) and vice versa. Indeed, in the exact case there always exists a trivial latent solution $Z^{*}=\Theta^{*}$. We show in what follows that this is not true in general, when an exact decomposition does not exist.
We demonstrate first that there exists a specific matrix $X$ such that the rank-1 Latent-ReLU-NMD is ill-posed since the infimum of (2) is not attained (see Lemma 3.1). Moreover, by showing that for the same matrix $X$, ReLU-NMD admits a non-empty set of optimal solutions (Lemma 3.3), we establish that the two formulations do not necessarily yield the same optimal approximations. Note that we could not establish the general well-poseness of ReLU-NMD, which is a direction of future research.
Lemma 3.1.
Let $0<\epsilon<1/\sqrt{2}$ be a fixed parameter, $r=1$, and
(8) $X=\begin{pmatrix}1&0\\ \epsilon&1\end{pmatrix}.$
Then the infimum of Latent-ReLU-NMD in (2) is not attained.
Proof 3.2.
We need to show that there exists no solution of Latent-ReLU-NMD attaining the infimum of (2). Let us prove the result by contradiction, and assume there exists a solution to (2), denoted $(\widehat{Z},\widehat{\Theta})$. First note that, given $\widehat{Z}$, the optimal $\widehat{\Theta}$ must be a best rank-1 approximation of $\widehat{Z}$. Hence $\widehat{Z}$ is such that the distance with its best rank-$1$ approximation is minimized and such that it satisfies the constraint $\max(0,\widehat{Z})=X$. Let us denote the $i$-th singular values of $Z$ as $\sigma_{i}(Z)$, and $\widehat{Z}_{1}$ be a best rank-one approximation of $\widehat{Z}$. By the Eckart-Young theorem, $\|\widehat{Z}-\widehat{Z}_{1}\|_{F}^{2}=\sigma_{2}(Z)^{2}$ since $\widehat{Z}$ is a 2-by-2 matrix. Hence $\widehat{Z}$ must belong to
(9) $\operatorname{argmin}_{Z}\sigma_{2}(Z)^{2}\quad\mbox{such that }\quad\max(0,Z)=X.$
Equivalently, we want to find $b\leq 0$ such that the matrix $Z=\begin{pmatrix}1&b\\ \epsilon&1\end{pmatrix}$ has the lowest possible second singular value. Since $Z$ is a $2\times 2$ matrix, we can compute the eigenvalues of $ZZ^{T}$, solve det$(ZZ^{T}-\sigma I)=0$ and get an explicit expression for the second singular value of $Z$: for a general matrix $\begin{pmatrix}a&b\\ c&d\end{pmatrix},$ the second singular value is given by
$\sigma_{2}=\sqrt{\frac{S_{1}-S_{2}}{2}},$
where
$S_{1}=a^{2}+b^{2}+c^{2}+d^{2},\quad S_{2}=\sqrt{(a^{2}+b^{2}-c^{2}-d^{2})^{2}+4(ac+bd)^{2}}.$
We rewrite (9) as
(10) $\operatorname{argmin}_{b<0}\;\ell(b):=\frac{2+b^{2}+\epsilon^{2}-\sqrt{(b^{2}-\epsilon^{2})^{2}+4(b+\epsilon)^{2}}}{2}.$
One can verify that $\lim_{b\rightarrow-\infty}\ell(b)=\epsilon^{2}$, while $\ell(b)>\epsilon^{2}$ for every $\epsilon\leq 1/\sqrt{2}$ and $b<0$. This means the infimum of $\ell(b)$ is not attained by any finite value of $b$, and hence $\widehat{Z}$ does not exist.
We demonstrate in the following lemma that rank-1 ReLU-NMD admits an optimal solution for $X$ as in (8).
Lemma 3.3.
Let $X$ as in (8) and $r=1$, then for every $0<\epsilon<\frac{1}{\sqrt{2}}$, the infimum of ReLU-NMD is attained by all the matrices of the form
(11) $\Theta=\begin{pmatrix}1&-v\\ -\frac{1}{v}&1\end{pmatrix},\quad v\in\mathbb{R_{+}},$
with optimal value equal to $\epsilon^{2}.$
Proof 3.4.
Let us denote the error function of ReLU-NMD as $F(\Theta)=\lVert X-\max(0,\Theta)\rVert_{F}^{2}$. Note that $2\times 2$ rank-1 matrices have limited patterns that the signs of their entries must follow. For instance, if the matrix does not contain zeros, the number of positive and negative entries must be even, otherwise the rank cannot be equal to one. We proceed by analyzing three different cases.
Case 1: The rank-1 matrix $\Theta$ contains one zero column or row. In this case, the residual $F(\Theta)$ cannot be less than 1 since the nonlinear approximation $\max(0,\Theta)$ has at least one zero in the diagonal.
Case 2: The rank-1 matrix $\Theta$ has no zero rows or columns, and it contains two negative entries. If there is a negative entry on the diagonal, then the error is larger than one. Otherwise, the positive entries on the diagonal can be set to 1 to minimize the error, which is exactly the matrix in (11), with error $\epsilon^{2}$.
Case 3: The rank-1 matrix $\Theta$ contains only positive entries. If $\Theta\geq 0$, ReLU-NMD corresponds to the following problem without nonlinear activations:
$\operatorname{argmin}_{\Theta}\lVert X-\Theta\rVert_{F}^{2}\ \text{ such that }\ \Theta\geq 0.$
Therefore, we need to show that there exists no rank-1 matrix with nonnegative entries that achieves a residual value $F(\Theta)$ strictly smaller than $\epsilon^{2}$. Let $\widehat{\sigma}_{2}$ be the second singular value of $X$ as in (8). If rank($\Theta$)=1, then $\lVert X-\Theta\rVert_{F}^{2}\geq\widehat{\sigma}_{2}^{2}$ and using equation (10),
$\widehat{\sigma}_{2}^{2}=\frac{2+\epsilon^{2}-\epsilon\sqrt{\epsilon^{2}+4}}{2}.$
Simple calculations show that $\widehat{\sigma}_{2}^{2}$ is strictly larger than $\epsilon^{2}$ for any $0<\epsilon<\frac{1}{\sqrt{2}}$.
As a consequence of Lemma 3.1 and Lemma 3.3, we have the following corollary.
Corollary 3.5.
ReLU-NMD and Latent-ReLU-NMD may not share the same set of optimal low-rank solutions.
We established the non equivalence between the low-rank solutions of ReLU-NMD and Latent-ReLU-NMD. However, in practice, we are interested in computing accurate solutions of ReLU-NMD even though they might not be optimal. Numerous numerical results show that the ReLU decomposition obtained by Latent-ReLU-NMD achieves a small residual error also for ReLU-NMD [34, 35, 41]. The following theorem formalizes this observation, showing that any feasible point $(\widehat{\Theta},\widehat{Z})$ of the Latent model has a residual value for $\widehat{\Theta}$ in the original ReLU-NMD that is upper bounded by twice the residual of Latent-ReLU-NMD.
Theorem 3.6.
Let $X\in\mathbb{R}^{n\times m}$ be a nonnegative matrix and denote as $(\widehat{\Theta},\widehat{Z})$ a feasible point of the Latent-ReLU-NMD model in (2), then
(12) $\lVert X-\max(0,\widehat{\Theta})\rVert_{F}^{2}\leq 2\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}.$
Proof 3.7.
We have
(13) $\lVert X-\max(0,\widehat{\Theta})\rVert_{F}^{2}$ $=\lVert X-\widehat{Z}+\widehat{Z}-\max(0,\widehat{\Theta})-\min(0,\widehat{\Theta})+\min(0,\widehat{\Theta})\rVert_{F}^{2}$
$\leq\lVert X-\widehat{Z}+\min(0,\widehat{\Theta})\rVert_{F}^{2}+\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}.$
Let us define $\Omega=\{(i,j)\ |\ X(i,j)>0\}$ and $\widehat{\Omega}=\{(i,j)\ |\ \widehat{\Theta}(i,j)>0\}$, and $\Omega^{C}$ and $\widehat{\Omega}^{C}$ their respective complements. By assumption $\widehat{Z}$ is a feasible solution of (2) and hence $\max(0,\widehat{Z})=X$, therefore
$\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}=$ $\sum_{(i,j)\in\Omega\cap\widehat{\Omega}}(X_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}(X_{ij}-\widehat{\Theta}_{ij})^{2}+$
$+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}.$
Similarly, observe that $Z_{i,j}\leq 0$ whenever $X_{i,j}=0$, and hence
(14) $\lVert X-\widehat{Z}+\min(0,\widehat{\Theta})\rVert_{F}^{2}=$
$=\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}\widehat{\Theta}_{ij}^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}\widehat{Z}_{ij}^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}$
$\leq\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}(X_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}$
$\leq\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}.$
The first inequality holds because
$\widehat{\Theta}_{ij}^{2}\leq(X_{ij}-\widehat{\Theta}_{ij})^{2}\mbox{ for }(i,j)\in\Omega\cap\widehat{\Omega}^{C},\text{ and }$
$\widehat{Z}_{ij}^{2}\leq(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}\mbox{ for }(i,j)\in\Omega^{C}\cap\widehat{\Omega}.$
Combining (13) and (14), we get (12).
A straightforward consequence of Theorem 3.6 is that a solution of Latent-ReLU-NMD with small error provides a ReLU-NMD solution with small error, even though the two models do not necessarily yield the same approximation.
4 Algorithms for the ReLU decomposition
This section is dedicated to the algorithms for computing ReLU decompositions and to the study of their convergence. In Section 4.1, we briefly review the state-of-the-art algorithms for solving Latent-ReLU-NMD and 3B-ReLU-NMD. In Section 4.2, we discuss the convergence of the BCD-NMD scheme. Finally, we present our new algorithm eBCD-NMD along with its convergence in Section 4.3.
4.1 Previous works
Let us start with Latent-ReLU-NMD, since we use the algorithms designed to solve Latent-ReLU-NMD as baselines in the numerical comparison in Section 5. To our knowledge, all existing algorithms for Latent-ReLU-NMD are alternate optimization methods in which each block of variables is updated sequentially [16, 17, 32, 31, 40] and lack convergence guarantees.
The first approach is the so-called Naive-ReLU-NMD, introduced by Saul in [34], which employs a basic alternating scheme over $Z$ and $\Theta$ with closed-form solutions for both. Similar approaches have been studied in matrix completion, see for example [5, 8, 10, 15]. At each iteration, firstly $\Theta$ is fixed and then the optimal $Z$ is computed solving the subproblem
(15) $\operatorname{argmin}_{Z}\lVert Z-\Theta\rVert_{F}^{2}\quad\mbox{such that}\quad\max(0,Z)=X.$
The solution of (15) is given by the projection of $\Theta$ over the set $\{Z\ :\ \max(0,Z)=X\}$, that is,
$Z_{ij}=\begin{cases}X_{ij}\quad&\mbox{if }X_{ij}>0\\ \min(0,\Theta_{ij})&\mbox{if }X_{ij}=0.\\ \end{cases}$
Once $Z$ is updated, $\Theta$ is updated solving
$\operatorname{argmin}_{\Theta}\lVert Z-\Theta\rVert_{F}^{2}\quad\mbox{such that}\quad\operatorname{rank}(\Theta)\leq r,$
whose optimal solution(s) can be obtained with the TSVD of $Z$. Despite its simplicity, this scheme is highly reliable in practice, with the most costly operation at each iteration being the computation of a TSVD.
An accelerated, heuristic version of the Naive-NMD scheme was proposed in [35]. Specifically, the so-called Aggressive-ReLU-NMD (A-NMD) employs an extrapolation step for both $Z$ and $\Theta$ while the extrapolation parameter changes adaptively at each iteration, depending on the value of the residual. A-NMD is an example where, despite its lack of theoretical guarantees, extrapolation can be considerably faster [2, 43].
Another algorithm for solving Latent-ReLU-NMD is the so-called expectation-minimization NMD (EM-NMD) [34]. This approach is based on a Gaussian latent variable model parametrized by the low-rank matrix $\Theta$ and a parameter $\sigma$. For each entry of the original matrix, a Gaussian latent variable $Z_{ij}=\mathcal{N}(\Theta_{ij},\sigma^{2})$ is sampled. The method estimates $\Theta$ and $\sigma$ by maximizing the likelihood of the data with respect to $\Theta$ and $\sigma$. The optimization is carried out using the Expectation-Minimization (EM-NMD) scheme that consists of two main steps: the E-step which computes posterior mean and variance of the Gaussian latent variable model, and the M-step re-estimates the parameters of the model from the posterior mean and variance.
The BCD scheme for solving 3B-ReLU-NMD in (5), referred to as BCD-NMD, consists of an alternating optimization procedure over three blocks of variables $(Z,W,H)$. BCD-NMD belongs to the class of exact BCD methods, where a global optimizer for each of the subproblem is computed sequentially. Let $A^{\dagger}$ denote the Moore-Penrose inverse of $A$ and let $(H^{k},Z^{k},W^{k})$ be the current iteration. The $(k+1)$-th step of BCD-NMD is as follows:
(16) $Z^{k+1}=$ $\operatorname{argmin}_{Z}\{\lVert Z-W^{k}H^{k}\rVert_{F}^{2}\quad\text{s.t.}\quad\max(0,Z)=X\}=P_{\Omega}(X)+P_{\Omega^{C}}(\min(0,W^{k}H^{k})),$
$W^{k+1}=$ $\operatorname{argmin}_{W}\lVert Z^{k+1}-WH^{k}\rVert_{F}^{2}=Z^{k+1}(H^{k})^{\dagger},$
$H^{k+1}=$ $\operatorname{argmin}_{H}\lVert Z^{k+1}-W^{k+1}H\rVert_{F}^{2}=((W^{k+1})^{T})^{\dagger}Z^{k+1}.$
The solution of the subproblems for $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$ are analogous. In particular, the first is separable by rows of $W$, leading to $m$ least squares problems in $r$ variables with the data vector of dimension $n$, while the subproblem for $H$ is separable by column and equivalent to $n$ least squares problems of $r$ variables and data vector of dimension $m$. Both these subproblems do not necessarily have a unique solution. In fact, uniqueness is guaranteed only if $W^{k}$ and $H^{k}$ remain full rank throughout the iterations. Alternatively, uniqueness is ensured by adding a Tikhonov regularization term or a proximal term in each of the subproblems [16, 17, 3]; however, it is not necessary to prove the convergence of the BCD-NMD algorithm.
On the contrary, the subproblem in the latent variable $Z$ admits a unique global minimum, which is the projection of the product $W^{k}H^{k}$ onto the feasible, convex set
(17) $\{Z\in\mathbb{R}^{m\times n}\ :\ P_{\Omega}(Z)=P_{\Omega}(X),\ P_{\Omega^{C}}(Z)\leq 0\}.$
The overall BCD-NMD algorithm is reported in Algorithm 1 and has a computational cost of $O(r|\Omega^{C}|+(m+n)r^{2})$.
Algorithm 1 BCD-NMD: block coordinate descent for ReLU-NMD
0:  $X$, $W^{0}$, $H^{0}$, maxit.
0:  low-rank factors $W^{k}$ and $H^{k}$.
1:  Set $\Omega=\{(i,j)\ |\ X_{ij}>0\}$
2:  for $k=0,1,\dots,$ maxit do
3:     $Z^{k+1}=P_{\Omega}(X)+P_{\Omega^{C}}(\min(0,W^{k}H^{k}))$
4:     $W^{k+1}=Z^{k+1}(H^{k})^{\dagger}$
5:     $H^{k+1}=((W^{k+1})^{T})^{\dagger}Z^{k+1}$
6:  end for
Similarly to the Naive scheme for Latent-ReLU-NMD, adding extrapolation might be beneficial also for BCD-NMD. In particular, the work [35] proposes to perform an extrapolation step on the latent variable $Z$ and on the low-rank product $WH$, resulting in the so-called 3B-NMD algorithm. Moreover, Wang et al. [41, 42] added a Tikhonov regularization for both the variables $W$ and $H$ in the 3B-ReLU-NMD formulation. Their algorithm is dubbed as ReLU-NMD-T and it is based on the 3B-NMD scheme but performs two more extrapolation steps following the computation of $W$ and $H$. Even though numerical experience shows that both the accelerated algorithms outperform the original BCD-NMD, convergence still remains an open problem. For the numerical experiments in Section 5, we consider only 3B-NMD since we are interested in comparing algorithms that solve the non-regularized 3B-ReLU-NMD.
4.2 Convergence of BCD-NMD
In this section, we discuss the convergence to a stationary point of BCD-NMD (Algorithm 1) which, to the best of our knowledge, has not been studied before. Indeed, our goal is to prove that BCD-NMD is convergent under the assumption that a limit point of the sequence generated by the scheme exists. We proceed by showing that the method falls into the more general framework of exact block coordinate descent (BCD), where one of the subproblems is strictly quasi-convex.
Specifically, the general BCD scheme is studied by Grippo et al. [16] when the objective function is continuously differentiable over the Cartesian product of $m$ closed, convex sets, one for each block of variables. We first observe that the 3B-ReLU-NMD objective function is continuously differentiable and defined over three blocks of variables $Z$, $W$, and $H$. The latent variable $Z$ is the only constrained variable, and it is easy to show that the constraint set in (17) is convex. Therefore, the 3B-ReLU-NMD is a particular instance of the more general problem considered in [16].
Theorem 4.1.
Any limit point of the sequence generated by BCD-NMD (Algorithm 1) is a stationary point of problem (5).
Proof 4.2.
The global convergence of the BCD scheme for a general function with $m$ blocks of variables is analyzed under generalized convexity assumptions in [16] (Proposition 5). The key hypotheses are: 1) the existence of a limit point of the sequence generated by the BCD algorithm, and 2) strict quasi-convexity of at least $m-2$ subproblems. Therefore, since 3B-ReLU-NMD has three blocks of variables, we just need one subproblem to be strictly quasi-convex. The first subproblem in the variable $Z$ is a convex, quadratic problem over a convex set, making it strictly quasi-convex, implying the convergence of BCD-NMD.
The assumption of the existence of the limit point is an undesirable hypothesis. However, it is necessary because the level sets of the objective function are not bounded. Therefore, similarly to Latent-ReLU-NMD, 3B-ReLU-NMD might have instances in which the minimum is not attained (see Theorem 3.3). This means that, in pathological cases, the sequence generated by BCD-NMD may diverge simply because the infimum of the function is not attained by any finite value. Thus, the hypothesis regarding the existence of a limit point of the BCD-NMD sequence cannot be weakened. One possible way to overcome this limitation is by adding a regularization term to the objective function (5), which makes the level sets compact, as in [41]. However, this is not the purpose of this work which focuses on the non-regularized formulation.
4.3 Extrapolated block coordinate descent NMD (eBCD-NMD)
We present now a new extrapolated version of Algorithm 1, namely extrapolated BCD-NMD (eBCD-NMD). Our goal is to accelerate BCD-NMD while preserving convergence to a stationary point under mild assumptions. The eBCD-NMD method is an adaptation of the well-known LMaFit algorithm [43] for matrix completion. The standard matrix completion formulation in [43] is similar to the one we consider in (5), although it does not include the inequality constraint $P_{\Omega^{C}}(Z)\leq 0$, which models the crucial information that the missing entries are negative. Consequently, we derive a new optimization scheme, differing from the one in [43], as the update of $Z$ presents an additional nonlinear term; see equation (18). Moreover, we relax the hypothesis in [43] that require $\{W^{k}\}$ and $\{H^{k}\}$ to remain full-rank for all iterations. We present two different versions of the eBCD-NMD algorithm that, starting from the same point, produce the same low-rank approximation $W^{k}H^{k}$, yet the factors $W^{k}$ and $H^{k}$ vary. In what follows, we omit the iteration index, $k$, to simplify the notation, as we focus on a single iteration eBCD-NMD. We assume we have a point $(Z,W,H)$ from which we perform one iteration.
4.3.1 First version of eBCD-NMD
Let $\alpha\geq 1$, then the first version of the eBCD-NMD scheme uses the following update: given the iterate $(Z,W,H)$, it computes
(18) $Z_{\alpha}=\alpha Z+(1-\alpha)WH,$
$\widehat{W}(\alpha)=Z_{\alpha}H^{\dagger},$
$\widehat{H}(\alpha)=(\widehat{W}(\alpha)^{T})^{\dagger}Z_{\alpha},$
$\widehat{Z}(\alpha)=P_{\Omega}(X)+P_{\Omega^{C}}(\min(0,\widehat{W}(\alpha)\widehat{H}(\alpha))),$
where $\left(\widehat{Z}(\alpha),\widehat{W}(\alpha),\widehat{H}(\alpha)\right)$ provides the next iterate. Similarly to BCD-NMD, the main computational cost for the scheme in (18) is solving the two least squares problems for computing $\widehat{W}(\alpha)$ and $\widehat{H}(\alpha)$. The optimization scheme in (18) differs from the one in (16) because of the new variable $Z_{\alpha}$ which is used in place of the original latent variable $Z$ to update $W$ and $H$. Moreover, if we assume the matrix $H$ to be full-rank, we can derive an intuitive interpretation of the update of $Z_{\alpha}$ as an indirect extrapolation step. Since $H$ is full-rank we have $H^{\dagger}=H^{T}(HH^{T})^{-1}$ and thus
$\widehat{W}(\alpha)=Z_{\alpha}H^{T}(HH^{T})^{-1}=\alpha ZH^{T}(HH^{T})^{-1}+(1-\alpha)WHH^{T}(HH^{T})^{-1}=\alpha W_{\text{BCD}}+(1-\alpha)W,$
where $W_{\text{BCD}}$ is the update of the variable $W$ that we would have had if we performed one step of the BCD-NMD scheme (16). Moreover, if we define $\beta=\alpha-1$, then
(19) $\widehat{W}(\alpha)=\alpha W_{\text{BCD}}+(1-\alpha)W=(1+\beta)W_{\text{BCD}}-\beta W=W_{\text{BCD}}+\beta(W_{\text{BCD}}-W),$
that is, an extrapolation step for the BCD-NMD update of $W$, provided that $\alpha\in(1,2)$, that means $\beta\in(0,1)$. Unfortunately, such an intuitive interpretation does not apply to the update of the variable $H$. Note that the step in (19) is properly referred to as an extrapolation step if $\beta$ is in the interval (0,1) or equivalently $\alpha$ is in the interval (1,2), but our scheme potentially allows for larger values that exceed the upper bound.
4.3.2 Improved version of eBCD-NMD
The second version of the eBCD-NMD does not require solving any least squares problem and the main computational cost is associated with a single QR factorization of a matrix in $\mathbb{R}^{m\times r}$. We specify that this scheme can be adapted to design alternative versions of BCD-NMD and 3B-NMD as well. We state the following lemma from [43] that allows us to derive the second version of eBCD-NMD.
Lemma 4.3.
Let $\widehat{W}(\alpha)$ be the update of $W$ for the scheme in (18), then
$\mathcal{R}(\widehat{W}(\alpha))=\mathcal{R}(Z_{\alpha}H^{T}),$
where $\mathcal{R}(A)$ denotes the range of matrix $A$.
Proof 4.4.
Let $UDV^{T}=H$ be the economy form SVD of $H$, then
$\widehat{W}(\alpha)=Z_{\alpha}H^{\dagger}=Z_{\alpha}VD^{\dagger}U^{T}.$
Moreover, since $Z_{\alpha}H^{T}=Z_{\alpha}VDU^{T}$, we get $\mathcal{R}(\widehat{W}(\alpha))=\mathcal{R}(Z_{\alpha}H^{T})$.
Lemma 4.3 shows that the range space of $\widehat{W}(\alpha)$ can be obtained from that of $Z_{\alpha}H^{T}$. Let $Q(\alpha)$ be a matrix whose columns contain an orthogonal basis for $\mathcal{R}(Z_{\alpha}H^{T}).$ Then, using Lemma 4.3 and from the second and third lines in (18), we get
(20) $\widehat{W}(\alpha)\widehat{H}(\alpha)=\widehat{W}(\alpha)\widehat{W}(\alpha)^{\dagger}Z_{\alpha}=Q(\alpha)Q(\alpha)^{T}Z_{\alpha}.$
This shows that the low-rank approximation produced by the eBCD-NMD is nothing else than the orthogonal projection of $Z_{\alpha}$ into $\mathcal{R}(Z_{\alpha}H^{T})$. Justified by the relation in (20), we now state the improved version of eBCD-NMD. Given an iterate $(Z,W,H)$, we compute
(21) $Z_{\alpha}=\alpha Z+(1-\alpha)WH,$
$Q(\alpha)\rightarrow\text{ Orthogonal basis of }\mathcal{R}(Z_{\alpha}H^{T}),$
$W(\alpha)=Q(\alpha),$
$H(\alpha)=W(\alpha)^{T}Z_{\alpha}=Q(\alpha)^{T}Z_{\alpha},$
$Z(\alpha)=P_{\Omega}(X)+P_{\Omega^{C}}(\min(0,W(\alpha)H(\alpha))),$
to obtain the next iterate $\left(Z(\alpha),W(\alpha),H(\alpha)\right)$. Equation (20) guarantees that the low-rank approximations computed by the scheme in (18) and in (21) are equal, that is $\widehat{W}(\alpha)\widehat{H}(\alpha)=W(\alpha)H(\alpha)$. In practice, $Q(\alpha)$ can be computed from the economy QR decomposition of $Z_{\alpha}H^{T}$. Note that if $\operatorname{rank}(Z_{\alpha}H^{T})=r_{1}<r$, the first $r_{1}$ columns of the $Q$ matrix from the economy QR decomposition of $Z_{\alpha}H^{T}$ are not, in general, an orthogonal basis of the range space of the matrix. To ensure this property, a modified version of the QR decomposition using column pivoting must be used instead. Specifically, there exists $Q(\alpha)\in\mathbb{R}^{m\times r_{1}}$, containing an orthogonal basis of $Z_{\alpha}H^{T}$, such that $Z_{\alpha}H^{T}\Pi=Q(\alpha)R(\alpha)$, where $\Pi$ is a permutation matrix and $R(\alpha)\in\mathbb{R}^{r_{1}\times r}$ [14]. Therefore, we potentially allow for a drop in the rank of the approximation whenever $Z_{\alpha}H^{T}$ is rank-deficient. However, we have never encountered this case in practical situations because $Z_{\alpha}H^{T}$ is likely to have rank at least $r$ while the factors $Q(\alpha)$ and $H(\alpha)$ have rank exactly $r$.
We proceed by defining the residual of our algorithm, denoted as $S(\alpha)$,
(22) $S(\alpha)=Z(\alpha)-W(\alpha)H(\alpha)=P_{\Omega}(X-W(\alpha)H(\alpha))-P_{\Omega^{C}}(\max(0,W(\alpha)H(\alpha))).$
We highlight an interesting relation between the variable $Z_{\alpha}$ and the residual. Let $S$ denote the residual matrix of the iterate $(Z,W,H)$, defined as in (22) with $Z$, $W$, and $H$ instead of $Z(\alpha)$, $W(\alpha)$ and $H(\alpha)$ respectively, then
(23) $Z_{\alpha}$ $=\alpha Z+(1-\alpha)WH=\alpha(P_{\Omega}(X)+P_{\Omega^{C}}(\min(0,WH))-WH)+WH$
$=\alpha(P_{\Omega}(X-WH)-P_{\Omega^{C}}(\max(0,WH)))+WH$
$=WH+\alpha S.$
Equation (23) shows that $Z_{\alpha}$ is indeed the low-rank approximation $WH$ at step $k$, to which we add a multiple of the residual $S$. This suggests a second interpretation of the update of $Z_{\alpha}$ as a residual corrected step of the low-rank approximation $WH$.
Restarting scheme of the extrapolation parameter
We need now to determine a rule to select the extrapolation parameter $\alpha$ at each iteration. Extrapolating a variable means adding information from the past iterate, thus the extrapolation parameter $\alpha$ affects how much relevant is the past information in the current update. Of course, not all the values of $\alpha$ are acceptable to progress in minimizing the residual, and a good choice of the parameter is crucial. We present now the updating scheme of the extrapolation parameter $\alpha$, inspired by the one introduced in [43]. The main idea is to choose the parameter $\alpha$ such that the residual of the function is decreased sufficiently between two consecutive iterations.
At each step, we start from a candidate value for $\alpha$, and evaluate the ratio $\delta(\alpha)$ between the norms of the residual $S(\alpha)$ at iteration $k+1$ and $S$ at iteration $k$, defined as
(24) $\delta(\alpha)=\frac{\lVert S(\alpha)\rVert_{F}}{\lVert S\rVert_{F}}.$
If $\delta(\alpha)<1$, we declare our update successful and the new iterate is set to $(W^{k+1},H^{k+1},Z^{k+1})=(W(\alpha),H(\alpha),Z(\alpha))$. Furthermore, let $\bar{\delta}$ be a fixed threshold, if $\delta(\alpha)>\bar{\delta}\in(0,1)$, meaning we do not consider the decrease of the residual fast enough, we increase $\alpha$ by setting $\alpha\leftarrow\min(\alpha+\mu,\bar{\alpha})$, where $\mu=\max(\mu,0.25(\alpha-1))$. If $\alpha$ reaches the upper bound $\bar{\alpha}$, we restart the parameter by setting $\alpha=1$. This correction does not appear in [43], but we added it to guarantee the convergence of the our algorithm. On the contrary, if $\delta(\alpha)\geq 1$, the step is declared unsuccessful, we do not accept the update, and we compute a standard BDC-NMD update instead by setting $\alpha=1$. Consequently, either we find an $\alpha\in(1,\bar{\alpha})$ that allows a decrease of the residual or we recover one iteration of the BCD-NMD, which is guaranteed to decrease the residual. Therefore, the new iterate of the algorithm produces a lower value in the residual with respect to the previous iteration.
The overall eBCD-NMD scheme is described in Algorithm 2, where we consider the updating scheme in (21). We choose the improved version of eBCD-NMD because (a) it exhibits faster performance in practice; (b) computing one QR factorization is in general more stable than solving two consecutive least squares problems; (c) one of the factors in the decomposition is orthogonal and this helps demonstrate crucial properties in the convergence analysis, such as boundness of the sequence generated by the eBCD-NMD algorithm. Let us also mention that, in [43], convergence is proved for a scheme similar to the one in (18), instead of the scheme in (21) as done in this paper.
Algorithm 2 eBCD-NMD
0:  $X$, $Z^{0}$,$W^{0}$, $H^{0}$, $\bar{\alpha}$ (default $=4$), $\mu$ (default $=0.3$), $\bar{\delta}$ (default $=0.8$), maxit
1:  Set $\Omega=\{(i,j)\ |\ X_{ij}>0\}$ and $\alpha_{0}=1$.
2:  for $k=0,\dots,$ maxit  do
3:     $Z_{\alpha_{k}}=\alpha_{k}Z^{k}+(1-\alpha_{k})W^{k}H^{k}$
4:     Compute $Q(\alpha_{k})$ orthogonal basis of $\mathcal{R}(Z_{\alpha_{k}}(H^{k})^{T})$
5:     $W(\alpha_{k})=Q(\alpha_{k})$
6:     $H(\alpha_{k})=Q(\alpha_{k})^{T}Z_{\alpha_{k}}$
7:     $Z(\alpha_{k})=P_{\Omega}(X)+P_{\Omega^{C}}(\min(0,W(\alpha_{k})H(\alpha_{k})))$
8:     if $\delta(\alpha_{k})\geq 1$ then
9:        Set $\alpha_{k}$ to 1
10:        $(W^{k+1},H^{k+1},Z^{k+1})=(W^{k},H^{k},Z^{k})$
11:     else
12:        $(W^{k+1},H^{k+1},Z^{k+1})=(W(\alpha_{k}),H(\alpha_{k}),Z(\alpha_{k}))$
13:        if $\delta(\alpha_{k})\geq\bar{\delta}$ then
14:           $\mu=\max(\mu,0.25(\alpha_{k}-1))$ and $\alpha_{k+1}=\min(\alpha_{k}+\mu,\bar{\alpha})$
15:           if $\alpha_{k+1}=\bar{\alpha}$ then
16:              $\alpha_{k+1}=1,$
17:           end if
18:        else
19:           $\alpha_{k+1}=\alpha_{k}.$
20:        end if
21:     end if
22:  end for
4.3.3 Convergence analysis
We present now the convergence analysis of Algorithm 2. The proofs of the lemmas can be found in the Appendix. The main steps of the convergence analysis are summarized as follows.
  1. Step 1
    We first prove that there exists a range of values larger than 1 for $\alpha$ such that the scheme (21) decreases the residual, that is,
    (25) $\lVert S\rVert_{F}^{2}-\lVert S(\alpha)\rVert_{F}^{2}>0.$
    Specifically, we split the residual reduction in (25) into the contribution given by the update of each variable separately. In particular, Lemma 4.5 characterizes the residual reduction after the update of $W$ and $H$, showing that it is nonnegative. Lemma 4.6 and Lemma 4.7 describe the residual reduction after updating the entries of $Z$.
  2. Step 2
    We provide the KKT conditions for problem (5); see (32).
  3. Step 3
    We use the residual reduction that we characterized in Step 1 to prove subsequence convergence of the eBCD-NMD algorithm in Theorem 4.9, showing that a subsequence of the eBCD-NMD iterates satisfies the KKT conditions at the limit. This is done analyzing all possible situations that the restarting scheme arises.
Step 1
Let us denote $\sigma_{WH}(\alpha)$ the variation of the residual after the update of $W$ and $H$ and $\sigma_{Z}(\alpha)$ the variations after updating the entries of $Z$, that is,
$\lVert S\rVert_{F}^{2}-\lVert S(\alpha)\rVert_{F}^{2}=\underbrace{\lVert S\rVert_{F}^{2}-\frac{1}{\alpha^{2}}\lVert W(\alpha)H(\alpha)-Z_{\alpha}\lVert_{F}^{2}}_{=:\sigma_{WH}(\alpha)}+\underbrace{\frac{1}{\alpha^{2}}\lVert W(\alpha)H(\alpha)-Z_{\alpha}\lVert_{F}^{2}-\rVert S(\alpha)\rVert_{F}^{2}}_{=:\sigma_{Z}(\alpha)}$
Let $U$ be an orthogonal basis of $\mathcal{R}(H^{T})$. If $H$ is full rank, we can define the orthogonal projections onto $\mathcal{R}(W(\alpha))$ and $\mathcal{R}(H^{T})$ as:
(26) $E(\alpha)=W(\alpha)W(\alpha)^{T}=Q(\alpha)Q(\alpha)^{T},$
$P=UU^{T}=H^{T}(HH^{T})^{-1}H.$
The following three lemmas characterize $\sigma_{WH}(\alpha)$ and $\sigma_{Z}(\alpha)$ (the proofs can be found in the appendix).
Lemma 4.5.
Given $(Z,W,H)$ where $H$ is full rank, let $(W(\alpha),H(\alpha))$ be computed by the scheme (21), then for any $\alpha\geq 1$,
(27) $\sigma_{WH}(\alpha)=\frac{1}{\alpha^{2}}\lVert W(\alpha)H(\alpha)-WH\rVert_{F}^{2}=\lVert SP\rVert_{F}^{2}+\lVert E(\alpha)S(I-P)\rVert_{F}^{2}\geq 0,$
where $E(\alpha)$ and $P$ are defined in (26).
Therefore, after the first two steps, the residual reduction is strictly positive unless $W(\alpha)H(\alpha)=WH$. We look at the residual reduction after updating the entries of $Z$ in $\Omega^{C}$ (Lemma 4.6) and in $\Omega$ (Lemma 4.7). In the following lemmas, we assume that there exists an interval $[1,\widehat{\alpha}]$, such that $Z_{\alpha}H^{T}$ and $ZH^{T}$ have the same rank. This hypothesis is needed to ensure the continuity of the residual reduction $\sigma_{Z}(\alpha)$ in a range of $\alpha>1$; more details are given in the Appendix.
Lemma 4.6.
Let $(W(\alpha),H(\alpha))$ be generated by the scheme in (21). Furthermore, assume there exists $\widehat{\alpha}>1$ such that $\operatorname{rank}(Z_{\alpha}H^{T})=\operatorname{rank}(ZH^{T})$ for $\alpha\in[1,\widehat{\alpha}]$, then there exists $\alpha_{*}>1$ such that for every $\alpha\in(1,\min(\alpha_{*},\widehat{\alpha})]$,
(28) $\frac{1}{\alpha^{2}}\lVert P_{\Omega^{C}}(W(\alpha)H(\alpha)-Z_{\alpha})\rVert_{F}^{2}-\lVert P_{\Omega^{C}}(S(\alpha))\rVert_{F}^{2}\geq 0.$
Moreover, if $P_{\Omega^{C}}(\min(0,W(1)H(1)))\neq P_{\Omega^{C}}((\min(0,WH))$, a strict inequality holds.
A similar result is proved in [43] for matrix completion. In their less restrictive setting, the residual reduction after updating the entries of $Z$ in $\Omega^{C}$ is independent of $\alpha$ and nonnegative. However, due to the additional constraint $P_{\Omega^{C}}(Z)\leq 0$ in the 3B-ReLU-NMD formulation, the residual reduction in (28) of the eBCD-NMD algorithm is only locally nonnegative in a range where $\alpha$ is larger than one.
Analogously to Lemma 4.6, we now state a similar result that characterizes the residual reduction when the entries of $Z$ in $\Omega$ are updated.
Lemma 4.7.
Let $(W(\alpha),H(\alpha))$ be generated by the scheme in (21). Furthermore, assume there exists $\widehat{\alpha}>1$ such that rank($Z_{\alpha}H^{T}$)=rank($ZH^{T}$), for $\alpha\in[1,\widehat{\alpha}]$. Then
(29) $\lim_{\alpha\to 1^{+}}\frac{1}{\alpha^{2}}\lVert P_{\Omega}(W(\alpha)H(\alpha)-Z_{\alpha})\rVert_{F}^{2}-\lVert P_{\Omega}(S(\alpha))\rVert_{F}^{2}=0.$
The following corollary gathers the findings from Lemma 4.5, Lemma 4.6, and Lemma 4.7.
Corollary 4.8.
Let $(W(\alpha),H(\alpha),Z(\alpha))$ be generated by the scheme in (21). Furthermore, assume there exists $\widehat{\alpha}>1$ such that rank($Z_{\alpha}H^{T}$)=rank($ZH^{T}$) for $\alpha\in[1,\widehat{\alpha}]$. If $P_{\Omega^{C}}(\min(0,WH))\neq P_{\Omega^{C}}(\min(0,W(1)H(1)))$, then there exists $\tilde{\alpha}>1$ such that
(30) $\lVert S\rVert_{F}^{2}-\lVert S(\alpha)\rVert_{F}^{2}=\sigma_{WH}(\alpha)+\sigma_{Z}(\alpha)>0\quad\forall\alpha\in[1,\min(\widehat{\alpha},\tilde{\alpha})].$
Corollary 4.8 guarantees that there exist some values of the extrapolation parameter $\alpha$ strictly larger than 1 resulting in a positive residual reduction for eBCD-NMD (21).
Step 2
Before stating the actual convergence theorem, we need to introduce the optimality conditions or KKT conditions for problem (5). Let us define the Lagrangian function as
(31) $\mathcal{L}(W,H,Z,\Lambda,\Sigma)=\frac{1}{2}\lVert Z-WH\rVert_{F}^{2}+\langle\Lambda,P_{\Omega}(Z-X)\rangle+\langle\Sigma,P_{\Omega^{C}}(Z)\rangle,$
where the dual variables are such that $\Lambda=P_{\Omega}(\Lambda)$ and $\Sigma=P_{\Omega^{C}}(\Sigma)$:
(32) Stationarity condition: $\nabla_{W}\mathcal{L}(Z,W,H)=-(Z-WH)H^{T}=0,$
$\nabla_{H}\mathcal{L}(Z,W,H)=-W^{T}(Z-WH)=0,$
$\nabla_{Z}\mathcal{L}(Z,W,H,\Lambda,\Sigma)=Z-WH+\Lambda+\Sigma=0,$
Primal feasibility: $P_{\Omega}(Z-X)=0,$
$P_{\Omega^{C}}(Z)\leq 0,$
Dual feasibility: $\Sigma\geq 0,$
Complementary slackness: $\langle\Sigma,P_{\Omega^{C}}(Z)\rangle=0,$
where the multiplier matrices $\Lambda$ and $\Omega$ measure the residual in $\Omega$ and $\Omega^{C}$ respectively, that is, $\Lambda=P_{\Omega}(WH-Z)$ and $\Sigma=P_{\Omega^{C}}(\max(0,WH))$. Our goal is to prove that, under some mild assumptions, a subsequence of the sequence generated by eBCD-NMD satisfies the KKT condition at the limit. Specifically, one can verify that the eBCD-NMD sequence satisfies the complementary slackness condition and the stationarity condition for $Z$ at any point. Therefore, we just need to show that the optimality conditions for $W$ and $H$ are satisfied by a subsequence of the eBCD-NMD scheme at the limit, that is,
(33) $\lim_{k\to\infty}\nabla_{W}\mathcal{L}(Z^{k},W^{k},H^{k})=0,\qquad\lim_{k\to\infty}\nabla_{H}\mathcal{L}(Z^{k},W^{k},H^{k})=0.$
Step 3
We are now ready to prove the convergence result for Algorithm 2.
Theorem 4.9.
Let $(W^{k+1},H^{k+1},Z^{k+1})$ be generated by eBCD-NMD (Algorithm 2), and assume that $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ is bounded for all $k$. Then there exists a bounded subsequence of $(W^{k+1},H^{k+1},Z^{k+1})$ whose limit point satisfies the KKT conditions (32).
Proof 4.10.
We first discuss the boundness of the two sequences $\{W^{k+1}\}$ and $\{H^{k+1}\}$ produced by Algorithm 2. Te boundness of $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ implies that $\{Z^{k}\}$ and $\{W^{k}H^{k}\}$ are bounded. This also means that $\{Z_{\alpha_{k}}\}$ is a bounded sequence by construction. Moreover, the sequence $\{W^{k+1}\}$ contains orthogonal matrices whose columns are orthogonal bases for $\mathcal{R}(Z_{\alpha_{k}}(H^{k})^{T})$, thus it is bounded. Finally, $H^{k+1}=(W^{k+1})^{T}Z_{\alpha_{k}}$ and since $\{W^{k+1}\}$ and $\{Z_{\alpha_{k}}\}$ are bounded, then also $\{H^{k+1}\}$ is bounded.
We observe that at every iteration $H^{k+1}$ has always the same rank as $W^{k+1}$. This means that $H^{k+1}$ is full rank for every $k$ and Lemma 4.5 can be applied. In fact, it follows from Lemma 4.3 that $\mathcal{R}(Q^{k+1})=\mathcal{R}(Z_{\alpha_{k}}(H^{k})^{T})\subseteq\mathcal{R}(Z_{\alpha_{k}})$; that implies
$\mathcal{R}(I)=\mathcal{R}((Q^{k+1})^{T}Q^{k+1})\subseteq\mathcal{R}((Q^{k+1})^{T}Z_{\alpha_{k}})=\mathcal{R}(H^{k+1}),$
where $I$ is the identity matrix of dimension equal to the number of columns of $Q^{k+1}$ or equivalently to the rank of $Z_{\alpha_{k}}(H^{k})^{T}$.
Next, we define the set of the iteration indices where the update of $Z^{k}$ results in a positive residual reduction. Let =$\mathcal{F}=\left\{k\ :\ \sigma_{Z}(\alpha_{k})\geq 0\right\}$ and $\mathcal{F}^{C}$ its complement. Therefore, denoting $E^{k}=E(\alpha_{k})$, since $H^{k+1}$ is full rank, we can use Lemma 4.5 and summing over all $k\in\mathcal{F}$, we get
(34) $\lVert S^{0}\rVert_{F}^{2}\geq\sum_{k\in\mathcal{F}}\sigma_{WH}(\alpha_{k})=\sum_{k\in\mathcal{F}}\lVert S^{k}P^{k}\rVert_{F}^{2}+\lVert E^{k}S^{k}(I-P^{k})\rVert_{F}^{2}.$
We analyze separately two different scenarios.
Case 1: assume $|\mathcal{F}^{C}|<\infty$. It follows from (34) that
(35) $\lim_{k\to\infty,k\in\mathcal{F}}\lVert S^{k}P^{k}\rVert_{F}^{2}=0,\quad\lim_{k\to\infty,k\in\mathcal{F}}\lVert E^{k}S^{k}(I-P^{k})\rVert_{F}^{2}=0,$
which implies
$\lim_{k\to\infty,k\in\mathcal{F}}\lVert E^{k}S^{k}\rVert_{F}^{2}=0.$
Let us recall that $P_{\Omega}(X)=P_{\Omega}(Z^{k})$, $P_{\Omega^{C}}(Z^{k})=P_{\Omega^{C}}(\min(0,W^{k}H^{k}))$ and using (26), we have
$S^{k}P^{k}=(Z^{k}-W^{k}H^{k})P^{k}=(Z^{k}-W^{k}H^{k})U^{k}(U^{k})^{T}.\$
Since $U^{k}$ is an orthonormal basis for $\mathcal{R}((H^{k})^{T})$ and $\{H^{k}\}$ bounded,
$\lim_{k\to\infty,k\in\mathcal{F}}(Z^{k}-W^{k}H^{k})(H^{k})^{T}=-\nabla_{W}\mathcal{L}(Z^{k},W^{k},H^{k})=0.$
Moreover,
(36) $E^{k}S^{k}=E^{k}S^{k+1}+E^{k}(S^{k}-S^{k+1})=W^{k+1}(W^{k+1})^{T}(Z^{k+1}-W^{k+1}H^{k+1})+E^{k}(S^{k}-S^{k+1}).$
Furthermore
(37) $\lVert S^{k}-S^{k+1}\rVert_{F}^{2}=$ $\lVert P_{\Omega}(W^{k+1}H^{k+1}-W^{k}H^{k})\rVert_{F}^{2}+\lVert P_{\Omega^{C}}(\max(0,W^{k+1}H^{k+1})-\max(0,W^{k}H^{k}))\rVert_{F}^{2}$
$\leq$ $\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}+\lVert\max(0,W^{k+1}H^{k+1})-\max(0,W^{k}H^{k})\rVert_{F}^{2}$
$\leq$ $2\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}.$
Using equation (27) and the fact that $\alpha_{k}\leq\bar{\alpha}$ by construction, we have
(38) $\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}\leq(\bar{\alpha})^{2}(\lVert S^{k}P^{k}\rVert_{F}^{2}+\lVert E^{k}S^{k}(I-P^{k})\rVert_{F}^{2}).$
Combining (38) and (35) we get
$\lim_{k\to\infty,k\in\mathcal{F}}\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}=0,$
thus, from (37),
(39) $\lim_{k\to\infty,k\in\mathcal{F}}\lVert S^{k}-S^{k+1}\rVert_{F}^{2}=0.$
Therefore, from (39) and (36), along with the boundness of $\{W^{k}\}$, we conclude that
$\lim_{k\to\infty,k\in\mathcal{F}}(W^{k+1})^{T}(Z^{k+1}-W^{k+1}H^{k+1})=\lim_{k\to\infty,k\in\mathcal{F}}-\nabla_{H}\mathcal{L}(Z^{k+1},W^{k+1},H^{k+1})=0.$
Case 2: $|\mathcal{F}^{C}|=\infty$. Let us analyze two subcases.
Case 2a: |$|\{k\in\mathcal{F}^{C}\ |\ \delta(\alpha_{k})>\bar{\delta}\}|<\infty$. For $k$ sufficiently large, we have
$\lVert S^{k+1}\rVert_{F}^{2}\leq\bar{\delta}\lVert S^{k}\rVert_{F}^{2}.$
Therefore, it holds $\lim_{k\to\infty,k\in\mathcal{F}^{C}}\lVert S^{k}\rVert_{F}^{2}=0,$ which means that the sequence converges to a global minimizer of the problem.
Case 2b: |$|\{k\in\mathcal{F}^{C}\ |\ \delta(\alpha_{k})>\bar{\delta}\}|=\infty$. The eBCD-NMD set the acceleration parameter $\alpha_{k}$ to 1 for an infinite number of iterations. Denote $\mathcal{J}_{1}$ the set of indices in which $\alpha_{k}=1$, a similar relation to (34) holds:
$\lVert S^{0}\rVert_{F}^{2}\geq\sum_{k\in\mathcal{J}_{1}}\sigma_{WH}(\alpha_{k})=\sum_{k\in\mathcal{J}_{1}}\lVert S^{k}P^{k}\rVert_{F}^{2}+\lVert E^{k}S^{k}(I-P^{k})\rVert_{F}^{2}.$
Therefore, one can replicate the proof for Case 1 and the subsequence whose indexes are in $\mathcal{J}_{1}$ satisfies the optimality conditions at the limit.
We emphasize that, similarly to BCD-NMD, we need a strong assumption, namely the boundness of $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ which guarantees that the eBCD-NMD sequence remains bounded. This hypothesis serves the same purpose as the existence of the limit point in the BCD-NMD algorithm, since it allows us to exclude pathological cases where the infimum of the problem is not attained.
5 Numerical experiments
In this section, we present numerical experiments on synthetic and real-world data sets, for different applications of ReLU decomposition: matrix completion with ReLU sampling, the recovery of Euclidean distance matrices, low-dimensional embedding on text data, and the compression sparse dictionaries. Our main purpose is to show that eBCD-NMD behaves favorably compared to the state of the art, namely Naive-NMD [34], EM-NMD [34], A-NMD [35], and 3B-NMD [35], while having convergence guarantees.
In all experiments, we use random initialization: entries of $W$ and $H$ are sampled from a Gaussian distribution, W = randn(m,r) and H = randn(r,n) in MATLAB. We then scale them so that the initial point matches the magnitude of the original matrix, that is, we multiply $W$ by $\sqrt{\lVert X\rVert_{F}}/\lVert W\rVert_{F}$ and $H$ by $\sqrt{\lVert X\rVert_{F}}/\lVert H\rVert_{F}$. If not otherwise specified, we stop each algorithm when we reach a small relative residual, namely $\Gamma_{k}=\lVert Z^{k}-W^{k}H^{k}\rVert_{F}/\lVert X\rVert_{F}\leq 10^{-9}$, or a fixed time limit is reached. The time limit will be selected depending on the size of the data. For the algorithms solving Latent-ReLU-NMD, the low-rank product $W^{k}H^{k}$ in $\Gamma_{k}$ is substituted with the matrix $\Theta^{k}$. The first stopping criteria can be achieved only if the target matrix admits an (almost) exact ReLU decomposition. The algorithms are implemented in MATLAB R2021b on a 64-bit Samsung/Galaxy with 11th Gen Intel(R) Core(TM) i5-1135G7 @ 2.40GHz and 8 GB of RAM, under Windows 11 version 23H2. The code is available online from https://github.com/giovanniseraghiti/ReLU-NMD .
The parameters of eBCD-NMD are the ones specified in Algorithm 2. For the parameters of the other methods, we use the setting from [35].
5.1 Matrix completion with ReLU sampling
In this subsection, we consider the same experiment used in [26] for symmetric matrix completion with ReLU sampling, adapted to the non-symmetric case. Specifically, we randomly generate the entries $W$ and $H$ from a Gaussian distribution, as for our initializations, and we construct $\Theta^{t}=WH$, and set $X=\max(0,\Theta^{t})$ in the noiseless case and $\max(0,\Theta^{t}+N)$ in the noisy case, where $N=\sigma*\tilde{N}*\lVert\Theta^{t}\rVert_{F}/\lVert\tilde{N}\rVert_{F}$, $\tilde{N}=$ randn(m,n), and $\sigma\in(0,1)$. In our experiment, we set $m=n=1000$, $r=20$ and $\sigma=10^{-2}$. Note that since $W$, $H$ and $N$ are Gaussian, $X$ is approximately $50\%$ sparse. We present the loglog plot of the average residual $\Gamma_{k}$ against the average CPU time over 20 different randomly generated matrices $X$ in Figure 2. The algorithms are stopped when the objective functions in (2) and (3) achieve a relative residual $\Gamma_{k}$ below $10^{-9}$ in the noiseless case, or below $\sigma=10^{-2}$ in the noisy case. We also report the average CPU time to reach these values in Table 1.
Figure 2: Matrix completion with ReLU sampling: evolution of the average relative residual $\Gamma_{k}$ w.r.t. CPU time for the noiseless case (left image) and the noisy case (right image).
BCD-NMD eBCD-NMD 3B-NMD Naive-NMD A-NMD EM-NMD
Noiseless 7.92 $\pm$ 0.16 0.90 $\pm$ 0.04 1.95 $\pm$ 0.04 22.0 $\pm$ 0.78 6.37 $\pm$ 0.33 26.7 $\pm$ 0.70
Noisy 1.14 $\pm$ 0.05 0.19 $\pm$ 0.01 0.66 $\pm$ 0.02 3.17 $\pm$ 0.12 1.54 $\pm$ 0.19 2.70 $\pm$ 0.15
Table 1: Average CPU time in seconds and standard deviation to reach a tolerance on the relative residual $\Gamma_{k}$ of $10^{-9}$ in the noiseless case, and of $10^{-2}$ in the noisy case. Best values are highlighted in bold.
Figure 2 shows that all the algorithms reach the tolerance on the relative residual. Moreover, eBCD-NMD outperforms all the other methods, being considerably faster both in the noiseless and noisy cases, as Table 1 confirms. In fact, compared to the second best, eBCD-NMD is on average more than twice faster in the noiseless case, and more than three time faster in the noisy case (each time, the second best is 3B-NMD).
5.2 EDMC
[…]
Figure 3: Euclidean distance matrix completion. Distribution of points on the left and relative error with the Euclidean distance matrix of the points in the right. The $x$-axis represents the percentage of known entries.
We observe from Figure 3 that the distribution of the points affects the recovery performance of the algorithms:
  • When the points are randomly distributed, $30\%$ of the entries needs to be observed for the eBCD-NMD to achieve an error below $10^{-7}$.
  • When the points are clustered, $50\%$ of the entries needs to be observed for the eBCD-NMD to achieve an error below $10^{-7}$.
Including the rank-1 modification in the model allows for a significantly more accurate reconstruction than solving a rank-($r+1$) 3B-ReLU-NMD. In general, A-NMD, eBCD-NMD and 3B-NMD provide the best reconstructions. However, 3B-NMD fails at least once in both examples when the percentage of observed entries is above $80\%$. This behavior might be due to the aggressive extrapolation procedure that 3B-NMD uses. We emphasize that neither the 3B-ReLU-NMD nor the Latent-ReLU-NMD formulations impose any constraints to ensure that the solution is symmetric. However, if the number of observed entries is sufficiently large, the algorithms still recover the original symmetric solution.
5.3 Low-dimensional embedding
In this subsection, we test eBCD-NMD within the threshold similarity matching (TSM) framework of Saul [33] for dimensionality reduction. Specifically, ReLU-NMD is employed to find a lower-dimensional representation of the similarity matrix of the input points ; see Section 2 for the details. Our focus is on text data, where each row of the data matrix corresponds to a document, each column to a word, and the entries $(i,j)$ of the matrix is the number of occurrences of the $j$th word in the $i$th document. We consider two different well-known and already pre-pocessed text-word data sets from [44], that are k1b which has 2340 documents and 21839 words and hitech that has 2301 documents and 10080 words. The threshold parameter $\tau$ in the TSM framework is selected according to the statistical analysis proposed in [33], that yields $\tau=0.17$ for k1b and $\tau=0.08$ for hitech.
Our objective is to evaluate how eBCD-NMD performs in comparison to the state-of-the-art algorithms for computing ReLU decomposition in this context. The quality of the embedding is evaluated using the mean angular deviation (MAD) metric as suggested in [33]. The MAD measures the difference between the angles that the algorithm tries to preserve (namely the ones below $\cos^{-1}(\tau)$). Define $\mathcal{P}(\tau)=\{(i,j)\ |\ \langle z_{i},z_{j}\rangle-\tau\lVert z_{j}\rVert\lVert z_{i}\rVert>0\}$, the MAD is given by
$\Delta=\frac{1}{|\mathcal{P}(\tau)|}\sum_{(i,j)\in\mathcal{P}(\tau)}\left|\cos^{-1}\left(\frac{\langle z_{i},z_{j}\rangle}{\lVert z_{i}\rVert\lVert z_{j}\rVert}\right)-\cos^{-1}\left(\frac{\langle y_{i},y_{j}\rangle}{\lVert y_{i}\rVert\lVert y_{j}\rVert}\right)\right|,$
where the $z_{i}$’s are the higher-dimensional points, and the $y_{i}$’s the lower-dimensional embeddings. Figure 4 displays the MAD and average running time per iteration for increasing values of the approximation rank. We run all the algorithms with a fixed time limit of 60 seconds.
Figure 4: MAD analysis and average iteration time for increasing values of the rank in the $x$-axes. k1b dataset on top and hitech dataset at the bottom.
Even though we are not able to identify one algorithm that outperforms the others, Figure 4 shows that solving 3B-ReLU-NMD is to be preferred in this application. Indeed, BCD-NMD, eBCD-NMD and 3B-NMD achieve the lowest MAD in both experiments. In fact, these three algorithms have the lowest average iteration time, being more than one second faster per iteration than Naive-NMD, A-NMD and EM-NMD. For this reason, BCD-NMD, eBCD-NMD and 3B-NMD perform more iterations than Naive-NMD, A-NMD and EM-NMD in 60 seconds, resulting in more accurate approximations.
5.4 Compression of sparse data
The last application is the compression of sparse data. We consider sparse greyscale images: 10000 randomly selected images from MNIST and Fashion MNIST (fMNIST), both of size $784\times 10000$, and on two well-known sparse images, Phantom ($256\times 256$) and Satellite ($256\times 256$). We also use sparse, structured matrices from various applications from [10]: Trec11 ($235\times 1138$), mycielskian ($767\times 767$), lock ($1074\times 1074$), and beaconfd ($173\times 295$).
We choose the approximation rank $r$ to ensure a desired level of compression. Specifically, when computing the ReLU factors $W$ and $H$ of a sparse matrix $X\in\mathbb{R}^{m\times n}$ with $\operatorname{nnz}(X)$ non zero entries, the number of entries of $WH$ is $r(n+m)$. Hence, the rank $r$ is set to the closest integer such that $r\leq 0.5\cdot\operatorname{nnz}(X)/(n+m)$, which means we reduced the memory requirement by $50\%$. We compare eBCD-NMD with state-of-the-art methods for ReLU decomposition, and we compute the TSVD of the same rank as baseline.
Table 2 shows the relative error $\frac{\|X-\max(0,\bar{W}\bar{H})\|_{F}}{\|X\|_{F}}$, where $\bar{W}$ and $\bar{H}$ are the factors of the rank-$r$ approximation produced by any algorithm. We run each algorithm for 120 seconds and display the average error over 10 different random initializations.
Data set BCD-NMD eBCD-NMD 3B-NMD Naive-NMD A-NMD EM-NMD TSVD
MNIST 12.4% 11.8% 11.6% 12.5% 11.7% 13.2% 25.8%
fMNIST 9.7% 9.3% 9.3% 9.7% 9.2% 9.9% 14.0%
Phantom 5.1% 4.1% 4.3% 5.4% 4.5% 5.6% 19.2%
Satellite 14.0% 13.6% 13.7% 14.4% 14.0% 14.8% 26.0%
Trec11 27.4% 26.5% 26.5% 27.2% 26.3% 31.1% 57.3%
mycielskian 0.37% 0.05% 0.06% 0.53% 0.06% 3.6% 58.0%
lock1074 0.3% 0.001% 0.001% 1.5% 0.006% 66.7% 67.2%
beaconfd 20.6% 20.4% 20.0% 20.9% 20.7% 26.8% 39.1%
Table 2: Sparse matrices compression. Comparison of the ReLU decomposition error in percentage on various data sets. The best values are highlighted in bold.
We observe that ReLU models outperform the linear alternative TSVD, as all the algorithms yield approximations that are significantly more accurate than TSVD. Additionally, eBCD-NMD and 3B-NMD produce the most accurate solutions in seven out of eight examples; moreover, eBCD-NMD outperforms BCD-NMD, achieving a smaller compression error for all the data sets considered. We also observe that the structure of the matrix has an impact on the effectiveness of the compression using ReLU decomposition.
Finally, we propose two visual examples of $50\%$ compression for MNIST and Phantom in Figure 5, where we compare TSVD with the approximation $\max(0,\Theta)$ provided by the eBCD-NMD algorithm. In both instances, one can visually appreciate how the eBCD-NMD approximation is consistently more accurate than TSVD, as further confirmed by the average error reported in Table 2.
Original digits
TSVD compression
eBCD compression
Original
TSVD compression
eBCD compression
Figure 5: Comparison between $50\%$ compressed images using TSVD and eBCD on MNIST and Phantom data sets.
6 Conclusion
The ReLU decomposition is a relatively recent matrix decomposition model (Saul, 2022 [33]) that finds applications in the compression of sparse data, entry-dependent matrix completion, and manifold learning. Moreover, we showed, for the first time, how it can be used for the completion of Euclidean distance matrices when only the smallest (or largest) entries are observed. Our first main contribution is to establish the non-equivalence of f the two main models used to solve the ReLU problem, namely ReLU-NMD and Latent-ReLU-NMD (Corollary 3.5). Furthermore, we derive an explicit connection between the optimal function values of the two, demonstrating that the solution of Latent-ReLU-NMD is in the worst case twice larger than that of ReLU-NMD (Theorem 3.6). Our second main contribution is to prove convergence of the BCD-NMD scheme for solving 3B-ReLU-NMD by demonstrating that it falls into the general framework of exact BCD, where one of the subproblems is strictly quasi-convex (Theorem 4.1). Our third and most important contribution is a novel algorithm, eBCD-NMD, which uses extrapolation to accelerate eBCD-NMD. eBCD-NMD extends the well-known LMaFit algorithm from [43] for matrix completion. We proved subsequence convergence of the eBCD-NMD scheme to a stationary point under mild assumptions (Theorem 4.9). Finally, we provided extensive numerical results for 4 applications on synthetic and real data sets, showing that eBCD-NMD performs favorably compared to the to the state of the art, while having convergence guarantees which other algorithms lack.
Future work includes investigating the well-posedness of the standard ReLU-NMD in (1), designing randomized algorithms to tackle large-scale problems, and exploring more advanced ReLU decompositions such as $X\approx W_{1}H_{1}+\max(0,W_{2}H_{2})-\max(0,W_{3}H_{3})$ that could handle mixed-signed and non-sparse data.
Acknowledgments
G.S. and N.G. acknowledge the support by the European Union (ERC consolidator, eLinoR, no 101085607). The work of M.P. and G.S. was partially supported by INdAM-GNCS under the project CUP E53C23001670001. The research of M.P.  was partially granted by the Italian Ministry of University and Research (MUR) through the PRIN 2022 “MOLE: Manifold constrained Optimization and LEarning”, code: 2022ZK5ME7 MUR D.D. financing decree n. 20428 of November 6th, 2024 (CUP B53C24006410006).
7 Appendix
7.1 Proof of Lemma 4.5
Proof 7.1.
Let $(W(\alpha),H(\alpha))$ be generated by (21), and let $S=\mathcal{P}_{\Omega}(X-WH)-\mathcal{P}_{\Omega^{C}}(\max(0,WH))$ be the residual with respect to the $k$-th iterate $(W,H)$. Recall that $Z_{\alpha}=WH+\alpha S$, $P=H^{\top}(HH^{\top})^{-1}H$, and $E(\alpha)=W(\alpha)W(\alpha)^{T}$. We denote $Q(\alpha)R(\alpha)\Pi=Z_{\alpha}H^{T}$, the QR factorization of $Z_{\alpha}H^{T}$ with column pivoting, where $\Pi$ is a permutation matrix. Note that $\Pi$ is needed only if $Z_{\alpha}H^{T}$ is rank-deficient; otherwise, one can consider $\Pi$ as the identity matrix. We have
(40) $W(\alpha)H(\alpha)-WH=\left[W(\alpha)R(\alpha)\Pi(HH^{T})^{-1}-W\right]H+W(\alpha)\left[H(\alpha)-R(\alpha)\Pi(HH^{T})^{-1}H\right].$
The first term in (40) can be written as
(41) $\left[W(\alpha)R(\alpha)\Pi(HH^{T})^{-1}-W\right]H=$ $W(\alpha)R(\alpha)\Pi(HH^{T})^{-1}H-WH$
$=$ $Z_{\alpha}H^{T}(HH^{T})^{-1}H-W(HH^{T})(HH^{T})^{-1}H=Z_{\alpha}P-WHP$
$=$ $(Z_{\alpha}-WH)P=\alpha SP.$
Similarly, we want to write the second term of the summation in terms of the residual. We use (41) to get
$H(\alpha)=$ $W(\alpha)^{T}Z_{\alpha}=W(\alpha)^{T}(WH+\alpha S)=$
$=$ $W(\alpha)^{T}\left[W(\alpha)R(\alpha)\Pi(HH^{T})^{-1}H-(W(\alpha)R(\alpha)\Pi(HH^{T})^{-1}-W)H+\alpha S\right]$
$=$ $W(\alpha)^{T}\left[W(\alpha)R(\alpha)\Pi(HH^{T})^{-1}H+\alpha S(I-P)\right].$
We multiply both sides by $W(\alpha)$ to obtain
$W(\alpha)H(\alpha)=$ $W(\alpha)W(\alpha)^{T}\left[W(\alpha)R(\alpha)\Pi(HH^{T})^{-1}H+\alpha S(I-P)\right]$
$=$ $W(\alpha)R(\alpha)\Pi(HH^{T})^{-1}H+\alpha E(\alpha)S(I-P).$
Rearranging the terms, we have
(42) $W(\alpha)\left[H(\alpha)-R(\alpha)\Pi(HH^{T})^{-1}H\right]=\alpha E(\alpha)S(I-P).$
Finally, combining (40), (41), and (42) we obtain
(43) $W(\alpha)H(\alpha)-WH=$ $\left[W(\alpha)R(\alpha)\Pi(HH^{T})^{-1}-W\right]H+W(\alpha)\left[H(\alpha)-R(\alpha)\Pi(HH^{T})^{-1}H\right]$
$=$ $\alpha SP+\alpha Q(\alpha)S(I-P)$
$=$ $\alpha(I-Q(\alpha))SP+\alpha Q(\alpha)S.$
From the properties of the orthogonal projection, (43) implies
(44) $\lVert W(\alpha)H(\alpha)-WH\rVert_{F}^{2}=\alpha^{2}\lVert(I-Q(\alpha))SP\rVert_{F}^{2}+\alpha^{2}\lVert Q(\alpha)S\rVert_{F}^{2},$
which proves the second inequality in (27).
Let us now obtain the first inequality. From (43) and the properties of orthogonal projection,
(45) $\langle\alpha S,W(\alpha)H(\alpha)-WH\rangle=$ $\alpha^{2}\langle S,(I-E(\alpha))SP\rangle+\alpha^{2}\langle S,E(\alpha)S\rangle$
$=$ $\alpha^{2}\lVert(I-E(\alpha))SP\rVert_{F}^{2}+\alpha^{2}\lVert E(\alpha)S\rVert_{F}^{2}$
$=$ $\lVert W(\alpha)H(\alpha)-WH\rVert_{F}^{2},$
From the observation (23) on $Z_{\alpha}$, the definition of $\sigma_{WH}(\alpha)$, and from (45), we get
$\sigma_{WH}(\alpha)=$ $\lVert S\rVert_{F}^{2}-\frac{1}{\alpha^{2}}\lVert W(\alpha)H(\alpha)-Z_{\alpha}\rVert_{F}^{2}=\lVert S\rVert_{F}^{2}-\frac{1}{\alpha^{2}}\lVert W(\alpha)H(\alpha)-WH-\alpha S\rVert_{F}^{2}$
$=$ $\lVert S\rVert_{F}^{2}-\frac{1}{\alpha^{2}}\left[\lVert W(\alpha)H(\alpha)-WH\rVert_{F}^{2}+\alpha^{2}\lVert S\rVert_{F}^{2}-2\alpha\langle S,W(\alpha)H(\alpha)-WH\rangle\right]$
$=$ $\frac{1}{\alpha^{2}}\lVert W(\alpha)H(\alpha)-WH\rVert_{F}^{2},$
which proves the first equality in (27).
7.2 Proof of Lemma 4.6
Proof 7.2.
Let us first show the continuity of the function in (28) when $\alpha\to 1^{+}$. Recall that $Z_{1}=Z$ and that $W(\alpha)$ is an orthogonal basis of $\mathcal{R}(Z_{\alpha}H^{T})$. Therefore, the orthogonal projection operator into $\mathcal{R}(Z_{\alpha}H^{T})$ can be written as $W(\alpha)W(\alpha)^{T}=(Z_{\alpha}H^{T})(Z_{\alpha}H^{T})^{\dagger}$. If rank($Z_{\alpha}H^{T}$)=rank($ZH^{T}$) in some interval $[1,\widehat{\alpha}]$, then by the continuity of the Moore-Penrose inverse [37], we have
(46) $\lim_{\alpha\to 1^{+}}W(\alpha)W(\alpha)^{T}=\lim_{\alpha\to 1^{+}}(Z_{\alpha}H^{T})(Z_{\alpha}H^{T})^{\dagger}=(ZH^{T})(ZH^{T})^{\dagger}=W(1)W(1)^{T}.$
From (46) and the continuity of the product,
(47) $\lim_{\alpha\to 1^{+}}W(\alpha)H(\alpha)=$ $\lim_{\alpha\to 1^{+}}W(\alpha)W(\alpha)^{T}Z_{\alpha}=W(1)W(1)^{T}Z=W(1)H(1),$
which leads to the continuity of (28) when $\alpha\to 1^{+}$. Let us recall that
$Z_{\alpha}=WH+\alpha S,\ P_{\Omega^{C}}(S)=-P_{\Omega^{C}}(\max(0,WH)),\ P_{\Omega^{C}}(Z(\alpha))=P_{\Omega^{C}}(\min(0,(W(\alpha)H(\alpha))).$
We have
(48) $\frac{1}{\alpha^{2}}\lVert P_{\Omega^{C}}(W(\alpha)H(\alpha)-Z_{\alpha})\rVert_{F}^{2}-\lVert P_{\Omega^{C}}(S(\alpha))\rVert_{F}^{2}$
$=$ $\frac{1}{\alpha^{2}}\lVert P_{\Omega^{C}}(W(\alpha)H(\alpha)-Z_{\alpha})\rVert_{F}^{2}-\lVert P_{\Omega^{C}}(W(\alpha)H(\alpha)-Z(\alpha))\rVert_{F}^{2}$
$=$ $\frac{1}{\alpha^{2}}\lVert P_{\Omega^{C}}(W(\alpha)H(\alpha)-WH-\alpha S)\rVert_{F}^{2}-\lVert P_{\Omega^{C}}(W(\alpha)H(\alpha)-\min(0,(W(\alpha)H(\alpha)))\rVert_{F}^{2}$
$=$ $\frac{1}{\alpha^{2}}\lVert P_{\Omega^{C}}(W(\alpha)H(\alpha)-\min(0,WH)+(\alpha-1)\max(0,WH))\rVert_{F}^{2}-$
$-\lVert P_{\Omega^{C}}(W(\alpha)H(\alpha)-\min(0,(W(\alpha)H(\alpha)))\rVert_{F}^{2}.$
Since (28) is continuous for $\alpha\to 1^{+}$, we just need to show that it is larger than or equal to zero at $\alpha=1$. Using (47) and (48),
$\lVert P_{\Omega^{C}}(W(1)H(1)-\min(0,WH))\rVert_{F}^{2}-\lVert P_{\Omega^{C}}(\max(0,(W(1)H(1)))\rVert_{F}^{2}$
$=$ $\lVert P_{\Omega^{C}}(W(1)H(1))\rVert_{F}^{2}+\lVert P_{\Omega^{C}}(\min(0,WH))\rVert_{F}^{2}-2\langle P_{\Omega^{C}}(W(1)H(1)),P_{\Omega^{C}}(\min(0,WH))\rangle-$
$-\lVert P_{\Omega^{C}}(\max(0,(W(1)H(1)))\rVert_{F}^{2}$
$=$ $\lVert P_{\Omega^{C}}(\min(0,W(1)H(1)))\rVert_{F}^{2}+\lVert P_{\Omega^{C}}(\min(0,WH))\rVert_{F}^{2}-$
$-2\langle P_{\Omega^{C}}(\min(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle-2\langle P_{\Omega^{C}}(\max(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle$
$=$ $\lVert P_{\Omega^{C}}(\min(0,W(1)H(1))-(\min(0,WH))\rVert_{F}^{2}-2\langle P_{\Omega^{C}}(\max(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle\geq 0,$
which concludes the proof.
7.3 Proof of Lemma 4.7
Proof 7.3.
By an analogous argument to the one used in Lemma 4.7, we get the continuity of (29) as $\alpha\to 1^{+}$. Next, we show that (29) is zero at $\alpha=1$. By the definition of the residual in (22),
$\frac{1}{\alpha^{2}}\lVert P_{\Omega}(W(\alpha)H(\alpha)-Z_{\alpha})\rVert_{F}^{2}-\lVert P_{\Omega}(S(\alpha))\rVert_{F}^{2}$
$=$ $\lVert P_{\Omega}(W(\alpha)H(\alpha)-X+X-WH-\alpha S)\rVert_{F}^{2}-\lVert P_{\Omega}(S(\alpha))\rVert_{F}^{2}$
$=$ $\lVert-P_{\Omega}(S(\alpha))+(1-\alpha)P_{\Omega}(S))\rVert_{F}^{2}-\lVert P_{\Omega}(S(\alpha))\rVert_{F}^{2},$
which is equal to zero for $\alpha=1$.

### v2

@@
-An extrapolated and provably convergent algorithm for nonlinear matrix decomposition with the ReLU function
+package/hyperref/before
 Abstract
-Nonlinear matrix decomposition (NMD) with the ReLU function, denoted ReLU-NMD, is the following problem: given a sparse, nonnegative matrix $X$ and a factorization rank $r$, identify a rank-$r$ matrix $\Theta$ such that $X\approx\max(0,\Theta)$.
-This decomposition finds application in data compression, matrix completion with entries missing not at random, and manifold learning.
-The standard ReLU-NMD model minimizes the least squares error, that is, $\|X-\max(0,\Theta)\|_{F}^{2}$.
-The corresponding optimization problem is nondifferentiable and highly nonconvex.
-This motivated Saul to propose an alternative model, Latent-ReLU-NMD, where a latent variable $Z$ is introduced and satisfies $\max(0,Z)=X$ while minimizing $\|Z-\Theta\|_{F}^{2}$ (“A nonlinear matrix decomposition for mining the zeros of sparse data”, SIAM J.
+ReLU matrix decomposition (RMD) is the following problem: given a sparse, nonnegative matrix $X$ and a factorization rank $r$, identify a rank-$r$ matrix $\Theta$ such that $X\approx\max(0,\Theta)$.
+RMD is a particular instance of nonlinear matrix decomposition (NMD) that finds application in data compression, matrix completion with entries missing not at random, and manifold learning.
+The standard RMD model minimizes the least squares error, that is, $\|X-\max(0,\Theta)\|_{F}^{2}$.
+The corresponding optimization problem, least-squares RMD (LS-RMD), is nondifferentiable and highly nonconvex.
+This motivated Saul to propose an alternative model, dubbed Latent-RMD, where a latent variable $Z$ is introduced and satisfies $\max(0,Z)=X$ while minimizing $\|Z-\Theta\|_{F}^{2}$ (“A nonlinear matrix decomposition for mining the zeros of sparse data”, SIAM J.
 Math.
 Data Sci., 2022).
-Our first contribution is to show that the two formulations may yield different low-rank solutions $\Theta$; in particular, we show that Latent-ReLU-NMD can be ill-posed when ReLU-NMD is not, meaning that there are instances in which the infimum of Latent-ReLU-NMD is not attained while that of ReLU-NMD is.
-We also consider another alternative model, called 3B-ReLU-NMD, which parameterizes $\Theta=WH$, where $W$ has $r$ columns and $H$ has $r$ rows, allowing one to get rid of the rank constraint in Latent-ReLU-NMD.
-Our second contribution is to prove the convergence of a block coordinate descent (BCD) applied to 3B-ReLU-NMD and referred to as BCD-NMD.
-Our third contribution is a novel extrapolated variant of BCD-NMD, dubbed eBCD-NMD, which we prove is also convergent under mild assumptions.
-We illustrate the significant acceleration effect of eBCD-NMD compared to BCD-NMD, and also show that eBCD-NMD performs well against the state of the art on synthetic and real-world data sets.
+Our first contribution is to show that the two formulations may yield different low-rank solutions $\Theta$.
+We then consider a reparametrization of the Latent-RMD, called 3B-RMD, in which $\Theta$ is substituted by a low-rank product $WH$, where $W$ has $r$ columns and $H$ has $r$ rows.
+Our second contribution is to prove the convergence of a block coordinate descent (BCD) approach applied to 3B-RMD.
+Our third contribution is a novel extrapolated variant of BCD, dubbed eBCD, which we prove is also convergent under mild assumptions.
+We illustrate the significant acceleration effect of eBCD compared to BCD, and also show that eBCD performs well against the state of the art on synthetic and real-world data sets.
 keywords:
-Low-rank matrix approximations, nonlinear matrix decomposition, rectified liner unit, block coordinate descent, extrapolation
+Low-rank matrix approximations, nonlinear matrix decomposition, rectified linear unit, block coordinate descent, extrapolation.
 {AMS}
@@
 1 Introduction
-Nonlinear matrix decomposition (NMD) can be defined as follows: given a data matrix $X\in\mathbb{R}^{n\times m}$, a factorization rank $r$, NMD looks for a rank-$r$ matrix $\Theta\in\mathbb{R}^{m\times n}$ such that $X\approx f(\Theta)$, where $f(\Theta)$ applies the scalar function $f$ element-wise on $\Theta$, that is, $[f(\Theta)]_{i,j}=f(\Theta_{i,j})$ for all $i,j$.
+Nonlinear matrix decomposition (NMD) can be defined as follows: given a data matrix $X\in\mathbb{R}^{n\times m}$, a factorization rank $r$, NMD looks for a rank-$r$ matrix $\Theta\in\mathbb{R}^{m\times n}$ such that $X\approx f(\Theta)$, where $f(\Theta)$ applies the scalar function $f$ element-wise on $\Theta$, that is, $f(\Theta)]_{i,j}=f(\Theta_{i,j})$ for all $i,j$.
 The choice of $f$ depends on the type of data at hand.
-For example, one might choose $f(x)=\operatorname{sign}(x)$ for binary data in $\{-1,1\}$ [11], or if $X$ has a multiplicative structure, one can consider the Hadamard product decomposition with $f(x)=x^{2}$, which is also efficient for the compression of dense images  [9, 18] and for representing probabilistic circuits [27].
+For example, one might choose $f(x)=\operatorname{sign}(x)$ for binary data in $\{-1,1\}$ [delsarte1989low], or $f(x)=x^{2}$ for nonnegative data with a multiplicative structure, resulting in the Hadamard product decomposition which is efficient for the compression of dense images [ciaperoni2024hadamard, hyeon2021fedpara, wertz2025efficient] and for the representation of probabilistic circuits [loconte2024subtractive].
+For a more detailed description of the choices for the function $f$; see [awari2025alternating].
 In this work, we assume $X$ to be nonnegative and sparse, opting for $f$ to be the ReLU function, $f(\cdot)=\max(0,\cdot)$.
-This NMD decomposition, with $X\approx\max(0,\Theta)$, is referred to as the ReLU decomposition.
-Intuitively, the ReLU decomposition aims at finding a matrix $\Theta$ that substitutes the zero entries of the target matrix $X$ with negative values in order to decrease its rank.
-Minimizing the least squares error leads to an optimization problem referred to as ReLU-NMD:
+This NMD, with $X\approx\max(0,\Theta)$, is referred to as the ReLU matrix decomposition (RMD).
+This model has a close connection with a two-layer neural network, using ReLU as the activation function in the output layer; see [saul2022geometrical, saul2022nonlinear] for more details.
+Intuitively, RMD aims at finding a matrix $\Theta$ that substitutes the zero entries of the target matrix $X$ with negative values in order to decrease its rank.
+Note that some degree of sparsity on the matrix $X$ is an essential condition for RMD to be meaningful in practice.
+In fact, if $X$ has mostly positive entries, then $\Theta$ will also have mostly positive entries, and the ReLU nonlinear activation will be active only for few entries, leading to a model that is similar to the truncated singular value decomposition (TSVD).
+Minimizing the least squares error leads to an optimization problem referred to as LS-RMD:
 (1) $\min_{\Theta}\lVert X-\max(0,\Theta)\rVert_{F}^{2}\quad\mbox{such that }\quad\operatorname{rank}(\Theta)\leq r.$
-Note that some degree of sparsity on the matrix $X$ is an essential condition for ReLU-NMD to be meaningful in practice.
-In fact, if $X$ has mostly positive entries, then $\Theta$ will also have mostly positive entries and the ReLU nonlinear activation will be active only for few entries, leading to a model that is similar to standard low-rank matrix approximations such as the truncated singular value decomposition (TSVD).
-However, it is interesting to note that ReLU-NMD will always lead to a lower error than the TSVD.
-In fact, let $X_{r}$ be an optimal rank-$r$ approximation of $X\geq 0$ (e.g., obtained with the TSVD), then $\|X-X_{r}\|_{F}^{2}\geq\|X-\max(X_{r},0)\|_{F}^{2}$ since negative entries in $X_{r}$, if any, are replaced by zeros which better approximate the corresponding nonnegative entries in $X$.
-ReLU-NMD was introduced by Saul in [34, 33].
+It is interesting to note that LS-RMD will always lead to a lower error than the TSVD. In fact, let $X_{r}$ be an optimal rank-$r$ approximation of $X\geq 0$ (e.g., obtained with the TSVD), then
+$\|X-X_{r}\|_{F}^{2}\geq\|X-\max(X_{r},0)\|_{F}^{2}$ since negative entries in $X_{r}$, if any, are replaced by zeros which better approximate the corresponding nonnegative entries in $X$.
+LS-RMD was introduced by Saul in [saul2022geometrical, saul2022nonlinear].
 It is non-differentiable and non-convex, and hence difficult to tackle.
-Indeed, to the best of our knowledge, there exists only one algorithm that addresses (1) directly, by parametrizing $\Theta=WH$ and using a coordinate descent approach, optimizing one entry of $W$ and $H$ at a time; see [4].
-All other existing approaches rely on Saul’s alternative formulation, dubbed Latent-ReLU-NMD, and defined as
+Indeed, to the best of our knowledge, there exist only a few algorithms that address (1) directly [awari2025alternating, awaricoordinate].
+All other existing approaches rely on Saul’s alternative formulation, dubbed Latent-RMD, and defined as
 (2) $\min_{Z,\Theta}\lVert Z-\Theta\rVert_{F}^{2}\quad\mbox{such that}\quad\begin{cases}\mbox{rank}(\Theta)\leq r,\\ \max(0,Z)=X.\end{cases}$
-The objective function in (2) is differentiable, while the subproblems in variables $Z$ and $\Theta$ can be solved up to global optimality; see Section 4.
+The objective function in (2) is differentiable, while the subproblems in the variables $Z$ and $\Theta$ can be solved up to global optimality; see Section 4.
 However, for a general matrix $X$, the link between the solutions of (1) and (2) has not been explored in the literature.
-Our first contribution, in Section 3, is to understand better the connection between the two.
-A third formulation was proposed in [35] using the parametrization $\Theta=WH$, where $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$, which is common practice for low-rank matrix approximation, as the rank of $\Theta$ does not exceed $r$ [43, 38].
-This formulation, referred to as 3B-ReLU-NMD model, is
-(3) $\min_{Z,W,H}\lVert Z-WH\rVert_{F}^{2}\quad\mbox{such that}\quad\max(0,Z)=X,$
-where 3B stands for 3 blocks of variables.
-The objective function in (3) is again jointly nonconvex, but it is convex with respect to each matrix variable, $W$, $H$ and $Z$, and can be solved computationally faster, as the rank constraint is not explicitly involved.
-Problem (3) can be tackled using the block coordinate descent (BCD) method, optimizing each matrix variable sequentially [35, 41].
+Our first contribution, in Section 3, is to better understand the connection between the two.
+If we reparametrize $\Theta=WH$ in (2), where $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$, as suggested in [seraghiti2023accelerated], we obtain the following equivalent problem:
+(3) $\min_{Z,W,H}\lVert Z-WH\rVert_{F}^{2}\quad\text{ such that }\quad\max(0,Z)=X.$
+We denote this reformulation 3B-RMD, where 3B stands for 3 blocks of variables.
+The objective function in (3) is again jointly nonconvex, but it is convex with respect to each matrix variable, $W$, $H$ and $Z$, making it more friendly from an algorithmic point of view, as the rank constraint is not explicitly involved.
+Problems (2) and (3) can be tackled using the block coordinate descent (BCD) method, optimizing each matrix variable alternatively [seraghiti2023accelerated, wang2024momentum].
 Outline and contribution
-Section 2 motivates the study of ReLU-NMD by presenting several applications.
-In Section 3, we clarify the relationship between ReLU-NMD and Latent ReLU-NMD.
-In particular, we provide a matrix $X$ for which the infimum of the rank-1 Latent-ReLU-NMD is not attained, showing the ill-poseness of Latent-ReLU-NMD (Lemma 3.1).
-On the other hand, for the same matrix $X$, ReLU-NMD has a non-empty set of optimal rank-1 solutions (Lemma 3.3).
+Section 2 motivates the study of RMD by presenting several applications.
+In Section 3, we clarify the relationship between LS-RMD and Latent-RMD.
+In particular, we provide a matrix $X$ for which the infimum of the rank-1 Latent-RMD is not attained, showing the ill-poseness of this model ( Example 3.1).
+On the other hand, for the same matrix $X$, LS-RMD has a non-empty set of optimal rank-1 solutions ( Example 3.3).
 This shows that, in general, the two models do not share the same set of optimal low-rank solutions, that is, the sets
 […]
-do not necessarily coincide. However, we show that any feasible point $(\widehat{Z},\widehat{\Theta})$ of the Latent-ReLU-NMD model satisfies the relation
-$\lVert X-\max(0,\widehat{\Theta})\rVert_{F}\leq 2\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}.$
-This guarantees that any approximated solution of Latent-ReLU-NMD provides an approximated solution for ReLU-NMD whose residual is at most twice that of Latent-ReLU-NMD.
-In Section 4, we first briefly review the state-of-the-art algorithms for solving Latent-ReLU-NMD and 3B-ReLU-NMD (Section 4.1).
-We then prove convergence of the BCD scheme for solving 3B-ReLU-NMD (BCD-NMD) (Section 4.2), demonstrating that it falls within the framework of BCD with strictly quasi-convex subproblems [16].
-In addition, we introduce a novel extrapolated version of BCD-NMD (eBCD-NMD) (Section 4.3), adapting the well-known LMaFit algorithm for matrix completion [43].
-We prove subsequence convergence of eBCD-NMD (Theorem 4.9).
-To the best of our knowledge, BCD-NMD and eBCD-NMD are the first algorithms for solving 3B-ReLU-NMD with convergence guarantees.
-In Section 5, we provide extensive numerical experiments comparing our new approach, eBCD-NMD, with the state of the art.
+do not necessarily coincide. However, we show that any feasible point $(\bar{Z},\bar{\Theta})$ of the Latent-RMD model satisfies the relation
+[…]
+This guarantees that any approximated solution of Latent-RMD provides an approximated solution for LS-RMD whose residual is at most four times that of Latent-RMD.
+In Section 4, we first briefly review the state-of-the-art algorithms for solving Latent-RMD and 3B-RMD (Section 4.1).
+We then prove the convergence of the BCD scheme for solving 3B-RMD (Section 4.2), demonstrating that it falls within the framework of BCD with strictly quasi-convex subproblems [grippo2000convergence].
+In addition, we introduce a novel extrapolated version of BCD ( eBCD) (Section 4.3), adapting the well-known LMaFit algorithm for matrix completion [wen2012solving].
+We prove the subsequence convergence of eBCD (Theorem 4.9).
+To the best of our knowledge, BCD and eBCD are the first algorithms for solving 3B-RMD with convergence guarantees.
+In Section 5, we provide extensive numerical experiments comparing our new approach, eBCD, with the state of the art.
 We test the algorithms on various applications: matrix completion with ReLU sampling, Euclidean distance matrix completion (EDMC), compression of sparse data, and low-dimensional embedding.
-In particular, our experiments show that eBCD-NMD consistently accelerates BCD-NMD and performs favorably with the state of the art, while having stronger theoretical guarantees.
-2 Applications of ReLU-NMD
-In this section, we first show the link between the ReLU decomposition and neural networks.
-Then we present some applications of the ReLU decomposition, illustrating how nonlinear models can be used in a variety of different contexts.
-2.1 Neural network interpretation
-The increasing interest in nonlinear models can be attributed to the rising popularity of neural networks, which use nonlinear functions as activations in their hidden layers.
-Furthermore, low-rank nonlinear models, involving matrix decompositions, have been studied in the context of neural networks [11, 23, 28, 36].
-Let us discuss the link between ReLU-NMD and ReLU networks.
-Figure 1 illustrates a one hidden-layer neural network.
-$g$$f$$W$$Y$$a_{i}$$b_{i}$
-Figure 1: One hidden-layer neural network with 5 input and output nodes and 3 hidden nodes.
-Let $a_{i}\in\mathbb{R}^{p}$ be $p$-dimensional input and $b_{i}\in\mathbb{R}^{m}$ the corresponding output vectors, for $i=1,\dots,n$.
-Define the input and output matrices $A=[a_{1},\dots,a_{n}]\in\mathbb{R}^{p\times n}$ and $B=[b_{1},\dots,b_{n}]\in\mathbb{R}^{m\times n}$, respectively.
-We also denote $Y\in\mathbb{R}^{r\times p}$ the matrix of the weights between the input layer and hidden layer, and $W\in\mathbb{R}^{m\times r}$ the matrix of the weights between the hidden layer and the output layer.
-The functions $g$ and $f$ are the activation functions for the first and second layer, respectively.
-We are interested in finding an approximation of a given data $X$ using this neural network with one hidden layer.
-If we choose the error measure to be the least squares error, we need to minimize
-(4) $\lVert X-B\rVert_{F}^{2}\;=\;\lVert X-f(Wg(YA))\rVert_{F}^{2}.$
-This shows that the hidden representation of each data point is a nonlinear combination of the inputs of the network, that is, $h_{i}=g(Ya_{i})\in\mathbb{R}^{r}$, or $H=g(YA)\in\mathbb{R}^{r\times n}$ in matrix form.
-NMD is a relaxation of (4) in which the hidden representation of the input matrix, $H$, can be any real matrix of appropriate dimensions.
-Furthermore, the number of hidden neurons is equal to the rank of the decomposition, and it needs to be chosen a priori as when designing the structure of the network.
-If the number of hidden neurons is smaller than the dimension of the inputs, the hidden layer is a low-dimensional embedding of the data.
+In particular, our experiments show that eBCD consistently accelerates BCD and performs favorably compared to the state of the art, while having stronger theoretical guarantees.
+2 Applications of RMD
+In this section, we present some applications of RMD, illustrating how nonlinear models can be used in a variety of contexts.
 2.1 Compression of sparse data
-An intuitive application of ReLU decomposition models is the compression of sparse data.
+The first application is the compression of sparse data.
 Let $X\in\mathbb{R}^{m\times n}$ be a sparse and nonnegative matrix with $\operatorname{nnz}(X)$ non-zero entries, and let $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$ be such that $X\approx\max(0,WH)$.
 As long as $r(m+n)<\operatorname{nnz}(X)$, storing the factors $W$ and $H$ requires less memory than storing $X$.
-ReLU models can outperform the standard linear compression techniques such as the TSVD on sparse data, such as sparse images or dictionaries [34, 33, 35, 41, 4].
-An interesting example is the identity matrix of any dimension that can be exactly reconstructed using a rank-3 ReLU decomposition [11, 34].
+RMD typically outperforms standard linear compression techniques, including the TSVD on sparse data, such as sparse images or dictionaries [awaricoordinate, saul2022geometrical, saul2022nonlinear, seraghiti2023accelerated, wang2024momentum].
+An interesting example is the identity matrix of any dimension, which can be exactly reconstructed using a rank-3 RMD [delsarte1989low, saul2022nonlinear].
 2.2 Matrix completion with ReLU sampling
-The class of matrix completion problems with entries missing not at random (MNR) is characterized by the fact that the probability of an entry being missing depends on the matrix itself.
-For example, when completing a survey, people are likely to share nonsensitive information such as the name or the surname, while they might be more hesitant to disclose their phone number.
-Therefore, we might expect the missing entries to be concentrated in some area of the matrix more than others.
-Even though MNR matrix completion is still a relatively unexplored area, dropping the assumption that missing entries are independent of the matrix is a challenging problem that attracts more and more people as the increasing number of related works testifies [6, 29, 13, 7, 26].
+The class of matrix completion problems with entries missing not at random (MNR) is characterized by the probability of an entry being missing that depends on the matrix itself.
+For example, when completing a survey, people are likely to share nonsensitive information such as their name or surname, while they might be more hesitant to disclose their phone number.
+Therefore, we might expect the missing entries to be concentrated in some areas of the matrix more than others.
+Even though MNR matrix completion is still a relatively unexplored area, dropping the assumption that missing entries are independent of the matrix is a challenging problem that attracts more and more researchers as the increasing number of related works testifies [bhattacharya2022matrix, bi2021absence, ganti2015matrix, liu2024symmetric, naik2022truncated].
 Among them, Liu et al.
-[26] investigate the matrix completion problem when only positive values are observed, matrix completion with ReLU sampling.
+[liu2024symmetric] investigate the matrix completion problem when only positive values are observed, namely matrix completion with ReLU sampling.
 This instance of matrix completion falls within the MNR problems since the missing entries depend on the sign of the target matrix.
 Assume that $X$ is a nonnegative and sparse matrix and that there exists a rank-$r$ matrix $\Theta$ such that $X=\max(0,\Theta)$.
-In this case, finding a ReLU decomposition is equivalent to recovering the negative values, observing only the positive ones; thus, it is equivalent to matrix completion with ReLU sampling.
-This implies that we can use the algorithms for computing a ReLU decomposition to address this matrix completion problem, as illustrated in [26].
-Inspired by the connection with ReLU sampling matrix completion, we introduce an equivalent formulation of 3B-ReLU-NMD that we will use throughout the paper.
-Let $\Omega=\{(i,j)\ |\ X_{ij}>0\}$ be the set of positive entries of $X$ and let $P_{\Omega}(X)$ be equal to $X$ in $\Omega$ and zero elsewhere.
-We rewrite 3B-ReLU-NMD in (3) as
-(5) $\operatorname{argmin}_{Z,W,H}\frac{1}{2}\lVert Z-WH\rVert_{F}^{2}\quad\mbox{such that}\quad P_{\Omega}(Z)=P_{\Omega}(X),\quad P_{\Omega^{C}}(Z)\leq 0,$
+In this case, finding an RMD is equivalent to recovering the negative values observing the positive ones; thus, it is equivalent to matrix completion with ReLU sampling.
+This implies that we can use the algorithms for RMD to address this matrix completion problem, as illustrated in [liu2024symmetric].
+Inspired by the connection to ReLU sampling matrix completion, we introduce an equivalent formulation of 3B-RMD that we will use throughout the paper.
+Let $\Omega=\{(i,j)\ |\ X_{ij}>0\}$ be the set of positive entries of $X$, and let $P_{\Omega}(X)$ be equal to $X$ in $\Omega$ and zero elsewhere.
+We rewrite 3B-RMD in (3) as
+(4) $\min_{Z,W,H}\frac{1}{2}\lVert Z-WH\rVert_{F}^{2}\quad\mbox{such that}\quad P_{\Omega}(Z)=P_{\Omega}(X),\quad P_{\Omega^{C}}(Z)\leq 0,$
 where $\Omega^{C}$ is the complement of $\Omega$ and a reformulation of the constraint $\max(0,Z)=X$ is introduced.
-Using the matrix factorization $\Theta=WH$ is a standard approach in matrix completion [5, 19, 20, 21, 38], see [30, 25] for detailed surveys on the topic.
-To our knowledge, the problem (5), without the constraint $P_{\Omega^{C}}(Z)\leq 0$, was first addressed in [43] in the context of matrix completion.
-The novelties in our model are the additional information that the missing entries have negative signs, which translates into the constraints $P_{\Omega^{C}}(Z)\leq 0$, and the explicit dependence of the known index set $\Omega$ from the target matrix $X$.
-Since the formulations (3) and (5) are equivalent, we will refer to both as 3B-ReLU-NMD.
+To our knowledge, the problem (4), without the constraint $P_{\Omega^{C}}(Z)\leq 0$, was first addressed in [wen2012solving] in the context of matrix completion.
+The novelties in our model are the additional information that the missing entries have negative signs, which translates into the constraints $P_{\Omega^{C}}(Z)\leq 0$, and the explicit dependence of the known index set $\Omega$ on the target matrix $X$.
+Since the formulations (3) and (4) are equivalent, we will refer to both as 3B-RMD.
 2.3 Euclidean distance matrix completion (EDMC)
-We provide now one practical example of the MNR matrix completion problem that can be solved using the ReLU decomposition.
-It is the recovery of a low-rank matrix for which only the smallest or the largest entries are observed.
+We provide one practical example of MNR matrix completion problem that can be solved using RMD.
+Specifically, we address the recovery of a low-rank matrix for which only the smallest or the largest entries are observed.
 Let $\Theta$ be a low-rank matrix, and assume that we observe all the entries in $\Theta$ smaller than a known threshold $d$.
 We want to recover the missing entries, which are the ones larger than $d$.
-We can model this problem using a rank-1 modified ReLU decomposition: $X=\max(0,dee^{T}-\Theta)$, where $e$ denotes the vector of all ones of appropriate dimension.
+We can model this problem using a rank-1 modified RMD: $X=\max(0,dee^{T}-\Theta)$, where $e$ denotes the vector of all ones of appropriate dimension.
 The matrix $X$ is nonzero exactly in the entries where $\Theta$ is smaller than $d$.
 Of course, an analogous model can be derived when only the largest entries are observed.
-The 3B-ReLU-NMD formulation in (5) can be easily adapted to include the rank-1 modification, that is,
-(6) $\operatorname{argmin}_{Z,W,H}\frac{1}{2}\lVert dee^{T}-WH-Z\rVert_{F}^{2}\quad\mbox{such that}\quad P_{\Omega}(Z)=P_{\Omega}(X),\quad P_{\Omega^{C}}(Z)\leq 0.$
-The same idea applies also to the Latent-ReLU-NMD and all the algorithms can be adjusted to take into account the additional rank-one term.
+The 3B-RMD formulation in (4) can be easily adapted to include the rank-1 modification:
+(5) $\min_{Z,W,H}\frac{1}{2}\lVert dee^{T}-WH-Z\rVert_{F}^{2}\quad\mbox{such that}\quad P_{\Omega}(Z)=P_{\Omega}(X),\quad P_{\Omega^{C}}(Z)\leq 0.$
+The same idea applies also to the Latent-RMD and all the algorithms can be adjusted to take into account the additional rank-one term.
 This formulation is particularly interesting when the matrix $\Theta$ contains the squared distances between a collection of points $p_{1},\dots,p_{n}\in\mathbb{R}^{d}$.
-This problem is often referred to as Euclidean distance matrix completion (EDMC) [1, 12, 22, 39], and is particularly meaningful in sensor network localization problems where sensors can only communicate with nearby sensors.
+This problem is often referred to as Euclidean distance matrix completion (EDMC) [alfakih1999solving, fang2012Euclidean, krislock2012euclidean, tasissa2018exact], and is particularly meaningful in sensor network localization problems where sensors can only communicate with nearby sensors.
 Let $P=[p_{1},\dots,p_{n}]\in\mathbb{R}^{d\times n}$, then the corresponding Euclidean distance matrix is $\Theta_{i,j}=\|p_{i}-p_{j}\|_{2}^{2}$ for all $i,j$, so that $\Theta=e\text{diag}(P^{T}P)^{T}+\text{diag}(P^{T}P)e^{T}+2P^{T}P$ and it has at most rank $r=d+2$ and thus the rank of the factorization in (5) is known in advance.
-We will show in Section 5 that this slightly modifed ReLU-NMD model allows us to solve EDMC in this context.
+We will show in Section 5 that this slightly modified RMD allows us to solve EDMC in this context.
 2.4 Manifold learning
-Manifold learning aims at learning a norm-preserving and neighborhood-preserving mapping of high-dimensional inputs into a lower-dimensional space [24].
-Given a collection of points $\{z_{1},\dots,z_{m}\}$ in $\mathbb{R}^{N}$, Saul [33] looks for a faithful representation $\{y_{1},\dots,y_{m}\}$ in $\mathbb{R}^{r}$, with $r<N$.
-Let $\tau\in(0,1)$, Saul defines an embedding to be $\tau$-faithful if
+Manifold learning aims at learning a norm-preserving and neighborhood-preserving mapping of high-dimensional inputs into a lower-dimensional space [lee2007nonlinear].
+Given a collection of points $\{z_{1},\dots,z_{m}\}$ in $\mathbb{R}^{N}$, Saul [saul2022geometrical] looks for a faithful representation $\{y_{1},\dots,y_{m}\}$ in $\mathbb{R}^{r}$, with $r<N$.
+Let $\tau\in(0,1)$, an embedding is $\tau$-faithful if
 (6) $\max(0,\langle y_{i},y_{j}\rangle-\tau\lVert y_{i}\rVert\lVert y_{j}\rVert)=\max(0,\langle z_{i},z_{j}\rangle-\tau\lVert z_{i}\rVert\lVert z_{j}\rVert).$
-Equation (7) implies that the embedding must preserve norms (take $i=j$ to obtain $(1-\tau)\lVert y_{i}\rVert^{2}$ $=$ $(1-\tau)\lVert z_{i}\rVert^{2}$) as well as the small angles between two points $(z_{i},z_{j})$: when $\cos(z_{i},z_{j})>\tau$, we need $\cos(z_{i},z_{j})=\cos(y_{i},y_{j})$.
+Equation (6) implies that the embedding must preserve norms (take $i=j$ to obtain $(1-\tau)\lVert y_{i}\rVert^{2}$ == $(1-\tau)\lVert z_{i}\rVert^{2}$) as well as the small angles between two points $(z_{i},z_{j})$: when $\cos(z_{i},z_{j})>\tau$, we need $\cos(z_{i},z_{j})=\cos(y_{i},y_{j})$.
 Large angles have to remain large but do not need to be preserved, that is, $\cos(z_{i},z_{j})\leq\tau$ only requires $\cos(y_{i},y_{j})\leq\tau$.
@@
 $X_{ij}=\max(0,\langle z_{i},z_{j}\rangle-\tau\lVert z_{i}\rVert\lVert z_{j}\rVert)\quad\text{ for all }i,j.$
-Second, compute a ReLU-decomposition, $X\approx\max(0,\Theta)$, and the matrix $\max(0,\Theta)$ represents the similarities between points in the lower-dimensional space.
-Finally, the embedded points are recovered from the low-rank approximation $\Theta$; see [33] for more details.
-3 ReLU-NMD vs Latent-ReLU-NMD
-In this section, we explore the link between ReLU-NMD in (1) and Latent-ReLU-NMD in (2).
+Second, compute an RMD, $X\approx\max(0,\Theta)$, and the matrix $\max(0,\Theta)$ represents the similarities between points in the lower-dimensional space.
+Finally, the embedded points are recovered from the low-rank approximation $\Theta$; see [saul2022geometrical] for more details.
+3 LS-RMD vs Latent-RMD
+In this section, we explore the link between LS-RMD in (1) and Latent-RMD in (2).
 In particular, we aim to determine if the optimal low-rank approximations provided by the two models coincide.
-First, when $X$ contains negative values, Latent-ReLU-NMD is infeasible, since there exists no latent variable $Z$ such that $X=\max(0,Z)$.
-In contrast, ReLU-NMD still admits feasible solutions.
-Of course, the ReLU decomposition is more meaningful when $X$ is nonnegative which is the case we focus on in the following.
-In the exact case, when there exists a low-rank matrix $\Theta$ such that $X=\max(0,\Theta)$, ReLU-NMD and Latent-ReLU-NMD are equivalent: for any low-rank solution $\Theta^{*}$ of (1), there exists a latent matrix $Z^{*}$, such that $(Z^{*},\Theta^{*})$ is a solution of (2) and vice versa.
-Indeed, in the exact case there always exists a trivial latent solution $Z^{*}=\Theta^{*}$.
-We show in what follows that this is not true in general, when an exact decomposition does not exist.
-We demonstrate first that there exists a specific matrix $X$ such that the rank-1 Latent-ReLU-NMD is ill-posed since the infimum of (2) is not attained (see Lemma 3.1).
-Moreover, by showing that for the same matrix $X$, ReLU-NMD admits a non-empty set of optimal solutions (Lemma 3.3), we establish that the two formulations do not necessarily yield the same optimal approximations.
-Note that we could not establish the general well-poseness of ReLU-NMD, which is a direction of future research.
-Lemma 3.1.
+First, when $X$ contains negative values, Latent-RMD is infeasible since there exists no latent variable $Z$ such that $X=\max(0,Z)$.
+In contrast, LS-RMD still admits feasible solutions.
+Of course, the RMD is more meaningful when $X$ is nonnegative, which is the case we focus on in the following.
+In the exact case, if there exists a low-rank matrix $\Theta$ such that $X=\max(0,\Theta)$, LS-RMD and Latent-RMD are equivalent: for any low-rank solution $\Theta^{*}$ of (1), there exists a latent matrix $Z^{*}$, such that $(Z^{*},\Theta^{*})$ is a solution of (2) and vice versa.
+Indeed, in this case, there always exists a trivial latent solution $Z^{*}=\Theta^{*}$.
+We show in what follows that this is not always true if an exact decomposition does not exist.
+We demonstrate first that there exists a specific matrix $X$ such that the rank-1 Latent-RMD is ill-posed since the infimum of (2) is not attained (see Example 3.1).
+Moreover, by showing that for the same matrix $X$, LS-RMD admits a non-empty set of optimal solutions ( Example 3.3), we establish that the two formulations do not necessarily yield the same optimal approximations.
+Note that we could not establish the general well-poseness of LS-RMD, which is a direction of future research.
+Example 3.1.
 Let $0<\epsilon<1/\sqrt{2}$ be a fixed parameter, $r=1$, and
 (7) $X=\begin{pmatrix}1&0\\ \epsilon&1\end{pmatrix}$.
-Then the infimum of Latent-ReLU-NMD in (2) is not attained.
+Then the infimum of Latent-RMD in (2) is not attained.
 Proof 3.2.
-We need to show that there exists no solution of Latent-ReLU-NMD attaining the infimum of (2).
-Let us prove the result by contradiction, and assume there exists a solution to (2), denoted $(\widehat{Z},\widehat{\Theta})$.
-First note that, given $\widehat{Z}$, the optimal $\widehat{\Theta}$ must be a best rank-1 approximation of $\widehat{Z}$.
-Hence $\widehat{Z}$ is such that the distance with its best rank-$1$ approximation is minimized and such that it satisfies the constraint $\max(0,\widehat{Z})=X$.
-Let us denote the $i$-th singular values of $Z$ as $\sigma_{i}(Z)$, and $\widehat{Z}_{1}$ be a best rank-one approximation of $\widehat{Z}$.
-By the Eckart-Young theorem, $\|\widehat{Z}-\widehat{Z}_{1}\|_{F}^{2}=\sigma_{2}(Z)^{2}$ since $\widehat{Z}$ is a 2-by-2 matrix.
-Hence $\widehat{Z}$ must belong to
+[…]
 (8) $\operatorname{argmin}_{Z}\sigma_{2}(Z)^{2}\quad\mbox{such that }\quad\max(0,Z)=X.$
-Equivalently, we want to find $b\leq 0$ such that the matrix $Z=\begin{pmatrix}1&b\\ \epsilon&1\end{pmatrix}$ has the lowest possible second singular value.
-Since $Z$ is a $2\times 2$ matrix, we can compute the eigenvalues of $ZZ^{T}$, solve det$(ZZ^{T}-\sigma I)=0$ and get an explicit expression for the second singular value of $Z$: for a general matrix $\begin{pmatrix}a&b\\ c&d\end{pmatrix},$ the second singular value is given by
-$\sigma_{2}=\sqrt{\frac{S_{1}-S_{2}}{2}},$
-where
-$S_{1}=a^{2}+b^{2}+c^{2}+d^{2},\quad S_{2}=\sqrt{(a^{2}+b^{2}-c^{2}-d^{2})^{2}+4(ac+bd)^{2}}.$
-We rewrite (9) as
+[…]
 (9) $\operatorname{argmin}_{b<0}\;\ell(b):=\frac{2+b^{2}+\epsilon^{2}-\sqrt{(b^{2}-\epsilon^{2})^{2}+4(b+\epsilon)^{2}}}{2}.$
@@
 This means the infimum of $\ell(b)$ is not attained by any finite value of $b$, and hence $\widehat{Z}$ does not exist.
-We demonstrate in the following lemma that rank-1 ReLU-NMD admits an optimal solution for $X$ as in (8).
-Lemma 3.3.
-Let $X$ as in (8) and $r=1$, then for every $0<\epsilon<\frac{1}{\sqrt{2}}$, the infimum of ReLU-NMD is attained by all the matrices of the form
+We now demonstrate that rank-1 LS-RMD admits an optimal solution for $X$ as in (7).
+Example 3.3.
+Let $X$ as in (7) and $r=1$, then for every $0<\epsilon<\frac{1}{\sqrt{2}}$, the infimum of LS-RMD is attained by all the matrices of the form
 (10) $\Theta=\begin{pmatrix}1&-v\\ -\frac{1}{v}&1\end{pmatrix},\quad v\in\mathbb{R_{+}},$
@@
 Proof 3.4.
-Let us denote the error function of ReLU-NMD as $F(\Theta)=\lVert X-\max(0,\Theta)\rVert_{F}^{2}$.
+Let us denote the error function of LS-RMD as $F(\Theta)=\lVert X-\max(0,\Theta)\rVert_{F}^{2}$.
 Note that $2\times 2$ rank-1 matrices have limited patterns that the signs of their entries must follow.
-For instance, if the matrix does not contain zeros, the number of positive and negative entries must be even, otherwise the rank cannot be equal to one.
+For instance, if the matrix does not contain zeros, the number of positive and negative entries must be even; otherwise, the rank cannot be equal to one.
 We proceed by analyzing three different cases.
@@
 Otherwise, the positive entries on the diagonal can be set to 1 to minimize the error, which is exactly the matrix in (10), with error $\epsilon^{2}$.
-Case 3: The rank-1 matrix $\Theta$ contains only positive entries. If $\Theta\geq 0$, ReLU-NMD corresponds to the following problem without nonlinear activations:
-$\operatorname{argmin}_{\Theta}\lVert X-\Theta\rVert_{F}^{2}\ \text{ such that }\ \Theta\geq 0.$
-Therefore, we need to show that there exists no rank-1 matrix with nonnegative entries that achieves a residual value $F(\Theta)$ strictly smaller than $\epsilon^{2}$.
+Case 3: The rank-1 matrix $\Theta$ contains only positive entries.
+If $\Theta\geq 0$, the nonlinear activation has no effect on $\Theta$ and the ReLU-NMD corresponds to minimizing the function $\lVert X-\Theta\rVert_{F}^{2}$.
 Let $\widehat{\sigma}_{2}$ be the second singular value of $X$ as in (7).
-If rank($\Theta$)=1, then $\lVert X-\Theta\rVert_{F}^{2}\geq\widehat{\sigma}_{2}^{2}$ and using equation (10),
-$\widehat{\sigma}_{2}^{2}=\frac{2+\epsilon^{2}-\epsilon\sqrt{\epsilon^{2}+4}}{2}.$
-Simple calculations show that $\widehat{\sigma}_{2}^{2}$ is strictly larger than $\epsilon^{2}$ for any $0<\epsilon<\frac{1}{\sqrt{2}}$.
-As a consequence of Lemma 3.1 and Lemma 3.3, we have the following corollary.
+If rank($\Theta$)=1, then $\lVert X-\Theta\rVert_{F}^{2}\geq\widehat{\sigma}_{2}^{2}$ and using equation (9), we have $\widehat{\sigma}_{2}^{2}=\frac{2+\epsilon^{2}-\epsilon\sqrt{\epsilon^{2}+4}}{2}.$ Simple calculations show that $\widehat{\sigma}_{2}^{2}$ is strictly larger than $\epsilon^{2}$ for any $0<\epsilon<\frac{1}{\sqrt{2}}$.
+As a consequence of Example 3.1 and Example 3.3, we have the following corollary.
 Corollary 3.5.
-ReLU-NMD and Latent-ReLU-NMD may not share the same set of optimal low-rank solutions.
-We established the non equivalence between the low-rank solutions of ReLU-NMD and Latent-ReLU-NMD.
-However, in practice, we are interested in computing accurate solutions of ReLU-NMD even though they might not be optimal.
-Numerous numerical results show that the ReLU decomposition obtained by Latent-ReLU-NMD achieves a small residual error also for ReLU-NMD [34, 35, 41].
-The following theorem formalizes this observation, showing that any feasible point $(\widehat{\Theta},\widehat{Z})$ of the Latent model has a residual value for $\widehat{\Theta}$ in the original ReLU-NMD that is upper bounded by twice the residual of Latent-ReLU-NMD.
+LS-RMD and Latent-RMD may not share the same set of optimal low-rank solutions.
+We established the non equivalence between the low-rank solutions of LS-RMD and Latent-RMD.
+However, in practice, we are interested in computing accurate solutions of even though they might not be optimal.
+Numerous numerical results show that the decompositions obtained by solving the Latent-RMD model achieves a small residual error also for LS-RMD [saul2022nonlinear, seraghiti2023accelerated, wang2024momentum].
+The following theorem formalizes this observation, showing that any feasible point $(\bar{\Theta},\bar{Z})$ of the latent model has a residual value for in the original LS-RMD that is upper bounded by four times the residual of Latent-RMD.
 Theorem 3.6.
-Let $X\in\mathbb{R}^{n\times m}$ be a nonnegative matrix and denote as $(\widehat{\Theta},\widehat{Z})$ a feasible point of the Latent-ReLU-NMD model in (2), then
-(12) $\lVert X-\max(0,\widehat{\Theta})\rVert_{F}^{2}\leq 2\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}.$
+Let $X\in\mathbb{R}^{n\times m}$ be a nonnegative matrix and denote as $(\bar{\Theta},\bar{Z})$ a feasible point of the Latent-RMD model in (2); then
+(11) $\lVert X-\max(0,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}})\rVert_{F}^{2}\leq{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}}\lVert{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}\rVert_{F}^{2}$.
 Proof 3.7.
-We have
-(13) $\lVert X-\max(0,\widehat{\Theta})\rVert_{F}^{2}$ $=\lVert X-\widehat{Z}+\widehat{Z}-\max(0,\widehat{\Theta})-\min(0,\widehat{\Theta})+\min(0,\widehat{\Theta})\rVert_{F}^{2}$
-$\leq\lVert X-\widehat{Z}+\min(0,\widehat{\Theta})\rVert_{F}^{2}+\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}.$
-Let us define $\Omega=\{(i,j)\ |\ X(i,j)>0\}$ and $\widehat{\Omega}=\{(i,j)\ |\ \widehat{\Theta}(i,j)>0\}$, and $\Omega^{C}$ and $\widehat{\Omega}^{C}$ their respective complements.
-By assumption $\widehat{Z}$ is a feasible solution of (2) and hence $\max(0,\widehat{Z})=X$, therefore
-$\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}=$ $\sum_{(i,j)\in\Omega\cap\widehat{\Omega}}(X_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}(X_{ij}-\widehat{\Theta}_{ij})^{2}+$
-$+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}.$
-Similarly, observe that $Z_{i,j}\leq 0$ whenever $X_{i,j}=0$, and hence
-(14) $\lVert X-\widehat{Z}+\min(0,\widehat{\Theta})\rVert_{F}^{2}=$
-$=\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}\widehat{\Theta}_{ij}^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}\widehat{Z}_{ij}^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}$
-$\leq\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}(X_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}$
-$\leq\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}.$
-The first inequality holds because
-$\widehat{\Theta}_{ij}^{2}\leq(X_{ij}-\widehat{\Theta}_{ij})^{2}\mbox{ for }(i,j)\in\Omega\cap\widehat{\Omega}^{C},\text{ and }$
-$\widehat{Z}_{ij}^{2}\leq(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}\mbox{ for }(i,j)\in\Omega^{C}\cap\widehat{\Omega}.$
+Using the matrix inequality $\lVert A+B\rVert_{F}^{2}\leq 2\lVert A\rVert_{F}^{2}+2\lVert B\rVert_{F}^{2}$, we have
+(12) $\lVert X-\max(0,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}})\rVert_{F}^{2}$ $=\lVert X-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\min(0,\bar{\Theta})-\min(0,\bar{\Theta})-\max(0,\bar{\Theta})}}\rVert_{F}^{2}$
+[…]
+[…]
+[…]
+$+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}.$
+Similarly, ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{i,j}\leq 0$ whenever $X_{i,j}=0$; thus,
+$\lVert X-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}+\min(0,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}})\rVert_{F}^{2}$ $=\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij}^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}$
+$\leq\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}(X_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}+\hskip-5.69046pt\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}$
+[…]
+where the first inequality holds because
+${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij}^{2}\leq(X_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}\mbox{ for }(i,j)\in\Omega\cap\widehat{\Omega}^{C},\quad\text{ and }\quad{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}^{2}\leq({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}\mbox{ for }(i,j)\in\Omega^{C}\cap\widehat{\Omega}$.
 Combining (12) and (13), we get (11).
-A straightforward consequence of Theorem 3.6 is that a solution of Latent-ReLU-NMD with small error provides a ReLU-NMD solution with small error, even though the two models do not necessarily yield the same approximation.
-4 Algorithms for the ReLU decomposition
-This section is dedicated to the algorithms for computing ReLU decompositions and to the study of their convergence.
-In Section 4.1, we briefly review the state-of-the-art algorithms for solving Latent-ReLU-NMD and 3B-ReLU-NMD.
-In Section 4.2, we discuss the convergence of the BCD-NMD scheme.
-Finally, we present our new algorithm eBCD-NMD along with its convergence in Section 4.3.
+A straightforward consequence of Theorem 3.6 is that a solution of Latent-RMD with a small error provides a LS-RMD solution with a small error, even though the two models do not necessarily yield the same approximation.
+We specify that the same results holds if we reparametrize $\Theta$ as $WH$; thus, the same bound holds for 3B-RMD as well.
+4 Algorithms for RMD
+This section is dedicated to the algorithms for computing RMDs and to the study of their convergence.
+In Section 4.1, we briefly review the state-of-the-art algorithms for solving Latent-RMD and 3B-RMD.
+In Section 4.2, we discuss the convergence of the BCD scheme.
+Finally, we present our new algorithm eBCD along with its convergence in Section 4.3.
 4.1 Previous works
-Let us start with Latent-ReLU-NMD, since we use the algorithms designed to solve Latent-ReLU-NMD as baselines in the numerical comparison in Section 5.
-To our knowledge, all existing algorithms for Latent-ReLU-NMD are alternate optimization methods in which each block of variables is updated sequentially [16, 17, 32, 31, 40] and lack convergence guarantees.
-The first approach is the so-called Naive-ReLU-NMD, introduced by Saul in [34], which employs a basic alternating scheme over $Z$ and $\Theta$ with closed-form solutions for both.
-Similar approaches have been studied in matrix completion, see for example [5, 8, 10, 15].
+Let us start with the algorithm designed for Latent-RMD since we use them as baselines in the numerical comparison in Section 5.
+To our knowledge, all existing algorithms are alternating optimization methods in which each block of variables is updated sequentially [grippo2000convergence, grippo2023introduction, patriksson1998decomposition, powell1973search, tseng1991decomposition] and lack convergence guarantees.
+The first approach is the so-called Naive algorithm, introduced by Saul in [saul2022nonlinear], which employs a basic alternating scheme over $Z$ and $\Theta$ with closed-form solutions for both.
+Similar approaches have been studied in matrix completion, see for example [balzano2010online, boumal2015low, dai2012geometric, goyens2023nonlinear].
 At each iteration, firstly $\Theta$ is fixed and then the optimal $Z$ is computed solving the subproblem
@@
 The solution of (14) is given by the projection of $\Theta$ over the set $\{Z\ :\ \max(0,Z)=X\}$, that is,
-$Z_{ij}=\begin{cases}X_{ij}\quad&\mbox{if }X_{ij}>0\\ \min(0,\Theta_{ij})&\mbox{if }X_{ij}=0.\\ \end{cases}$
-Once $Z$ is updated, $\Theta$ is updated solving
-$\operatorname{argmin}_{\Theta}\lVert Z-\Theta\rVert_{F}^{2}\quad\mbox{such that}\quad\operatorname{rank}(\Theta)\leq r,$
-whose optimal solution(s) can be obtained with the TSVD of $Z$.
+[…]
+Once $Z$ is updated, $\Theta$ is updated solving $\operatorname{argmin}_{\Theta}\lVert Z-\Theta\rVert_{F}^{2}$ such that $\operatorname{rank}(\Theta)\leq r$, whose optimal solution(s) can be obtained with the TSVD of $Z$.
 Despite its simplicity, this scheme is highly reliable in practice, with the most costly operation at each iteration being the computation of a TSVD.
-An accelerated, heuristic version of the Naive-NMD scheme was proposed in [35].
-Specifically, the so-called Aggressive-ReLU-NMD (A-NMD) employs an extrapolation step for both $Z$ and $\Theta$ while the extrapolation parameter changes adaptively at each iteration, depending on the value of the residual.
-A-NMD is an example where, despite its lack of theoretical guarantees, extrapolation can be considerably faster [2, 43].
-Another algorithm for solving Latent-ReLU-NMD is the so-called expectation-minimization NMD (EM-NMD) [34].
+An accelerated, heuristic version of the Naive scheme was proposed in [seraghiti2023accelerated].
+Specifically, the so-called Aggressive Naive (A-Naive) algorithm, employs an extrapolation step for both $Z$ and $\Theta$ while the extrapolation parameter changes adaptively at each iteration, depending on the value of the residual.
+A-Naive is an example where, despite its lack of convexity and theoretical guarantees, extrapolation can be considerably faster the non accelerated counterpart [ang2019accelerating, wen2012solving].
+Another algorithm for solving Latent-RMD is the expectation-minimization (EM) algorithm [saul2022nonlinear].
 This approach is based on a Gaussian latent variable model parametrized by the low-rank matrix $\Theta$ and a parameter $\sigma$.
@@
 The method estimates $\Theta$ and $\sigma$ by maximizing the likelihood of the data with respect to $\Theta$ and $\sigma$.
-The optimization is carried out using the Expectation-Minimization (EM-NMD) scheme that consists of two main steps: the E-step which computes posterior mean and variance of the Gaussian latent variable model, and the M-step re-estimates the parameters of the model from the posterior mean and variance.
-The BCD scheme for solving 3B-ReLU-NMD in (5), referred to as BCD-NMD, consists of an alternating optimization procedure over three blocks of variables $(Z,W,H)$.
-BCD-NMD belongs to the class of exact BCD methods, where a global optimizer for each of the subproblem is computed sequentially.
-Let $A^{\dagger}$ denote the Moore-Penrose inverse of $A$ and let $(H^{k},Z^{k},W^{k})$ be the current iteration.
-The $(k+1)$-th step of BCD-NMD is as follows:
+The optimization is carried out using the expectation-minimization scheme that consists of two main steps: the E-step which computes posterior mean and variance of the Gaussian latent variable model, and the M-step re-estimates the parameters of the model from the posterior mean and variance.
+Finally, we present the BCD algorithm for solving the 3B-RMD model.
+In what follows, we use the formulation of 3B-RMD introduced in (4) rather than the equivalent one in (3).
+The BCD algorithm consists of an alternating optimization procedure over three blocks of variables $(Z,W,H)$, and computes a global optimizer for each of the subproblems sequentially.
+Let $A^{\dagger}$ denote the Moore-Penrose inverse of $A$, and let $(H^{k},Z^{k},W^{k})$ be the current iteration.
+The $(k+1)$-th step of BCD is as follows:
 (15) $Z^{k+1}=$ $\operatorname{argmin}_{Z}\{\lVert Z-W^{k}H^{k}\rVert_{F}^{2}\quad\text{s.t.}\quad\max(0,Z)=X\}=P_{\Omega}(X)+P_{\Omega^{C}}(\min(0,W^{k}H^{k})),$
 $W^{k+1}=$ $\operatorname{argmin}_{W}\lVert Z^{k+1}-WH^{k}\rVert_{F}^{2}=Z^{k+1}(H^{k})^{\dagger},$
-$H^{k+1}=$ $\operatorname{argmin}_{H}\lVert Z^{k+1}-W^{k+1}H\rVert_{F}^{2}=((W^{k+1})^{T})^{\dagger}Z^{k+1}.$
-The solution of the subproblems for $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$ are analogous.
-In particular, the first is separable by rows of $W$, leading to $m$ least squares problems in $r$ variables with the data vector of dimension $n$, while the subproblem for $H$ is separable by column and equivalent to $n$ least squares problems of $r$ variables and data vector of dimension $m$.
-Both these subproblems do not necessarily have a unique solution.
+$H^{k+1}=$ $\operatorname{argmin}_{H}\lVert Z^{k+1}-W^{k+1}H\rVert_{F}^{2}=((W^{k+1})^{T})^{\dagger}Z^{k+1},$
+where $P_{\Omega}$ and $P_{\Omega}^{C}$ are defined as in (4).
+The solutions of the subproblems for $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$ are analogous.
+In particular, the first is separable by rows of $W$, leading to $m$ least squares problems in $r$ variables with data vector of dimension $n$, while the subproblem for $H$ is separable by column, and equivalent to $n$ least squares problems of $r$ variables with data vector of dimension $m$.
+Both these subproblems do not necessarily have unique solutions.
 In fact, uniqueness is guaranteed only if $W^{k}$ and $H^{k}$ remain full rank throughout the iterations.
-Alternatively, uniqueness is ensured by adding a Tikhonov regularization term or a proximal term in each of the subproblems [16, 17, 3]; however, it is not necessary to prove the convergence of the BCD-NMD algorithm.
+Alternatively, uniqueness is ensured by adding a Tikhonov regularization term or a proximal term in each of the subproblems [attouch2010proximal, grippo2000convergence, grippo2023introduction].
+However, regularization might slow down the algorithm, and it is not necessary to prove the convergence of the BCD scheme in this context.
 On the contrary, the subproblem in the latent variable $Z$ admits a unique global minimum, which is the projection of the product $W^{k}H^{k}$ onto the feasible, convex set
 (16) $\{Z\in\mathbb{R}^{m\times n}\ :\ P_{\Omega}(Z)=P_{\Omega}(X),\ P_{\Omega^{C}}(Z)\leq 0\}.$
-The overall BCD-NMD algorithm is reported in Algorithm 1 and has a computational cost of $O(r|\Omega^{C}|+(m+n)r^{2})$.
-Algorithm 1 BCD-NMD: block coordinate descent for ReLU-NMD
+The overall BCD algorithm is reported in Algorithm 1 and has a computational cost of $O(r|\Omega^{C}|+(m+n)r^{2})$.
+Algorithm 1 BCD: block coordinate descent for 3B-RMD
 0:$X$, $W^{0}$, $H^{0}$, maxit.
@@
 6:end for
-Similarly to the Naive scheme for Latent-ReLU-NMD, adding extrapolation might be beneficial also for BCD-NMD.
-In particular, the work [35] proposes to perform an extrapolation step on the latent variable $Z$ and on the low-rank product $WH$, resulting in the so-called 3B-NMD algorithm.
+Similarly to the Naive scheme for Latent-RMD, adding extrapolation might be beneficial also for BCD.
+In particular, the work [seraghiti2023accelerated] proposes to perform an extrapolation step on the latent variable $Z$ and on the low-rank product $WH$, resulting in the so-called extrapolated 3B (e3B) algorithm.
 Moreover, Wang et al.
-[41, 42] added a Tikhonov regularization for both the variables $W$ and $H$ in the 3B-ReLU-NMD formulation.
-Their algorithm is dubbed as ReLU-NMD-T and it is based on the 3B-NMD scheme but performs two more extrapolation steps following the computation of $W$ and $H$.
-Even though numerical experience shows that both the accelerated algorithms outperform the original BCD-NMD, convergence still remains an open problem.
-For the numerical experiments in Section 5, we consider only 3B-NMD since we are interested in comparing algorithms that solve the non-regularized 3B-ReLU-NMD.
-4.2 Convergence of BCD-NMD
-In this section, we discuss the convergence to a stationary point of BCD-NMD (Algorithm 1) which, to the best of our knowledge, has not been studied before.
-Indeed, our goal is to prove that BCD-NMD is convergent under the assumption that a limit point of the sequence generated by the scheme exists.
-We proceed by showing that the method falls into the more general framework of exact block coordinate descent (BCD), where one of the subproblems is strictly quasi-convex.
-Specifically, the general BCD scheme is studied by Grippo et al.
-[16] when the objective function is continuously differentiable over the Cartesian product of $m$ closed, convex sets, one for each block of variables.
-We first observe that the 3B-ReLU-NMD objective function is continuously differentiable and defined over three blocks of variables $Z$, $W$, and $H$.
+[wang2024momentum, wang2025accelerated] added a Tikhonov regularization for both the variables $W$ and $H$.
+Their algorithm is based on the e3B scheme but performs two additional extrapolation steps following the computation of $W$ and $H$.
+Even though numerical experience shows that both accelerated algorithms outperform the original BCD, convergence still remains an open problem.
+4.2 Convergence of BCD
+In this section, we discuss the convergence to a stationary point of the BCD algorithm (Algorithm 1), which, to the best of our knowledge, has not been studied before in the context of RMD.
+Indeed, our goal is to prove that BCD is convergent under the assumption that a limit point of the sequence generated by the scheme exists.
+We proceed by showing that the method falls into the more general framework of exact BCD, where one of the subproblems is strictly quasi-convex.
+Specifically, the general BCD scheme is studied by Grippo and Sciandrone [grippo2000convergence] when the objective function is continuously differentiable over the Cartesian product of $m$ closed, convex sets, one for each block of variables.
+We first observe that the 3B-RMD objective function is continuously differentiable and defined over three blocks of variables $Z$, $W$, and $H$.
 The latent variable $Z$ is the only constrained variable, and it is easy to show that the constraint set in (16) is convex.
-Therefore, the 3B-ReLU-NMD is a particular instance of the more general problem considered in [16].
+Therefore, 3B-RMD is a particular instance of the more general problem considered in [grippo2000convergence].
 Theorem 4.1.
-Any limit point of the sequence generated by BCD-NMD (Algorithm 1) is a stationary point of problem (5).
+Any limit point of the sequence generated by BCD (Algorithm 1) is a stationary point of problem (4).
 Proof 4.2.
-The global convergence of the BCD scheme for a general function with $m$ blocks of variables is analyzed under generalized convexity assumptions in [16] (Proposition 5).
-The key hypotheses are: 1) the existence of a limit point of the sequence generated by the BCD algorithm, and 2) strict quasi-convexity of at least $m-2$ subproblems.
-Therefore, since 3B-ReLU-NMD has three blocks of variables, we just need one subproblem to be strictly quasi-convex.
-The first subproblem in the variable $Z$ is a convex, quadratic problem over a convex set, making it strictly quasi-convex, implying the convergence of BCD-NMD.
+The global convergence of the BCD scheme for a general function with $m$ blocks of variables is analyzed under generalized convexity assumptions in [grippo2000convergence] (Proposition 5).
+The key hypotheses are: 1) the existence of a limit point of the sequence generated by the BCD algorithm, and 2) strict quasi-convexity of at least m2m-2 subproblems.
+Therefore, since 3B-RMD has three blocks of variables, we just need one subproblem to be strictly quasi-convex.
+The first subproblem in the variable $Z$ is a convex, quadratic problem over a convex set, making it strictly quasi-convex, implying the convergence to a stationary point.
 The assumption of the existence of the limit point is an undesirable hypothesis.
 However, it is necessary because the level sets of the objective function are not bounded.
-Therefore, similarly to Latent-ReLU-NMD, 3B-ReLU-NMD might have instances in which the minimum is not attained (see Theorem 3.3).
-This means that, in pathological cases, the sequence generated by BCD-NMD may diverge simply because the infimum of the function is not attained by any finite value.
-Thus, the hypothesis regarding the existence of a limit point of the BCD-NMD sequence cannot be weakened.
-One possible way to overcome this limitation is by adding a regularization term to the objective function (5), which makes the level sets compact, as in [41].
+Therefore, similarly to Latent-RMD, 3B-RMD might have instances in which the minimum is not attained (see Theorem 3.3).
+This means that, in pathological cases, the sequence generated by BCD may diverge simply because the infimum of the function is not attained by any finite value.
+Thus, the hypothesis regarding the existence of a limit point of the BCD sequence cannot be weakened.
+One possible way to overcome this limitation is by adding a regularization term to the objective function (4), which makes the level sets compact, as in [wang2024momentum].
 However, this is not the purpose of this work which focuses on the non-regularized formulation.
-4.3 Extrapolated block coordinate descent NMD (eBCD-NMD)
-We present now a new extrapolated version of Algorithm 1, namely extrapolated BCD-NMD (eBCD-NMD).
-Our goal is to accelerate BCD-NMD while preserving convergence to a stationary point under mild assumptions.
-The eBCD-NMD method is an adaptation of the well-known LMaFit algorithm [43] for matrix completion.
-The standard matrix completion formulation in [43] is similar to the one we consider in (5), although it does not include the inequality constraint $P_{\Omega^{C}}(Z)\leq 0$, which models the crucial information that the missing entries are negative.
-Consequently, we derive a new optimization scheme, differing from the one in [43], as the update of $Z$ presents an additional nonlinear term; see equation (18).
-Moreover, we relax the hypothesis in [43] that require $\{W^{k}\}$ and $\{H^{k}\}$ to remain full-rank for all iterations.
-We present two different versions of the eBCD-NMD algorithm that, starting from the same point, produce the same low-rank approximation $W^{k}H^{k}$, yet the factors $W^{k}$ and $H^{k}$ vary.
-In what follows, we omit the iteration index, $k$, to simplify the notation, as we focus on a single iteration eBCD-NMD.
+4.3 Extrapolated BCD (eBCD)
+We want to accelerate the BCD algorithm while preserving convergence to a stationary point under mild assumptions.
+The eBCD method is an adaptation of the well-known LMaFit algorithm [wen2012solving] for matrix completion.
+The standard matrix completion formulation in [wen2012solving] is similar to the one we consider in (4), although it does not include the inequality constraint $P_{\Omega^{C}}(Z)\leq 0$, which models the crucial information that the missing entries are negative.
+Consequently, we derive a new optimization scheme, differing from the one in [wen2012solving], as the update of $Z$ presents an additional nonlinear term; see equation (17).
+Moreover, we relax the hypothesis in [wen2012solving] that requires $\{W^{k}\}$ and $\{H^{k}\}$ to remain full-rank for all iterations.
+We present two different versions of the eBCD algorithm that, starting from the same point, produce the same low-rank approximation $W^{k}H^{k}$, yet the factors $W^{k}$ and $H^{k}$ vary.
+In what follows, we omit the iteration index, $k$, to simplify the notation, as we focus on a single iteration of the eBCD algorithm.
 We assume we have a point $(Z,W,H)$ from which we perform one iteration.
-4.3.1 First version of eBCD-NMD
-Let $\alpha\geq 1$, then the first version of the eBCD-NMD scheme uses the following update: given the iterate $(Z,W,H)$, it computes
+4.3.1 First version of eBCD
+Let $\alpha\geq 1$, then the first version of the eBCD scheme uses the following update: given the iterate $(Z,W,H)$, it computes
 (17) $Z_{\alpha}=\alpha Z+(1-\alpha)WH,$
@@
 where $\left(\widehat{Z}(\alpha),\widehat{W}(\alpha),\widehat{H}(\alpha)\right)$ provides the next iterate.
-Similarly to BCD-NMD, the main computational cost for the scheme in (18) is solving the two least squares problems for computing $\widehat{W}(\alpha)$ and $\widehat{H}(\alpha)$.
+Similarly to BCD, the main computational cost for the scheme in (17) is solving the two least squares problems for computing $\widehat{W}(\alpha)$ and $\widehat{H}(\alpha)$.
 The optimization scheme in (17) differs from the one in (15) because of the new variable $Z_{\alpha}$ which is used in place of the original latent variable $Z$ to update $W$ and $H$.
@@
 $\widehat{W}(\alpha)=Z_{\alpha}H^{T}(HH^{T})^{-1}=\alpha ZH^{T}(HH^{T})^{-1}+(1-\alpha)WHH^{T}(HH^{T})^{-1}=\alpha W_{\text{BCD}}+(1-\alpha)W,$
-where $W_{\text{BCD}}$ is the update of the variable $W$ that we would have had if we performed one step of the BCD-NMD scheme (16). Moreover, if we define $\beta=\alpha-1$, then
-(19) $\widehat{W}(\alpha)=\alpha W_{\text{BCD}}+(1-\alpha)W=(1+\beta)W_{\text{BCD}}-\beta W=W_{\text{BCD}}+\beta(W_{\text{BCD}}-W),$
-that is, an extrapolation step for the BCD-NMD update of $W$, provided that $\alpha\in(1,2)$, that means $\beta\in(0,1)$.
+where $W_{\text{BCD}}$ is the update of the variable $W$ that we would have had if we performed one step of the BCD scheme (15). Moreover, if we define $\beta=\alpha-1$, then
+(18) $\widehat{W}(\alpha)=\alpha W_{\text{BCD}}+(1-\alpha)W=(1+\beta)W_{\text{BCD}}-\beta W=W_{\text{BCD}}+\beta(W_{\text{BCD}}-W).$
+Thus, line (18) represents an extrapolation step for $W_{\text{BCD}}$, provided that $\alpha\in(1,2)$, that means $\beta\in(0,1)$.
 Unfortunately, such an intuitive interpretation does not apply to the update of the variable $H$.
-Note that the step in (19) is properly referred to as an extrapolation step if $\beta$ is in the interval (0,1) or equivalently $\alpha$ is in the interval (1,2), but our scheme potentially allows for larger values that exceed the upper bound.
-4.3.2 Improved version of eBCD-NMD
-The second version of the eBCD-NMD does not require solving any least squares problem and the main computational cost is associated with a single QR factorization of a matrix in $\mathbb{R}^{m\times r}$.
-We specify that this scheme can be adapted to design alternative versions of BCD-NMD and 3B-NMD as well.
-We state the following lemma from [43] that allows us to derive the second version of eBCD-NMD.
+Note that the step in (18) is properly referred to as an extrapolation step if $\beta$ is in the interval (0,1) or equivalently, $\alpha$ is in the interval (1,2), but our scheme potentially allows for larger values that exceed the upper bound.
+4.3.2 Improved version of eBCD
+The second version of the eBCD scheme does not require solving any least squares problem and the main computational cost is associated with a single QR factorization of a matrix in $\mathbb{R}^{m\times r}$.
+We specify that this scheme can be adapted to design alternative versions of BCD and e3B as well.
+We state the following lemma from [wen2012solving] that is used in what follows.
 Lemma 4.3.
@@
 Proof 4.4.
-Let $UDV^{T}=H$ be the economy form SVD of $H$, then
-$\widehat{W}(\alpha)=Z_{\alpha}H^{\dagger}=Z_{\alpha}VD^{\dagger}U^{T}.$
+Let $UDV^{T}=H$ be the economy form SVD of $H$, then $\widehat{W}(\alpha)=Z_{\alpha}H^{\dagger}=Z_{\alpha}VD^{\dagger}U^{T}$.
 Moreover, since $Z_{\alpha}H^{T}=Z_{\alpha}VDU^{T}$, we get $\mathcal{R}(\widehat{W}(\alpha))=\mathcal{R}(Z_{\alpha}H^{T})$.
 Lemma 4.3 shows that the range space of $\widehat{W}(\alpha)$ can be obtained from that of $Z_{\alpha}H^{T}$.
-Let $Q(\alpha)$ be a matrix whose columns contain an orthogonal basis for $\mathcal{R}(Z_{\alpha}H^{T}).$ Then, using Lemma 4.3 and from the second and third lines in (18), we get
+Let $Q(\alpha)$ be a matrix whose columns form an orthonormal basis for $\mathcal{R}(Z_{\alpha}H^{T}).$ Then, using Lemma 4.3 and from the second and third lines in (17), we get
 (19) $\widehat{W}(\alpha)\widehat{H}(\alpha)=\widehat{W}(\alpha)\widehat{W}(\alpha)^{\dagger}Z_{\alpha}=Q(\alpha)Q(\alpha)^{T}Z_{\alpha}.$
-This shows that the low-rank approximation produced by the eBCD-NMD is nothing else than the orthogonal projection of $Z_{\alpha}$ into $\mathcal{R}(Z_{\alpha}H^{T})$.
-Justified by the relation in (20), we now state the improved version of eBCD-NMD.
+This shows that the low-rank approximation produced by the eBCD is nothing else than the orthogonal projection of $Z_{\alpha}$ into $\mathcal{R}(Z_{\alpha}H^{T})$.
+Justified by the relation in (19), we now state the improved version of eBCD.
 Given an iterate $(Z,W,H)$, we compute
 (20) $Z_{\alpha}=\alpha Z+(1-\alpha)WH,$
-$Q(\alpha)\rightarrow\text{ Orthogonal basis of }\mathcal{R}(Z_{\alpha}H^{T}),$
-$W(\alpha)=Q(\alpha),$
+$W(\alpha)=Q(\alpha){{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ \leftarrow}}\text{ {{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0} Orthonormal}} basis of }\mathcal{R}(Z_{\alpha}H^{T}),$
 $H(\alpha)=W(\alpha)^{T}Z_{\alpha}=Q(\alpha)^{T}Z_{\alpha},$
@@
 In practice, $Q(\alpha)$ can be computed from the economy QR decomposition of $Z_{\alpha}H^{T}$.
-Note that if $\operatorname{rank}(Z_{\alpha}H^{T})=r_{1}<r$, the first $r_{1}$ columns of the $Q$ matrix from the economy QR decomposition of $Z_{\alpha}H^{T}$ are not, in general, an orthogonal basis of the range space of the matrix.
+Note that if $\operatorname{rank}(Z_{\alpha}H^{T})=r_{1}<r$, the first $r_{1}$ columns of the $Q$ matrix from the economy QR decomposition of $Z_{\alpha}H^{T}$ are not, in general, an orthonormal basis of the range space of the matrix.
 To ensure this property, a modified version of the QR decomposition using column pivoting must be used instead.
-Specifically, there exists $Q(\alpha)\in\mathbb{R}^{m\times r_{1}}$, containing an orthogonal basis of $Z_{\alpha}H^{T}$, such that $Z_{\alpha}H^{T}\Pi=Q(\alpha)R(\alpha)$, where $\Pi$ is a permutation matrix and $R(\alpha)\in\mathbb{R}^{r_{1}\times r}$ [14].
+Specifically, there exists $Q(\alpha)\in\mathbb{R}^{m\times r_{1}}$, containing an orthonormal basis of $Z_{\alpha}H^{T}$, such that $Z_{\alpha}H^{T}\Pi=Q(\alpha)R(\alpha)$, where $\Pi$ is a permutation matrix and $R(\alpha)\in\mathbb{R}^{r_{1}\times r}$ [golub2013matrix].
 Therefore, we potentially allow for a drop in the rank of the approximation whenever $Z_{\alpha}H^{T}$ is rank-deficient.
@@
 Restarting scheme of the extrapolation parameter
-We need now to determine a rule to select the extrapolation parameter $\alpha$ at each iteration.
-Extrapolating a variable means adding information from the past iterate, thus the extrapolation parameter $\alpha$ affects how much relevant is the past information in the current update.
+We need to determine a rule to select the extrapolation parameter $\alpha$ at each iteration.
+Extrapolating a variable means adding information from the past iterate; thus, the extrapolation parameter $\alpha$ affects the relevance of the past information in the current update.
 Of course, not all the values of $\alpha$ are acceptable to progress in minimizing the residual, and a good choice of the parameter is crucial.
-We present now the updating scheme of the extrapolation parameter $\alpha$, inspired by the one introduced in [43].
-The main idea is to choose the parameter $\alpha$ such that the residual of the function is decreased sufficiently between two consecutive iterations.
+The updating scheme of the extrapolation parameter $\alpha$, is inspired by the one introduced in [wen2012solving].
+The main idea is to choose the parameter $\alpha$ such that the residual of the function is sufficiently decreased between two consecutive iterations.
 At each step, we start from a candidate value for $\alpha$, and evaluate the ratio $\delta(\alpha)$ between the norms of the residual $S(\alpha)$ at iteration $k+1$ and $S$ at iteration $k$, defined as
 (23) $\delta(\alpha)=\frac{\lVert S(\alpha)\rVert_{F}}{\lVert S\rVert_{F}}.$
-If $\delta(\alpha)<1$, we declare our update successful and the new iterate is set to $(W^{k+1},H^{k+1},Z^{k+1})=(W(\alpha),H(\alpha),Z(\alpha))$.
+If $\delta(\alpha)<1$, we declare our update successful, and the new iterate is set to $(W^{k+1},H^{k+1},Z^{k+1})=(W(\alpha),H(\alpha),Z(\alpha))$.
 Furthermore, let $\bar{\delta}$ be a fixed threshold, if $\delta(\alpha)>\bar{\delta}\in(0,1)$, meaning we do not consider the decrease of the residual fast enough, we increase $\alpha$ by setting $\alpha\leftarrow\min(\alpha+\mu,\bar{\alpha})$, where $\mu=\max(\mu,0.25(\alpha-1))$.
 If $\alpha$ reaches the upper bound $\bar{\alpha}$, we restart the parameter by setting $\alpha=1$.
-This correction does not appear in [43], but we added it to guarantee the convergence of the our algorithm.
-On the contrary, if $\delta(\alpha)\geq 1$, the step is declared unsuccessful, we do not accept the update, and we compute a standard BDC-NMD update instead by setting $\alpha=1$.
-Consequently, either we find an $\alpha\in(1,\bar{\alpha})$ that allows a decrease of the residual or we recover one iteration of the BCD-NMD, which is guaranteed to decrease the residual.
+This correction does not appear in [wen2012solving], but we added it to guarantee the convergence of our algorithm.
+On the contrary, if $\delta(\alpha)\geq 1$, the step is declared unsuccessful, we do not accept the update, and we compute a standard BCD update instead by setting $\alpha=1$.
+Consequently, either we find an $\alpha\in(1,\bar{\alpha})$ that allows a decrease of the residual, or we recover one iteration of the BCD, which is guaranteed to decrease the residual.
 Therefore, the new iterate of the algorithm produces a lower value in the residual with respect to the previous iteration.
-The overall eBCD-NMD scheme is described in Algorithm 2, where we consider the updating scheme in (21).
-We choose the improved version of eBCD-NMD because (a) it exhibits faster performance in practice; (b) computing one QR factorization is in general more stable than solving two consecutive least squares problems; (c) one of the factors in the decomposition is orthogonal and this helps demonstrate crucial properties in the convergence analysis, such as boundness of the sequence generated by the eBCD-NMD algorithm.
-Let us also mention that, in [43], convergence is proved for a scheme similar to the one in (18), instead of the scheme in (21) as done in this paper.
-Algorithm 2 eBCD-NMD
-0:  $X$, $Z^{0}$,$W^{0}$, $H^{0}$, $\bar{\alpha}$ (default $=4$), $\mu$ (default $=0.3$), $\bar{\delta}$ (default $=0.8$), maxit
+The overall eBCD scheme is described in Algorithm 2, where we consider the updating scheme in (20).
+We choose the improved version of eBCD because (a) it exhibits faster performance in practice; (b) computing one QR factorization is, in general, more stable than solving two consecutive least squares problems; (c) one of the factors in the decomposition is orthogonal and this helps demonstrate crucial properties in the convergence analysis, such as boundness of the sequence generated by the eBCD algorithm.
+Let us also mention that, in [wen2012solving], convergence is proved for a scheme similar to the one in (17), instead of the scheme in (20) as done in this paper.
+Algorithm 2 eBCD for 3B-RMD
+0:$X$, $Z^{0}$$W^{0}$, $H^{0}$, $\bar{\alpha}$ (default =4=4), $\mu$ (default =0.3=0.3), $\bar{\delta}$ (default =0.8=0.8), maxit
 1: Set $\Omega=\{(i,j)\ |\ X_{ij}>0\}$ and $\alpha_{0}=1$.
@@
 3:  $Z_{\alpha_{k}}=\alpha_{k}Z^{k}+(1-\alpha_{k})W^{k}H^{k}$
-4:     Compute $Q(\alpha_{k})$ orthogonal basis of $\mathcal{R}(Z_{\alpha_{k}}(H^{k})^{T})$
+4:  Compute $Q(\alpha_{k})$ orthonormal basis of $\mathcal{R}(Z_{\alpha_{k}}(H^{k})^{T})$
 5:  $W(\alpha_{k})=Q(\alpha_{k})$
@@
   1. Step 1
-    We first prove that there exists a range of values larger than 1 for $\alpha$ such that the scheme (21) decreases the residual, that is,
-    (25) $\lVert S\rVert_{F}^{2}-\lVert S(\alpha)\rVert_{F}^{2}>0.$
-    Specifically, we split the residual reduction in (25) into the contribution given by the update of each variable separately.
-In particular, Lemma 4.5 characterizes the residual reduction after the update of $W$ and $H$, showing that it is nonnegative.
+    We first prove that there exists a range of values larger than 1 for $\alpha$ such that the scheme (20) decreases the residual, that is, $\lVert S\rVert_{F}^{2}-\lVert S(\alpha)\rVert_{F}^{2}>0$.
+Specifically, we split this residual reduction into the contribution given by the update of each variable separately: Lemma 4.5 characterizes the residual reduction after the update of $W$ and $H$, showing that it is nonnegative.
 Lemma 4.6 and Lemma 4.7 describe the residual reduction after updating the entries of $Z$.
@@
   3. Step 3
-    We use the residual reduction that we characterized in Step 1 to prove subsequence convergence of the eBCD-NMD algorithm in Theorem 4.9, showing that a subsequence of the eBCD-NMD iterates satisfies the KKT conditions at the limit.
-This is done analyzing all possible situations that the restarting scheme arises.
+    We use the residual reduction that we characterized in Step 1 to prove the subsequence convergence of the eBCD algorithm in Theorem 4.9, showing that a subsequence of the eBCD iterates satisfies the KKT conditions at the limit.
+This is done by analyzing all possible situations that the restarting scheme arises.
 Step 1
@@
 $\lVert S\rVert_{F}^{2}-\lVert S(\alpha)\rVert_{F}^{2}=\underbrace{\lVert S\rVert_{F}^{2}-\frac{1}{\alpha^{2}}\lVert W(\alpha)H(\alpha)-Z_{\alpha}\lVert_{F}^{2}}_{=:\sigma_{WH}(\alpha)}+\underbrace{\frac{1}{\alpha^{2}}\lVert W(\alpha)H(\alpha)-Z_{\alpha}\lVert_{F}^{2}-\rVert S(\alpha)\rVert_{F}^{2}}_{=:\sigma_{Z}(\alpha)}$
-Let $U$ be an orthogonal basis of $\mathcal{R}(H^{T})$. If $H$ is full rank, we can define the orthogonal projections onto $\mathcal{R}(W(\alpha))$ and $\mathcal{R}(H^{T})$ as:
-(26) $E(\alpha)=W(\alpha)W(\alpha)^{T}=Q(\alpha)Q(\alpha)^{T},$
-$P=UU^{T}=H^{T}(HH^{T})^{-1}H.$
+Let $U$ be an orthonormal basis of $\mathcal{R}(H^{T})$. If $H$ is full rank, we can define the orthogonal projections onto $\mathcal{R}(W(\alpha))$ and $\mathcal{R}(H^{T})$ as:
+(24) $E(\alpha)=W(\alpha)W(\alpha)^{T}=Q(\alpha)Q(\alpha)^{T},\qquad P=UU^{T}=H^{T}(HH^{T})^{-1}H.$
 The following three lemmas characterize $\sigma_{WH}(\alpha)$ and $\sigma_{Z}(\alpha)$ (the proofs can be found in the appendix).
@@
 We look at the residual reduction after updating the entries of $Z$ in $\Omega^{C}$ (Lemma 4.6) and in $\Omega$ (Lemma 4.7).
-In the following lemmas, we assume that there exists an interval $[1,\widehat{\alpha}]$, such that $Z_{\alpha}H^{T}$ and $ZH^{T}$ have the same rank.
+In the following lemmas, we assume that there exists an interval $1,\widehat{\alpha}$], such that $Z_{\alpha}H^{T}$ and $ZH^{T}$ have the same rank.
 This hypothesis is needed to ensure the continuity of the residual reduction $\sigma_{Z}(\alpha)$ in a range of $\alpha>1$; more details are given in the Appendix.
 Lemma 4.6.
-Let $(W(\alpha),H(\alpha))$ be generated by the scheme in (21).
-Furthermore, assume there exists $\widehat{\alpha}>1$ such that $\operatorname{rank}(Z_{\alpha}H^{T})=\operatorname{rank}(ZH^{T})$ for $\alpha\in[1,\widehat{\alpha}]$, then there exists $\alpha_{*}>1$ such that for every $\alpha\in(1,\min(\alpha_{*},\widehat{\alpha})]$,
+[…]
 (26) $\frac{1}{\alpha^{2}}\lVert P_{\Omega^{C}}(W(\alpha)H(\alpha)-Z_{\alpha})\rVert_{F}^{2}-\lVert P_{\Omega^{C}}(S(\alpha))\rVert_{F}^{2}\geq 0.$
 Moreover, if $P_{\Omega^{C}}(\min(0,W(1)H(1)))\neq P_{\Omega^{C}}((\min(0,WH))$, a strict inequality holds.
-A similar result is proved in [43] for matrix completion.
+A similar result is proved in [wen2012solving] for matrix completion.
 In their less restrictive setting, the residual reduction after updating the entries of $Z$ in $\Omega^{C}$ is independent of $\alpha$ and nonnegative.
-However, due to the additional constraint $P_{\Omega^{C}}(Z)\leq 0$ in the 3B-ReLU-NMD formulation, the residual reduction in (28) of the eBCD-NMD algorithm is only locally nonnegative in a range where $\alpha$ is larger than one.
+However, due to the additional constraint $P_{\Omega^{C}}(Z)\leq 0$ in the 3B-RMD formulation, the residual reduction in (26) of the eBCD algorithm is only locally nonnegative in a range where $\alpha$ is larger than one.
 Analogously to Lemma 4.6, we now state a similar result that characterizes the residual reduction when the entries of $Z$ in $\Omega$ are updated.
@@
 If $P_{\Omega^{C}}(\min(0,WH))\neq P_{\Omega^{C}}(\min(0,W(1)H(1)))$, then there exists $\tilde{\alpha}>1$ such that
-(30) $\lVert S\rVert_{F}^{2}-\lVert S(\alpha)\rVert_{F}^{2}=\sigma_{WH}(\alpha)+\sigma_{Z}(\alpha)>0\quad\forall\alpha\in[1,\min(\widehat{\alpha},\tilde{\alpha})].$
-Corollary 4.8 guarantees that there exist some values of the extrapolation parameter $\alpha$ strictly larger than 1 resulting in a positive residual reduction for eBCD-NMD (21).
+[…]
+Corollary 4.8 guarantees that there exist some values of the extrapolation parameter $\alpha$ strictly larger than 1 resulting in a positive residual reduction for eBCD (20).
 Step 2
@@
 $\nabla_{Z}\mathcal{L}(Z,W,H,\Lambda,\Sigma)=Z-WH+\Lambda+\Sigma=0,$
-Primal feasibility: $P_{\Omega}(Z-X)=0,$
-$P_{\Omega^{C}}(Z)\leq 0,$
+Primal feasibility: $P_{\Omega}(Z-X)=0,P_{\Omega^{C}}(Z)\leq 0,$
 Dual feasibility: $\Sigma\geq 0,$
@@
 where the multiplier matrices $\Lambda$ and $\Omega$ measure the residual in $\Omega$ and $\Omega^{C}$ respectively, that is, $\Lambda=P_{\Omega}(WH-Z)$ and $\Sigma=P_{\Omega^{C}}(\max(0,WH))$.
-Our goal is to prove that, under some mild assumptions, a subsequence of the sequence generated by eBCD-NMD satisfies the KKT condition at the limit.
-Specifically, one can verify that the eBCD-NMD sequence satisfies the complementary slackness condition and the stationarity condition for $Z$ at any point.
-Therefore, we just need to show that the optimality conditions for $W$ and $H$ are satisfied by a subsequence of the eBCD-NMD scheme at the limit, that is,
+Our goal is to prove that, under some mild assumptions, a subsequence of the sequence generated by eBCD satisfies the KKT condition at the limit.
+Specifically, one can verify that the eBCD sequence satisfies the complementary slackness condition and the stationarity condition for $Z$ at any point.
+Therefore, we just need to show that the optimality conditions for $W$ and $H$ are satisfied by a subsequence of the eBCD scheme at the limit, that is,
 (31) $\lim_{k\to\infty}\nabla_{W}\mathcal{L}(Z^{k},W^{k},H^{k})=0,\qquad\lim_{k\to\infty}\nabla_{H}\mathcal{L}(Z^{k},W^{k},H^{k})=0.$
@@
 Theorem 4.9.
-Let $(W^{k+1},H^{k+1},Z^{k+1})$ be generated by eBCD-NMD (Algorithm 2), and assume that $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ is bounded for all $k$.
-Then there exists a bounded subsequence of $(W^{k+1},H^{k+1},Z^{k+1})$ whose limit point satisfies the KKT conditions (32).
+Let $(W^{k+1},H^{k+1},Z^{k+1})$ be generated by eBCD (Algorithm 2), and assume that $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ is bounded for all $k$. Then there exists a bounded subsequence of
+$(W^{k+1},H^{k+1},Z^{k+1})$ whose limit point satisfies the KKT conditions (30).
 Proof 4.10.
@@
 This also means that $\{Z_{\alpha_{k}}\}$ is a bounded sequence by construction.
-Moreover, the sequence $\{W^{k+1}\}$ contains orthogonal matrices whose columns are orthogonal bases for $\mathcal{R}(Z_{\alpha_{k}}(H^{k})^{T})$, thus it is bounded.
+Moreover, the sequence $\{W^{k+1}\}$ contains orthogonal matrices whose columns form an orthonormal basis for $\mathcal{R}(Z_{\alpha_{k}}(H^{k})^{T})$, thus it is bounded.
 Finally, $H^{k+1}=(W^{k+1})^{T}Z_{\alpha_{k}}$ and since $\{W^{k+1}\}$ and $\{Z_{\alpha_{k}}\}$ are bounded, then also $\{H^{k+1}\}$ is bounded.
@@
 $\mathcal{R}(I)=\mathcal{R}((Q^{k+1})^{T}Q^{k+1})\subseteq\mathcal{R}((Q^{k+1})^{T}Z_{\alpha_{k}})=\mathcal{R}(H^{k+1}),$
-where $I$ is the identity matrix of dimension equal to the number of columns of $Q^{k+1}$ or equivalently to the rank of $Z_{\alpha_{k}}(H^{k})^{T}$.
-Next, we define the set of the iteration indices where the update of $Z^{k}$ results in a positive residual reduction.
-Let =$\mathcal{F}=\left\{k\ :\ \sigma_{Z}(\alpha_{k})\geq 0\right\}$ and $\mathcal{F}^{C}$ its complement.
-Therefore, denoting $E^{k}=E(\alpha_{k})$, since $H^{k+1}$ is full rank, we can use Lemma 4.5 and summing over all $k\in\mathcal{F}$, we get
+where $I$ is the identity matrix of dimension equal to the number of columns of $Q^{k+1}$ or equivalently, to the rank of $Z_{\alpha_{k}}(H^{k})^{T}$.
+Next, we define the set of indices where the update of $Z^{k}$ results in a positive residual reduction.
+Let $\mathcal{F}=\left\{k\ :\ \sigma_{Z}(\alpha_{k})\geq 0\right\}$ and $\mathcal{F}^{C}$ its complement.
+Therefore, denoting $E^{k}=E(\alpha_{k})$, since $H^{k+1}$ is full rank, we can use Lemma 4.5, and summing over all $k\in\mathcal{F}$, we get
 (32) $\lVert S^{0}\rVert_{F}^{2}\geq\sum_{k\in\mathcal{F}}\sigma_{WH}(\alpha_{k})=\sum_{k\in\mathcal{F}}\lVert S^{k}P^{k}\rVert_{F}^{2}+\lVert E^{k}S^{k}(I-P^{k})\rVert_{F}^{2}.$
@@
 (33) $\lim_{k\to\infty,k\in\mathcal{F}}\lVert S^{k}P^{k}\rVert_{F}^{2}=0,\quad\lim_{k\to\infty,k\in\mathcal{F}}\lVert E^{k}S^{k}(I-P^{k})\rVert_{F}^{2}=0,$
-which implies
-$\lim_{k\to\infty,k\in\mathcal{F}}\lVert E^{k}S^{k}\rVert_{F}^{2}=0.$
-Let us recall that $P_{\Omega}(X)=P_{\Omega}(Z^{k})$, $P_{\Omega^{C}}(Z^{k})=P_{\Omega^{C}}(\min(0,W^{k}H^{k}))$ and using (26), we have
-$S^{k}P^{k}=(Z^{k}-W^{k}H^{k})P^{k}=(Z^{k}-W^{k}H^{k})U^{k}(U^{k})^{T}.\$
+which implies $\lim_{k\to\infty,k\in\mathcal{F}}\lVert E^{k}S^{k}\rVert_{F}^{2}=0$.
+Recall that $P_{\Omega^{C}}(Z^{k})=P_{\Omega^{C}}(\min(0,W^{k}H^{k}))$ and $P_{\Omega}(X)=P_{\Omega}(Z^{k})$.
+Using (24), we have $S^{k}P^{k}=(Z^{k}-W^{k}H^{k})P^{k}=(Z^{k}-W^{k}H^{k})U^{k}(U^{k})^{T}$.
 Since $U^{k}$ is an orthonormal basis for $\mathcal{R}((H^{k})^{T})$ and $\{H^{k}\}$ bounded,
@@
 Using equation (25) and the fact that $\alpha_{k}\leq\bar{\alpha}$ by construction, we have
-(38) $\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}\leq(\bar{\alpha})^{2}(\lVert S^{k}P^{k}\rVert_{F}^{2}+\lVert E^{k}S^{k}(I-P^{k})\rVert_{F}^{2}).$
-Combining (38) and (35) we get
-$\lim_{k\to\infty,k\in\mathcal{F}}\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}=0,$
-thus, from (37),
+(36) $\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}\leq\bar{\alpha}^{2}(\lVert S^{k}P^{k}\rVert_{F}^{2}+\lVert E^{k}S^{k}(I-P^{k})\rVert_{F}^{2}).$
+Combining (36) and (33) we get $\lim_{k\to\infty,k\in\mathcal{F}}\lVert W^{k+1}H^{k+1}-W^{k}H^{k}\rVert_{F}^{2}=0$, thus, from (35),
 (37) $\lim_{k\to\infty,k\in\mathcal{F}}\lVert S^{k}-S^{k+1}\rVert_{F}^{2}=0.$
@@
 Case 2: $|\mathcal{F}^{C}|=\infty$. Let us analyze two subcases.
-Case 2a: |$|\{k\in\mathcal{F}^{C}\ |\ \delta(\alpha_{k})>\bar{\delta}\}|<\infty$. For $k$ sufficiently large, we have
-$\lVert S^{k+1}\rVert_{F}^{2}\leq\bar{\delta}\lVert S^{k}\rVert_{F}^{2}.$
-Therefore, it holds $\lim_{k\to\infty,k\in\mathcal{F}^{C}}\lVert S^{k}\rVert_{F}^{2}=0,$ which means that the sequence converges to a global minimizer of the problem.
-Case 2b: |$|\{k\in\mathcal{F}^{C}\ |\ \delta(\alpha_{k})>\bar{\delta}\}|=\infty$.
-The eBCD-NMD set the acceleration parameter $\alpha_{k}$ to 1 for an infinite number of iterations.
+Case 2a: $|\{k\in\mathcal{F}^{C}\ |\ \delta(\alpha_{k})>\bar{\delta}\}|<\infty$.
+For $k$ sufficiently large, $\lVert S^{k+1}\rVert_{F}^{2}\leq\bar{\delta}\lVert S^{k}\rVert_{F}^{2}$.
+Therefore, $\lim_{k\to\infty,k\in\mathcal{F}^{C}}\lVert S^{k}\rVert_{F}^{2}=0,$ which means that the sequence converges to a global minimizer of the problem.
+Case 2b: $|\{k\in\mathcal{F}^{C}\ |\ \delta(\alpha_{k})>\bar{\delta}\}|=\infty$.
+eBCD sets the parameter $\alpha_{k}$ to 1 for an infinite number of iterations.
 Denote $\mathcal{J}_{1}$ the set of indices in which $\alpha_{k}=1$, a similar relation to (32) holds:
 $\lVert S^{0}\rVert_{F}^{2}\geq\sum_{k\in\mathcal{J}_{1}}\sigma_{WH}(\alpha_{k})=\sum_{k\in\mathcal{J}_{1}}\lVert S^{k}P^{k}\rVert_{F}^{2}+\lVert E^{k}S^{k}(I-P^{k})\rVert_{F}^{2}.$
-Therefore, one can replicate the proof for Case 1 and the subsequence whose indexes are in $\mathcal{J}_{1}$ satisfies the optimality conditions at the limit.
-We emphasize that, similarly to BCD-NMD, we need a strong assumption, namely the boundness of $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ which guarantees that the eBCD-NMD sequence remains bounded.
-This hypothesis serves the same purpose as the existence of the limit point in the BCD-NMD algorithm, since it allows us to exclude pathological cases where the infimum of the problem is not attained.
+Therefore, one can replicate the proof for Case 1, and the subsequence whose indices are in $\mathcal{J}_{1}$ satisfies the optimality conditions at the limit.
+We emphasize that, similarly to BCD, we need a strong assumption, namely the boundness of $\{P_{\Omega^{C}}(W^{k}H^{k})\}$ which guarantees that the eBCD sequence remains bounded.
+This hypothesis serves the same purpose as the existence of the limit point in the BCD algorithm, since it allows us to exclude pathological cases where the infimum of the problem is not attained.
 5 Numerical experiments
-In this section, we present numerical experiments on synthetic and real-world data sets, for different applications of ReLU decomposition: matrix completion with ReLU sampling, the recovery of Euclidean distance matrices, low-dimensional embedding on text data, and the compression sparse dictionaries.
-Our main purpose is to show that eBCD-NMD behaves favorably compared to the state of the art, namely Naive-NMD [34], EM-NMD [34], A-NMD [35], and 3B-NMD [35], while having convergence guarantees.
+In this section, we present numerical experiments on synthetic and real-world data sets, for different applications of RMD: matrix completion with ReLU sampling, the recovery of Euclidean distance matrices, low-dimensional embedding on text data, and the compression sparse dictionaries.
+Our main purpose is to show that eBCD behaves favorably compared to the state of the art, namely the Naive approach [saul2022nonlinear], EM [saul2022nonlinear], A-Naive [seraghiti2023accelerated], and e3B [seraghiti2023accelerated], while having convergence guarantees.
 In all experiments, we use random initialization: entries of $W$ and $H$ are sampled from a Gaussian distribution, W = randn(m,r) and H = randn(r,n) in MATLAB.
@@
 The time limit will be selected depending on the size of the data.
-For the algorithms solving Latent-ReLU-NMD, the low-rank product $W^{k}H^{k}$ in $\Gamma_{k}$ is substituted with the matrix $\Theta^{k}$.
-The first stopping criteria can be achieved only if the target matrix admits an (almost) exact ReLU decomposition.
+For the algorithms solving Latent-RMD, the low-rank product $W^{k}H^{k}$ in $\Gamma_{k}$ is substituted with the matrix $\Theta^{k}$.
+The first stopping criteria can be achieved only if the target matrix admits an (almost) exact RMD.
 The algorithms are implemented in MATLAB R2021b on a 64-bit Samsung/Galaxy with 11th Gen Intel(R) Core(TM) i5-1135G7 @ 2.40GHz and 8 GB of RAM, under Windows 11 version 23H2.
 The code is available online from https://github.com/giovanniseraghiti/ReLU-NMD .
-The parameters of eBCD-NMD are the ones specified in Algorithm 2. For the parameters of the other methods, we use the setting from [35].
+The parameters of eBCD are the ones specified as default in Algorithm 2. For the parameters of the other methods, we use the setting from [seraghiti2023accelerated].
 5.1 Matrix completion with ReLU sampling
-In this subsection, we consider the same experiment used in [26] for symmetric matrix completion with ReLU sampling, adapted to the non-symmetric case.
-Specifically, we randomly generate the entries $W$ and $H$ from a Gaussian distribution, as for our initializations, and we construct $\Theta^{t}=WH$, and set $X=\max(0,\Theta^{t})$ in the noiseless case and $\max(0,\Theta^{t}+N)$ in the noisy case, where $N=\sigma*\tilde{N}*\lVert\Theta^{t}\rVert_{F}/\lVert\tilde{N}\rVert_{F}$, $\tilde{N}=$ randn(m,n), and $\sigma\in(0,1)$.
+In this subsection, we consider the same experiment used in [liu2024symmetric] for symmetric matrix completion with ReLU sampling, adapted to the non-symmetric case.
+Specifically, we randomly generate the entries $W$ and $H$ from a Gaussian distribution, as for our initializations, and we construct $\Theta^{t}=WH$, and set $X=\max(0,\Theta^{t})$ in the noiseless case and $\max(0,\Theta^{t}+N)$ in the noisy case, where $\sigma*\tilde{N}*\lVert\Theta^{t}\rVert_{F}/\lVert\tilde{N}\rVert_{F}, N~=\tilde{N}=$ randn(m,n), and $\sigma\in(0,1)$.
 In our experiment, we set $m=n=1000$, $r=20$ and $\sigma=10^{-2}$.
@@
 The algorithms are stopped when the objective functions in (2) and (3) achieve a relative residual $\Gamma_{k}$ below $10^{-9}$ in the noiseless case, or below $\sigma=10^{-2}$ in the noisy case.
-We also report the average CPU time to reach these values in Table 1.
+We also report the average CPU time and the standard deviation to reach these values in Table 1.
 Figure 1: Matrix completion with ReLU sampling: evolution of the average relative residual $\Gamma_{k}$ w.r.t. CPU time for the noiseless case (left image) and the noisy case (right image).
-BCD-NMD eBCD-NMD 3B-NMD Naive-NMD A-NMD EM-NMD
-Noiseless 7.92 $\pm$ 0.16 0.90 $\pm$ 0.04 1.95 $\pm$ 0.04 22.0 $\pm$ 0.78 6.37 $\pm$ 0.33 26.7 $\pm$ 0.70
-Noisy 1.14 $\pm$ 0.05 0.19 $\pm$ 0.01 0.66 $\pm$ 0.02 3.17 $\pm$ 0.12 1.54 $\pm$ 0.19 2.70 $\pm$ 0.15
-Table 1: Average CPU time in seconds and standard deviation to reach a tolerance on the relative residual $\Gamma_{k}$ of $10^{-9}$ in the noiseless case, and of $10^{-2}$ in the noisy case.
+BCD eBCD e3B Naive A-Naive EM
+[…]
+Noise iter 304 121 65 308 84 296
+[…]
+Noise iter 36 22 20 41 19 28
+Table 1: Average CPU time in seconds, standard deviation, and average iteration number to reach a tolerance on the relative residual $\Gamma_{k}$ of $10^{-9}$ in the noiseless case, and of $10^{-2}$ in the noisy case.
 Best values are highlighted in bold.
 Figure 1 shows that all the algorithms reach the tolerance on the relative residual.
-Moreover, eBCD-NMD outperforms all the other methods, being considerably faster both in the noiseless and noisy cases, as Table 1 confirms.
-In fact, compared to the second best, eBCD-NMD is on average more than twice faster in the noiseless case, and more than three time faster in the noisy case (each time, the second best is 3B-NMD).
+Moreover, eBCD outperforms all the other methods, being considerably faster both in the noiseless and noisy cases, as Table 1 confirms.
+In fact, compared to the second best, eBCD is on average more than twice faster in the noiseless case, and more than three time faster in the noisy case (each time, the second best is e3B).
+Table 1 shows also that A-Naive and e3B need less iterations to converge than eBCD.
+However, their computational cost per iteration is significantly larger.
 5.2 EDMC
-[…]
-Figure 3: Euclidean distance matrix completion.
-Distribution of points on the left and relative error with the Euclidean distance matrix of the points in the right.
-The $x$-axis represents the percentage of known entries.
-We observe from Figure 3 that the distribution of the points affects the recovery performance of the algorithms:
-  • When the points are randomly distributed, $30\%$ of the entries needs to be observed for the eBCD-NMD to achieve an error below $10^{-7}$.
-  • When the points are clustered, $50\%$ of the entries needs to be observed for the eBCD-NMD to achieve an error below $10^{-7}$.
-Including the rank-1 modification in the model allows for a significantly more accurate reconstruction than solving a rank-($r+1$) 3B-ReLU-NMD.
-In general, A-NMD, eBCD-NMD and 3B-NMD provide the best reconstructions.
-However, 3B-NMD fails at least once in both examples when the percentage of observed entries is above $80\%$.
-This behavior might be due to the aggressive extrapolation procedure that 3B-NMD uses.
-We emphasize that neither the 3B-ReLU-NMD nor the Latent-ReLU-NMD formulations impose any constraints to ensure that the solution is symmetric.
+We generate 200 points in two different ways in a three-dimensional space: (1) uniformly at random in $0,10]^{3}$, and (2) randomly clustered around six clusters (4 with 30 points, and two with 40) where the centroid $(\bar{x},\bar{y},\bar{z})$ of the $i$th cluster in sampled uniformly at random in $10,10]^{3}$, and we generate the points in MATLAB using $(\bar{x},\bar{y},\bar{z})+\text{std}*\texttt{randn(1,3)}$, where std denotes the standard deviation that we fixed to std=3.
+We compute the Euclidean distance matrix $\Theta^{t}$ containing the squared distance between all couples of points and we construct our target matrix $X=\max(0,dee^{T}-\Theta^{t})$, where $d$ represents the largest value in the matrix that we want to observe.
+The upper bound on the rank is fixed to $r=5$ which is equal to the rank of $\Theta^{t}$.
+In this experiment, we assume the threshold $d$ to be known, but one might be interested in considering $d$ as an unknown of the model.
+We aim to recover all the entries of $\Theta^{t}$ larger than $d$.
+In practice, the threshold $d$ is adapted to achieve the percentage of observed entries in the $x$-axis of the pictures on the right of Figure 2.
+Specifically, we display the average relative error of the low-rank approximation with the Euclidean distance matrix $\Theta^{t}$ for different percentages of observed entries, over 10 random experiments.
+Moreover, all algorithms have been adapted to include the rank-1 modification in (5).
+We excluded the EM algorithm from the analysis because adapting it to the rank-1 modified model appears less straightforward than for the other algorithms.
+We also consider as a baseline the rank-$(r+1)$ reconstruction given by the original BCD method solving the 3B-RMD of rank $r+1$, without considering any rank-1 modification (($r+1$)-BCD).
+In this case, we compute a rank-($r+1$) decomposition of $X$ and evaluate the relative error with $dee^{T}-\Theta^{t}$.
+We stop the algorithms when a tolerance of $10^{-9}$ on the relative residual $\Gamma_{k}$ is reached or after 60 seconds.
+Figure 2: Euclidean distance matrix completion: relative errors where the $x$-axis represents the percentage of known entries.
+The left figure is for the uniform distribution of the data points, the right figure for the clustered distribution.
+The distribution of the points affects the recovery performance of the algorithms, see Figure 2: When the points are randomly distributed, $30\%$ of the entries need to be observed for the eBCD to achieve an error below $10^{-7}$.
+When the points are clustered, only $50\%$ of the entries need to be observed for the eBCD to achieve an error below $10^{-7}$.
+Including the rank-1 modification in the model allows for a significantly more accurate reconstruction than solving a rank-($r+1$) 3B-RMD.
+In general, A-Naive, eBCD, and e3B provide the best reconstructions.
+However, e3B fails at least once in both examples when the percentage of observed entries is above $80\%$.
+This behavior might be due to the aggressive extrapolation procedure that e3B uses.
+We emphasize that neither the 3B-RMD nor the Latent-RMD formulations impose any constraints to ensure that the solution is symmetric.
 However, if the number of observed entries is sufficiently large, the algorithms still recover the original symmetric solution.
 5.3 Low-dimensional embedding
-In this subsection, we test eBCD-NMD within the threshold similarity matching (TSM) framework of Saul [33] for dimensionality reduction.
-Specifically, ReLU-NMD is employed to find a lower-dimensional representation of the similarity matrix of the input points ; see Section 2 for the details.
+RMD can be employed to find a lower-dimensional representation of the similarity matrix of the input points; see Section 2.4.
 Our focus is on text data, where each row of the data matrix corresponds to a document, each column to a word, and the entries $(i,j)$ of the matrix is the number of occurrences of the $j$th word in the $i$th document.
-We consider two different well-known and already pre-pocessed text-word data sets from [44], that are k1b which has 2340 documents and 21839 words and hitech that has 2301 documents and 10080 words.
-The threshold parameter $\tau$ in the TSM framework is selected according to the statistical analysis proposed in [33], that yields $\tau=0.17$ for k1b and $\tau=0.08$ for hitech.
-Our objective is to evaluate how eBCD-NMD performs in comparison to the state-of-the-art algorithms for computing ReLU decomposition in this context.
-The quality of the embedding is evaluated using the mean angular deviation (MAD) metric as suggested in [33].
-The MAD measures the difference between the angles that the algorithm tries to preserve (namely the ones below $\cos^{-1}(\tau)$).
-Define $\mathcal{P}(\tau)=\{(i,j)\ |\ \langle z_{i},z_{j}\rangle-\tau\lVert z_{j}\rVert\lVert z_{i}\rVert>0\}$, the MAD is given by
+We consider two well-known and already pre-pocessed text-word data sets from [zhong2005generative]: k1b with 2340 documents and 21839 words, and hitech with 2301 documents and 10080 words.
+The threshold parameter $\tau$ in the TSM framework is selected according to the statistical analysis proposed in [saul2022geometrical], that is, $\tau=0.17$ for k1b and $\tau=0.08$ for hitech.
+The quality of the embedding is evaluated using the mean angular deviation (MAD) metric as suggested in [saul2022geometrical]:
 $\Delta=\frac{1}{|\mathcal{P}(\tau)|}\sum_{(i,j)\in\mathcal{P}(\tau)}\left|\cos^{-1}\left(\frac{\langle z_{i},z_{j}\rangle}{\lVert z_{i}\rVert\lVert z_{j}\rVert}\right)-\cos^{-1}\left(\frac{\langle y_{i},y_{j}\rangle}{\lVert y_{i}\rVert\lVert y_{j}\rVert}\right)\right|,$
-where the $z_{i}$’s are the higher-dimensional points, and the $y_{i}$’s the lower-dimensional embeddings.
-Figure 4 displays the MAD and average running time per iteration for increasing values of the approximation rank.
-We run all the algorithms with a fixed time limit of 60 seconds.
-Figure 4: MAD analysis and average iteration time for increasing values of the rank in the $x$-axes. k1b dataset on top and hitech dataset at the bottom.
-Even though we are not able to identify one algorithm that outperforms the others, Figure 4 shows that solving 3B-ReLU-NMD is to be preferred in this application.
-Indeed, BCD-NMD, eBCD-NMD and 3B-NMD achieve the lowest MAD in both experiments.
-In fact, these three algorithms have the lowest average iteration time, being more than one second faster per iteration than Naive-NMD, A-NMD and EM-NMD.
-For this reason, BCD-NMD, eBCD-NMD and 3B-NMD perform more iterations than Naive-NMD, A-NMD and EM-NMD in 60 seconds, resulting in more accurate approximations.
+where the $z_{i}$’s are the higher-dimensional points, and the $y_{i}$’s the lower-dimensional embeddings, and $\mathcal{P}(\tau)=\{(i,j)\ |\ \langle z_{i},z_{j}\rangle-\tau\lVert z_{j}\rVert\lVert z_{i}\rVert>0\}$, Figure 3 displays the MAD and average run times per iteration for different ranks.
+We run all the algorithms for 60 seconds.
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 3: MAD analysis and average iteration time for increasing values of the rank: The top two figures are for the k1b dataset, and the bottom two for the hitech dataset.
+Even though no algorithm outperforms the others, Figure 3 shows that using the 3B-RMD is to be preferred in this application.
+Indeed, BCD, eBCD and e3B achieve the lowest MAD in both experiments.
+Moreover, these three algorithms have the lowest average iteration time, being more than one second faster per iteration than Naive, A-Naive, and EM.
+For this reason, BCD, eBCD, and e3B perform more iterations than Naive, A-Naive, and EM within 60 seconds, resulting in more accurate approximations.
 5.4 Compression of sparse data
 The last application is the compression of sparse data.
-We consider sparse greyscale images: 10000 randomly selected images from MNIST and Fashion MNIST (fMNIST), both of size $784\times 10000$, and on two well-known sparse images, Phantom ($256\times 256$) and Satellite ($256\times 256$).
-We also use sparse, structured matrices from various applications from [10]: Trec11 ($235\times 1138$), mycielskian ($767\times 767$), lock ($1074\times 1074$), and beaconfd ($173\times 295$).
+We consider sparse greyscale images: 10000 randomly selected images from MNIST and Fashion MNIST (fMNIST), both of size $784\times 10000$, and on two well-known sparse images: Phantom ($256\times 256$) and Satellite ($256\times 256$).
+We also use sparse, structured matrices from various applications from [dai2012geometric]: Trec11 ($235\times 1138$), mycielskian ($767\times 767$), lock ($1074\times 1074$), and beaconfd ($173\times 295$).
 We choose the approximation rank $r$ to ensure a desired level of compression.
-Specifically, when computing the ReLU factors $W$ and $H$ of a sparse matrix $X\in\mathbb{R}^{m\times n}$ with $\operatorname{nnz}(X)$ non zero entries, the number of entries of $WH$ is $r(n+m)$.
-Hence, the rank $r$ is set to the closest integer such that $r\leq 0.5\cdot\operatorname{nnz}(X)/(n+m)$, which means we reduced the memory requirement by $50\%$.
-We compare eBCD-NMD with state-of-the-art methods for ReLU decomposition, and we compute the TSVD of the same rank as baseline.
-Table 2 shows the relative error $\frac{\|X-\max(0,\bar{W}\bar{H})\|_{F}}{\|X\|_{F}}$, where $\bar{W}$ and $\bar{H}$ are the factors of the rank-$r$ approximation produced by any algorithm.
-We run each algorithm for 120 seconds and display the average error over 10 different random initializations.
-Data set BCD-NMD eBCD-NMD 3B-NMD Naive-NMD A-NMD EM-NMD TSVD
-MNIST 12.4% 11.8% 11.6% 12.5% 11.7% 13.2% 25.8%
-fMNIST 9.7% 9.3% 9.3% 9.7% 9.2% 9.9% 14.0%
-Phantom 5.1% 4.1% 4.3% 5.4% 4.5% 5.6% 19.2%
-Satellite 14.0% 13.6% 13.7% 14.4% 14.0% 14.8% 26.0%
-Trec11 27.4% 26.5% 26.5% 27.2% 26.3% 31.1% 57.3%
-mycielskian 0.37% 0.05% 0.06% 0.53% 0.06% 3.6% 58.0%
-lock1074 0.3% 0.001% 0.001% 1.5% 0.006% 66.7% 67.2%
-beaconfd 20.6% 20.4% 20.0% 20.9% 20.7% 26.8% 39.1%
-Table 2: Sparse matrices compression. Comparison of the ReLU decomposition error in percentage on various data sets. The best values are highlighted in bold.
-We observe that ReLU models outperform the linear alternative TSVD, as all the algorithms yield approximations that are significantly more accurate than TSVD.
-Additionally, eBCD-NMD and 3B-NMD produce the most accurate solutions in seven out of eight examples; moreover, eBCD-NMD outperforms BCD-NMD, achieving a smaller compression error for all the data sets considered.
-We also observe that the structure of the matrix has an impact on the effectiveness of the compression using ReLU decomposition.
-Finally, we propose two visual examples of $50\%$ compression for MNIST and Phantom in Figure 5, where we compare TSVD with the approximation $\max(0,\Theta)$ provided by the eBCD-NMD algorithm.
-In both instances, one can visually appreciate how the eBCD-NMD approximation is consistently more accurate than TSVD, as further confirmed by the average error reported in Table 2.
-Original digits
-TSVD compression
-eBCD compression
-Original
-TSVD compression
-eBCD compression
-Figure 5: Comparison between $50\%$ compressed images using TSVD and eBCD on MNIST and Phantom data sets.
+Specifically, when computing the RMD factors $W$ and $H$ of a sparse matrix $X\in\mathbb{R}^{m\times n}$ with $\operatorname{nnz}(X)$ non zero entries, the number of entries of $WH$ is $r(n+m)$.
+Hence, the rank $r$ is set to the closest integer such that $r\leq 0.5\cdot\operatorname{nnz}(X)/(n+m)$, which means a compression of $50\%$.
+Table 2 shows the relative error $\frac{\|X-\max(0,\bar{W}\bar{H})\|_{F}}{\|X\|_{F}}$, where $\bar{W}$ and $\bar{H}$ are the computed factors.
+Since we are considering real-world data of different dimensions, using a single time limit that is suitable for all the dataset is not meaningful.
+Therefore, we assign a time limit to all the algorithms (except the TSVD) scaled by the dimensions of each data set: we set a maximum time limit of $T=300$ seconds for the largest data sets, namely MNIST and fMNIST with $m_{\max}=784$ and $n_{\max}=10000$.
+Then, we assign a scaled time limit $T_{\text{d}}$ for a data set of dimension $m_{\text{d}}\times n_{\text{d}}$ using $T_{\text{d}}=T\frac{m_{\text{d}}\times n_{\text{d}}}{m_{\max}\times n_{\max}}$ seconds, and display the average relative error over 10 different random initializations.
+Data set BCD eBCD e3B Naive A-Naive EM TSVD
+MNIST (300 s.) err 12.2% 11.6% 11.5% 12.3% 11.6% 12.9% 25.8%
+$r=70$ iter 1215 2159 1075 1088 899 413 /
+fMNIST (300 s.) err 9.6% 9.1% 9.2% 9.5% 9.1% 9.7% 14.0%
+$r=181$ iter 703 1498 637 914 777 513 /
+Phantom (2.5 s.) err 9.0% 6.4% 7.1 9.6% 7.8% 9.8% 19.2%
+$r=26$ iter 540 2898 508 374 340 278 /
+Satellite (2.5 s.) err 16.1% 14.9% 15.0% 17.0% 15.6% 18.2% 26.0%
+$r=12$ iter 1052 4385 983 374 352 252 /
+Trec11 (10.2 s.) err 31.3% 28.9% 28.8% 30.6% 28.8% 35.7% 57.3%
+$r=13$ iter 526 902 398 691 461 228 /
+mycielskian (22.5 s.) err 3.6% 0.6% 0.9% 8.0% 1.3% 9.0% 58.0%
+$r=14$ iter 516 1021 400 232 213 106 /
+lock1074 (44.2 s.) err 6.3% 0.1% 0.2% 15.6% 1.9% 29.2% 67.3%
+$r=12$ iter 576 1158 432 238 214 96 /
+beaconfd (2.0 s.) err 22.8 % 22.0% 21.7% 23.3% 21.7% 37.3% 39.12%
+$r=3$ iter 998 1514 680 525 419 206 /
+Table 2: Compression of sparse matrices: we compare the average final relative error and the average iteration number, within the given time limit. The best values are highlighted in bold.
+We observe that RMD outperform the TSVD, as all the algorithms yield approximations that are significantly more accurate than the TSVD.
+Additionally, eBCD produces the most accurate solutions in five out of eight examples; moreover, eBCD outperforms BCD, achieving a smaller compression error for all the data sets considered.
+We also observe that the structure of the matrix has a crucial impact on the effectiveness of the compression using RMD.
+For example, for lock1074, the improvement of RMD is impressive: the relative error of eBCD is 0.1% while that of the TSVD is 67.3%.
+On the other hand, the improvement for fMNIST is less pronounced: from 14% relative error for the TSVD to 9.6% for RMD.
+Figure 4 visually illustrates the difference between TSVD and RMD for $50\%$ compression.
+One can visually appreciate that the eBCD approximation is consistently more accurate than the TSVD, as hinted by the errors reported in Table 2.
+Refer to caption Refer to caption Refer to caption
+Original images TSVD eBCD
+Figure 4: Visual illustration: TSVD (relative error = 25.8%) against RMD with eBCD (relative error = 11.6%) on the MNIST set with $50\%$ compression ($r=70$).
 6 Conclusion
-The ReLU decomposition is a relatively recent matrix decomposition model (Saul, 2022 [33]) that finds applications in the compression of sparse data, entry-dependent matrix completion, and manifold learning.
+The RMD is a relatively recent matrix decomposition model (Saul, 2022 [saul2022geometrical]) that finds applications in the compression of sparse data, entry-dependent matrix completion, and manifold learning.
 Moreover, we showed, for the first time, how it can be used for the completion of Euclidean distance matrices when only the smallest (or largest) entries are observed.
-Our first main contribution is to establish the non-equivalence of f the two main models used to solve the ReLU problem, namely ReLU-NMD and Latent-ReLU-NMD (Corollary 3.5).
-Furthermore, we derive an explicit connection between the optimal function values of the two, demonstrating that the solution of Latent-ReLU-NMD is in the worst case twice larger than that of ReLU-NMD (Theorem 3.6).
-Our second main contribution is to prove convergence of the BCD-NMD scheme for solving 3B-ReLU-NMD by demonstrating that it falls into the general framework of exact BCD, where one of the subproblems is strictly quasi-convex (Theorem 4.1).
-Our third and most important contribution is a novel algorithm, eBCD-NMD, which uses extrapolation to accelerate eBCD-NMD.
-eBCD-NMD extends the well-known LMaFit algorithm from [43] for matrix completion.
-We proved subsequence convergence of the eBCD-NMD scheme to a stationary point under mild assumptions (Theorem 4.9).
-Finally, we provided extensive numerical results for 4 applications on synthetic and real data sets, showing that eBCD-NMD performs favorably compared to the to the state of the art, while having convergence guarantees which other algorithms lack.
-Future work includes investigating the well-posedness of the standard ReLU-NMD in (1), designing randomized algorithms to tackle large-scale problems, and exploring more advanced ReLU decompositions such as $X\approx W_{1}H_{1}+\max(0,W_{2}H_{2})-\max(0,W_{3}H_{3})$ that could handle mixed-signed and non-sparse data.
+Our first main contribution is to establish the non-equivalence of the two main models used to to compute RMDs, namely LS-RMD and Latent-RMD (Corollary 3.5).
+Furthermore, we derive an explicit connection between the optimal function values of the two, demonstrating that the solution of Latent-RMD is, in the worst case, twice larger than that of LS-RMD (Theorem 3.6).
+Our second main contribution is to prove the convergence of the BCD scheme for solving 3B-RMD by demonstrating that it falls into the general framework of exact BCD, where one of the subproblems is strictly quasi-convex (Theorem 4.1).
+Our third and most important contribution is a novel algorithm, eBCD, which uses extrapolation to accelerate BCD.
+Our new algorithm extends the well-known LMaFit algorithm from [wen2012solving] for matrix completion.
+We proved the subsequence convergence of the eBCD-NMD scheme to a stationary point under mild assumptions (Theorem 4.9).
+Finally, we provided extensive numerical results for 4 applications on synthetic and real data sets, showing that eBCD performs favorably compared to the to the state of the art, while having convergence guarantees which other algorithms lack.
+Future work includes investigating the well-posedness of the LS-RMD in (1), designing randomized algorithms to tackle large-scale problems, and exploring more advanced RMDs such as $X\approx W_{1}H_{1}+\max(0,W_{2}H_{2})-\max(0,W_{3}H_{3})$ that could handle mixed-signed and non-sparse data.
 Acknowledgments
-G.S.
-and N.G.
-acknowledge the support by the European Union (ERC consolidator, eLinoR, no 101085607).
-The work of M.P.
-and G.S.
-was partially supported by INdAM-GNCS under the project CUP E53C23001670001.
-The research of M.P.
-was partially granted by the Italian Ministry of University and Research (MUR) through the PRIN 2022 “MOLE: Manifold constrained Optimization and LEarning”, code: 2022ZK5ME7 MUR D.D.
-financing decree n.
-20428 of November 6th, 2024 (CUP B53C24006410006).
+We are grateful to the anonymous reviewers who carefully read the manuscript, their feedback helped us improve our paper.
 7 Appendix
@@
 $=$ $(Z_{\alpha}-WH)P=\alpha SP.$
-Similarly, we want to write the second term of the summation in terms of the residual. We use (41) to get
-$H(\alpha)=$ $W(\alpha)^{T}Z_{\alpha}=W(\alpha)^{T}(WH+\alpha S)=$
+Similarly, we write the second term of the summation in terms of the residual. Using (39),
+$H(\alpha)=$ $W(\alpha)^{T}Z_{\alpha}=W(\alpha)^{T}(WH+\alpha S)$
 $=$ $W(\alpha)^{T}\left[W(\alpha)R(\alpha)\Pi(HH^{T})^{-1}H-(W(\alpha)R(\alpha)\Pi(HH^{T})^{-1}-W)H+\alpha S\right]$
@@
 Therefore, the orthogonal projection operator into $\mathcal{R}(Z_{\alpha}H^{T})$ can be written as $W(\alpha)W(\alpha)^{T}=(Z_{\alpha}H^{T})(Z_{\alpha}H^{T})^{\dagger}$.
-If rank($Z_{\alpha}H^{T}$)=rank($ZH^{T}$) in some interval $[1,\widehat{\alpha}]$, then by the continuity of the Moore-Penrose inverse [37], we have
+If rank($Z_{\alpha}H^{T}$)=rank($ZH^{T}$) in some interval $1,\widehat{\alpha}$], then by the continuity of the Moore-Penrose inverse [stewart1969continuity], we have
 (44) $\lim_{\alpha\to 1^{+}}W(\alpha)W(\alpha)^{T}=\lim_{\alpha\to 1^{+}}(Z_{\alpha}H^{T})(Z_{\alpha}H^{T})^{\dagger}=(ZH^{T})(ZH^{T})^{\dagger}=W(1)W(1)^{T}.$
@@
 $\lVert P_{\Omega^{C}}(W(1)H(1)-\min(0,WH))\rVert_{F}^{2}-\lVert P_{\Omega^{C}}(\max(0,(W(1)H(1)))\rVert_{F}^{2}$
-$=$ $\lVert P_{\Omega^{C}}(W(1)H(1))\rVert_{F}^{2}+\lVert P_{\Omega^{C}}(\min(0,WH))\rVert_{F}^{2}-2\langle P_{\Omega^{C}}(W(1)H(1)),P_{\Omega^{C}}(\min(0,WH))\rangle-$
+$=$ $\lVert P_{\Omega^{C}}(W(1)H(1))\rVert_{F}^{2}+\lVert P_{\Omega^{C}}(\min(0,WH))\rVert_{F}^{2}-2\langle P_{\Omega^{C}}(W(1)H(1)),P_{\Omega^{C}}(\min(0,WH))\rangle$
 $-\lVert P_{\Omega^{C}}(\max(0,(W(1)H(1)))\rVert_{F}^{2}$
 $=$ $\lVert P_{\Omega^{C}}(\min(0,W(1)H(1)))\rVert_{F}^{2}+\lVert P_{\Omega^{C}}(\min(0,WH))\rVert_{F}^{2}$-
-$-2\langle P_{\Omega^{C}}(\min(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle-2\langle P_{\Omega^{C}}(\max(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle$
-$=$ $\lVert P_{\Omega^{C}}(\min(0,W(1)H(1))-(\min(0,WH))\rVert_{F}^{2}-2\langle P_{\Omega^{C}}(\max(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle\geq 0,$
+$-2\langle P_{\Omega^{C}}(\min(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle$
+$-2\langle P_{\Omega^{C}}(\max(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle$
+$=$ $\lVert P_{\Omega^{C}}(\min(0,W(1)H(1))-(\min(0,WH))\rVert_{F}^{2}$
+$-2\langle P_{\Omega^{C}}(\max(0,W(1)H(1))),P_{\Omega^{C}}(\min(0,WH))\rangle\geq 0,$
 which concludes the proof.